### v1

# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

Abstract
Interactive agents capable of understanding and executing instructions in the physical world have long been a central goal in AI research. The Minecraft Collaborative Building Task (MCBT) provides one such setting to work towards this goal (Narayan-Chen, Jayannavar, and Hockenmaier, 2019). It is a two-player game in which an Architect (A) instructs a Builder (B) to construct a target structure in a simulated Blocks World Environment. We focus on the challenging Builder Action Prediction (BAP) subtask of predicting correct action sequences in a given multimodal game context with limited training data (Jayannavar, Narayan-Chen, and Hockenmaier, 2020). We take a closer look at evaluation and data for the BAP task, discovering key challenges and making significant improvements on both fronts to propose BAP v2, an upgraded version of the task. This will allow future work to make more efficient and meaningful progress on it. It comprises of: (1) an enhanced evaluation benchmark that includes a cleaner test set and fairer, more insightful metrics, and (2) additional synthetic training data generated from novel Minecraft dialogue and target structure simulators emulating the MCBT. We show that the synthetic data can be used to train more performant and robust neural models even with relatively simple training methods. Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. Although modeling is not the primary focus of this work, we also illustrate the impact of our data and training methodologies on a simple LLM- and transformer-based model, thus validating the robustness of our approach, and setting the stage for more advanced architectures and LLMs going forward.
issue: 1
1 Introduction
There has long been interest in developing interactive agents that can communicate with humans about and operate within the physical world (e.g., Winograd (1971)). The goal for these agents is not only to engage in rich natural language dialogue, but also to ground it to physical objects, and execute instructions in the real world. We work toward this goal using Minecraft (https://minecraft.net/), a popular multiplayer game where players control avatars to navigate a 3D world and build structures from block-like materials. Minecraft, in recent years, has garnered a lot of interest as an AI experimentation platform. Our previous work—the Minecraft Collaborative Building Task (MCBT) and the corresponding Minecraft Dialogue Corpus (MDC) Narayan-Chen, Jayannavar, and Hockenmaier (2019)—was among the first to use it to study grounded task-oriented dialogue. Jernite et al. (2019) is another such example. More recent works, such as Mohanty et al. (2024), have been directly inspired by the MCBT. Additionally, there now exists a well-developed set of platform and tools that support AI experimentation within Minecraft, including Malmo Johnson et al. (2016), Craftassist Gray et al. (2019), TaskWorldMod Ogawa et al. (2020), MC-Saar-Instruct Köhn et al. (2020), IGLU GridWorld Zholus et al. (2022), and Narayan-Chen, Jayannavar, and Hockenmaier (2019). The latter builds on the Malmo platform.
Figure 1: In the Minecraft Collaborative Building Task, the Architect (A) has to instruct a Builder (B) to build a target structure. A can observe B, but remains invisible to B. Both players communicate via a chat interface. (NB: We show B’s actions in the dialogue as a visual aid to the reader.)
The MCBT situates a dialogue task—designed to be asynchronous, asymmetric, and with minimal language constraints—in a simulated 3D Minecraft environment where participants, with constantly shifting perspectives, must refer to and modify a dynamic world. In this task, two players collaborate: an Architect (A) instructs a Builder (B) to construct a target structure using multi-colored blocks (Figure 1). The MDC comprises 509 human-human game logs for this task. Narayan-Chen, Jayannavar, and Hockenmaier (2019) focus on generating Architect utterances, and Jayannavar, Narayan-Chen, and Hockenmaier (2020) focus on end-to-end neural models for building an automated Builder agent. The latter defined the challenging Builder Action Prediction (BAP) task, aimed at predicting the sequence of actions (block placements and/or removals) a human Builder performed at a specific point in a human-human game. Despite the task’s inherent challenges and limited training data, we demonstrate that models trained with a suitable amount of game history, enriched world representations, and a sufficient amount of synthetic, diversified data (via simple data augmentation techniques) yield promising initial results. We recap relevant prior work on the Minecraft Builder in Section 2. Building on this foundation, we next take a closer look at evaluation and training data for the BAP task, uncovering key challenges and making significant improvements on both fronts to propose BAP v2, an upgraded version of the task. This will allow future work to make more efficient and meaningful progress on it.
Given the inherent complexity of the task, we find that evaluation is nuanced and requires closer examination. We identify key challenges that hinder fair and insightful evaluation, then systematically address them. To this end, we introduce a new, cleaner v2 test set to replace the original legacy test set for evaluating BAP models (Section 3). We also refine the F1 metric, proposing a fairer variant that accounts for nuances in the data. The current evaluation, which relies solely on an aggregated and opaque F1 metric, lacks detailed insights into model behavior. To enhance insightfulness, we introduce additional metrics that measure specific model capabilities. Together, these contributions form the updated BAP v2 evaluation benchmark (Section 4).
Our work on evaluation highlights spatial reasoning as a key bottleneck in model performance, alongside the challenge of limited training data (even with the aforementioned data augmentation). Instead of merely collecting more data (which can be impractical/expensive in realistic scenarios for such complex situated dialogue tasks), we create additional diverse synthetic training data generated from novel Minecraft dialogue and target structure simulators that emulate the MCBT (Section 5). This data, though naturally simpler than it, is rich in spatial relations and referential expressions and we show that it can be successfully used to train better models, even with simple training methods (Section 6). We demonstrate that jointly training on combined synthetic and original BAP (augmented) data already improves performance. Further, a straightforward Curriculum Learning method provides an additional boost. Together, the synthetic and augmented BAP data form the BAP v2 training set. Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. We provide a detailed analysis of model performance of our best performing model relative to the baseline in Section 7.
Although modeling is not the primary focus of this work, solely to test the robustness of our synthetic data and training methodologies and show their applicability to relatively contemporary model architectures, we experiment with very simple LLM (BERT)- and transformer-based variants of our neural models. The simplicity also allows us to assess whether/to what extent simpler LLMs like BERT and vanilla Transformers alone can already effectively handle the BAP task. Our results show that the synthetic data and training methodologies improve these models as well, validating the robustness of our approach (Section 6.3). Future work can build on these insights to explore more sophisticated LLMs and architectures while leveraging the BAP v2 framework.
Section 8 provides a detailed discussion on four topics – an additional key challenge in BAP evaluation, the importance of evaluation on the synthetic data, relevant concurrent work, and broader implications of our work that go beyond BAP. The latter further highlights the potential impact of our work and directions for future research. Related work is covered in Section 9. We conclude and discuss future work in Section 10.
2 Prior Work on the Minecraft Builder
This section provides a high-level overview of the dataset, task, and baseline model, summarizing our prior work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Jayannavar, Narayan-Chen, and Hockenmaier, 2020). We also introduce some necessary formalisms and figures for better exposition.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
Figure 2: A sample sequence of human-human game states. The game starts with an empty board and an initial A instruction (a), which B executes in the first action sequence (b) by placing a single block. In (c), B begins to execute the next A instruction given in (b). However, A interrupts B in (c), leading to two distinct B action sequences: (b)–(c) (single block placement), and (c)–(h) (multiple placements and removals).
2.1 The Minecraft Collaborative Building Task and Dialogue Corpus
The Minecraft Collaborative Building Task (MCBT) is a two-player game in which player A (the Architect) has to instruct player B (the Builder) to create a copy of a Target structure that is only shown to A. B controls a Minecraft agent that is given a fixed inventory of blocks. A has access to two Minecraft windows, one which contains the Target, and one in which it can observe B’s actions. A remains invisible to B and cannot place blocks itself. A and B can only communicate via a text-based chat interface that both can use freely throughout the game. The Minecraft Dialogue Corpus (MDC) consists of 509 human-human dialogues and game logs for this task, collected via a modification of the Malmo Minecraft client (Johnson et al., 2016). In the MCBT, each block $\mathbf{b}$ has one of six colors (red, orange, yellow, green, blue, purple), and B starts with an inventory of 120 blocks (twenty of each color) that it can place into or remove from a predefined BuildRegion. Like in standard Minecraft, a discrete grid that is superimposed on the environment defines where blocks can be placed. In the Minecraft API, grid cells are indexed by integer-valued coordinates $\mathbf{c}=\langle x,y,z\rangle$, where $y$ indicates vertical height, but these coordinates are not exposed in the graphical interface, and players are unable to use them to identify specific cells in their conversations. Blocks can only be placed in empty grid cells that are either on the ground ($y=1$) or adjacent to an existing block. If the Builder picks up (removes) a block, its grid cell becomes empty again, and the block returns to the inventory.
The games in the MDC consist mainly of A providing instructions, often involving multiple actions to be taken, and grounded in the Builder’s perspective, while B executes those instructions and resolves any confusion through clarification questions and further dialogue. They are based on 150 distinct target structures, split into disjoint test, training, and development sets such that training targets do not appear during test or development. Figure 1 (left) shows the perspectives seen by each player and a snippet of their conversation (right) taken from this corpus. The environment in which B operates contains the predefined BuildRegion (shown as the white square on the ground). The task is complete when the Built structure inside the Build Region matches the Target, allowing for translations and rotations in the horizontal plane.
2.1.1 Features and Challenges
Since target structures can be fairly complex, Architects typically give step-by-step instructions ("now a yellow 6") for different parts of the target. Builders should execute these instructions, but may also ask questions ("Where does the 6 start") which the Architects has to answer ("behind the 7 from your perspective"). Architects may also need to identify and correct mistakes the Builder may have made ("too much overlap unfortunately"). A can move around freely, but remains invisible to B and views the structure from behind B when giving instructions. As a result, A instructions frequently include spatial relations, both between pairs of blocks or substructures (“put … on top of…"), and relative to B ’s current position and perspective (“left", “right"). Humans also often use higher-level descriptions to refer to complex (sub)shapes (e.g. “staircase", “v"). Due to the asynchronous nature of the dialogue, the Architect often talks while the Builder is placing blocks (“so it will look like a v"), leading to an apparent interruption of the Builder’s action sequences. Figure 2 shows an example from the MDC that highlights some of these challenges.
Floating blocks
Blocks do not need to be placed on the ground if their cell is adjacent to a supporting block. If that supporting block is later removed, the remaining block (and any structure supported by it) will “float” in place. Thus, placing floating blocks needs the placement and subsequent removal of such placeholder supporting blocks. Instructions for floating structures vary greatly, ranging from step-by-step instructions involving temporary supporting blocks to single-shot descriptions such as, simply, “place a floating yellow block”.
2.1.2 MCBT subtasks: Architect Utterance Generation and Builder Action Prediction
We define the Architect Utterance Generation (AUG) Task as the task of generating a suitable Architect utterance at any point in a human-human game at which the human Architect produced an utterance (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), and the Builder Action Prediction (BAP) Task as the corresponding task of generating a suitable Builder action sequence (consisting of block placements/removals) at any point in a human-human game at which the human Builder placed or removed blocks (Jayannavar, Narayan-Chen, and Hockenmaier, 2020). Although both of these tasks consider only a single turn, and assume the context of a game and dialogue between two human players, they are important first steps towards the creation of agents that can complete an entire game in the MCBT with a human counterpart.
2.2 The Builder Action Prediction (BAP) Task
The Builder Action Prediction (BAP) task is defined as the task of predicting the actions (block placements and/or removals) that a human Builder performed at a particular point in a human-human game. Games start with an empty board, and consist of an alternating sequence of dialogues $D^{t}$ (which may each consist of a single utterance, or a back-and-forth exchange between the two players) and Builder actions $A^{t}$ that result in an updated built structure $S^{t}$. Each action sequence $A^{t}$ yields a BAP item $(\mathcal{H}^{t-1},S^{t-1},D^{t},A^{t},S^{t})$ that consists of the game history $\mathcal{H}^{t-1}$ that culminated in the previous built structure $S^{t-1}$, the new dialogue $D^{t}$, and the new action sequence $A^{t}$ that resulted in the new structure $S^{t}$. Given $\mathcal{H}^{t-1}$, $S^{t-1},$ and $D^{t}$, BAP models $M$ should predict an action sequence $A^{t}_{M}$ that yields $S^{t}$.
Dataset statistics
The training, test and development splits of the MDC contain 3709, 1616, and 1331 items respectively, and the average sequence length of an action sequence (in the development set) is 4.3 (with a std. deviation of 4.5). Target structures in the test data do not appear in the training or development data.
Example
To illustrate this task, Figure 3 shows how the game snippet shown in Figure 2 corresponds to three different items for the BAP task. The leftmost columns in each figure display the (truncated) game history ($S^{t-2}$, $D^{t-1}$, $S^{t-1}$, and $D^{t}$), while the rightmost column shows the resulting built structure $S^{t}$. Columns are read top to bottom and left to right. Note that in this figure (and similar ones in this work), we use $S_{H}$ instead of $S$ for clarity ($H$ indicates that the structure was built by a human, not a model). To highlight the effect of the final action sequence, previously placed blocks are shown in lighter colors, and each structure is rotated based on B’s gaze (only the yaw angle) at the end of the dialog snippet below it. B’s exact position and pitch angle are omitted for clarity. Any temporary supporting blocks are not shown as well. The game starts with an empty board $S^{0}$, and an initial dialogue between the two players $D^{1}$, and the first BAP item consists of predicting that first block placement $A^{1}$ (and consequent structure $S^{1}$) (Figure 3(a)). Then, another instruction follows, which the Builder executes by placing another block, resulting in the next structure to be predicted (Figure 3(b)), and the third instruction yields a complex action sequence that results in the desired v-shape structure (Figure 3(c)).
((a)) Item 1
((b)) Item 2
((c)) Item 3
Figure 3: The snippet in Figure 2 corresponds to three BAP task items as shown.
2.3 Formalizing the BAP task
The built structure
In Minecraft, blocks can be placed into the cells $\mathbf{c}=\langle x,y,z\rangle$ of a discrete 3D grid if $\mathbf{c}$ is empty and either on the ground or adjacent to any cell $\mathbf{c}^{\prime}$ that contains a block. A cell $\mathbf{c}=\langle x,y,z\rangle$ is on the ground if its height $y=1$, and is adjacent to any cell $\mathbf{c}^{\prime}=\langle x^{\prime},y^{\prime},z^{\prime}\rangle$ if they touch on one side. [^1] Cells $\langle x,y,z\rangle$ and $\langle x^{\prime},y^{\prime},z^{\prime}\rangle$ are adjacent if either $x=x^{\prime}\pm 1,y=y^{\prime}$, and $z=z^{\prime}$, or $x=x^{\prime},y=y^{\prime}\pm 1$, and $z=z^{\prime}$, or $x=x^{\prime},y=y^{\prime}$, and $z=z^{\prime}\pm 1$ In the MCBT, there are six different BlockColors, and Builders place blocks inside a predefined build region BuildRegion, an $11\times 9\times 11$ box that contains all grid cells $\mathbf{c}=\langle x,y,z\rangle$ from $\langle x_{\min},y_{\min},z_{\min}\rangle$ to $\langle x_{\max},y_{\max},z_{\max}\rangle$. [^2] In the Minecraft Dialogue Corpus, the BuildRegion’s horizontal coordinates $x$ and $z$ range from $x_{\min}=z_{\min}=-5$ to $x_{\max}=z_{\max}=+5$ and vertical coordinates $y$ range from $y_{\min}=1$ to $y_{\max}=9$. The current Built structure $S$ can be represented as a list of blocks and their locations:
BuildRegion $=_{\textit{def}}\{\mathbf{c}=\langle x,y,z\rangle\mid x_{\min}\leq x\leq x_{\max},y_{\min}\leq y\leq y_{\max},z_{\min}\leq z\leq z_{\max}\}$
BlockColors $=_{\textit{def}}\{\textsc{Red},\textsc{Orange},\textsc{Yellow},\textsc{Green},\textsc{Blue},\textsc{Purple}\}$
$S$ $=_{\textit{def}}\{(\mathbf{c},\mathsf{c})\mid\mathbf{c}\in\textsc{BuildRegion},\mathsf{c}\in\textsc{BlockColors}\}$
The Builder
The Builder can move freely around the environment. Since spatial relations in the instructions (“the block in front of you", “to your left") often depend on the Builder’s current location and viewpoint, we record the position and pose/orientation of the Builder agent $\textsc{Pos}_{\textbf{B}}=(\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle,\phi_{\textbf{B}},\gamma_{\textbf{B}})$, where $\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle$ is a (real-valued) location (that may or may not be inside the build region), $\phi_{\textbf{B}}\in[-90,...,+90]$ indicates a pitch (vertical rotation) and $\gamma_{\textbf{B}}\in[-180,...,+180]$ a yaw (horizontal orientation). Note that $\textsc{Pos}_{\textbf{B}}$ is a critical component of game context as it dictates how spatial relations are interpreted.
The Builder’s actions
[…]
2.4 Evaluating BAP models
[…]
Definition 2.1.
Given a BAP item consisting of a (human) reference action sequence $A_{H}=\langle a_{H}^{1},...,a_{H}^{k}\rangle$ that leads from $S$ to $S_{H}$, corresponding to a reference net action set $A_{H}^{\textrm{net}}$, and a (model) predicted action sequence $A_{M}=\langle a_{M}^{1},...,a_{M}^{l}\rangle$ that leads from $S$ to $S_{M}$, corresponding to a predicted net action set $A_{M}^{\textrm{net}}$, strict Precision, Recall and F1 scores assume that a Builder action $a_{M}^{m}=(t,c,\langle x,y,z\rangle)\in A_{M}^{\textrm{net}}$ is correct if and only if there is an equal reference action $a_{H}^{h}=(t,c,\langle x,y,z\rangle)\in A_{H}^{\textrm{net}}$:
Strict $\textbf{Strict Precision}~{}P_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{M}^{\textrm{net}}|}$
Strict $\textbf{Strict Recall}~{}R_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{H}^{\textrm{net}}|}$
[…]
2.5 Data Augmentation
Since the small size of the training set (3,709 examples) is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B) (details in appendix). Empirically, we found that increasing the training data to 14,836 (4x) items gave the best performance for our baseline model (Section 2.6). We will refer to this as the original BAP (augmented) data.
2.6 A CNN- and GRU-based baseline model
Figure 4: Baseline model in this work
In Jayannavar, Narayan-Chen, and Hockenmaier (2020), we proposed a neural model for the BAP task. In this work, we build upon it and refer to it as the baseline model. The model (Figure 4) is based on a recurrent encoder-decoder architecture (Sutskever, Vinyals, and Le, 2014; Cho et al., 2014) in which a GRU-based encoder (bottom left box) captures the game context (dialogue and action history) via GloVe embeddings (Pennington, Socher, and Manning, 2014), and a CNN-based encoder (top left box) captures the world state at each time step. The world state encompasses the current built structure, action history and the Builder’s position and orientation. The decoder (right box) predicts one action per time step, based on the game history, the world state at that time, and the last action taken. It consists of another GRU backbone over action sequences (bottom right), and a multi-class classifier that reads in the output of the GRU backbone as well as the world state encoding produced by the CNN to predict either the next action (block placement or removal) to be taken, or a special STOP token that terminates the action sequence. Actions are represented as feature vectors and embedded via feedforward layers before being fed to the decoder GRU. The world state representation gets updated and re-encoded after each predicted action. We train this model with teacher forcing and cross entropy loss on the original BAP (augmented) data, and use greedy decoding (max. sequence length of 10 actions) at test time. This yields an F1 score of 21.2% on the BAP test set. More details about the model are provided in the appendix.
3 Revisiting BAP Evaluation: A Cleaner Test Set
Since the MDC consists of real human-human game logs collected in a low-stakes, free-flowing setting with minimal constraints on the participants’ language and actions, rather than standalone instructions, this strategy does not always yield BAP items that are suitable for evaluation. First, Builders might move blocks even when no new instructions were given, or before it is completely clear what they should be building next (e.g. when the Architect only instructs the Builder to “build a rectangle to the right of that", without specifying the size or color of the rectangle, let alone its orientation and distance to “that"). In such cases, the corresponding BAP items have an unclear context. Second, Builders occasionally make mistakes, and are frequently interrupted by the Architect, who may want to provide more information or simply affirm that the Builder’s actions are correct. In those cases, the BAP item’s action sequence may lead to an incorrect structure. Finally, when the board is empty, we observe that Architects do not always identify a specific location or orientation for the blocks to be placed, even when they clearly describe the structure to be built (e.g., the instruction “place a red block on the ground"). This likely occurs because there are no blocks on the board or landmarks in the environment that could serve as spatial anchor to be mentioned, and because Architects are aware that the game only requires the Builder’s structure to match the Target as long as it appears anywhere inside the Build Region and in any orientation. In those situations, two different Builders might start the game by placing the same structure in two different locations and/or orientations and both might be correct, since the context provides multiple (valid) interpretations. This phenomenon arises almost exclusively with an empty board, since the blocks on a non-empty board and the current position of the Builder typically provide enough constraints for a clear instruction to only have a unique (valid interpretation) in that case. In this section, we perform an in-depth analysis of the items in the BAP test set that identifies these phenomena, and introduce a new, cleaner BAP test set that should be used to evaluate BAP models instead of the original, legacy test set. In the next section, we revisit the strict F1 metric, and define a fairer variant that accounts for the fact that some empty-board test items have multiple valid interpretations.
3.1 Annotating the test set
We manually annotate items in the test set so that we can identify cases that have to be distinguished for evaluation purposes. Annotators were asked to annotate all items in the same dialog in sequence. For each non-empty board (NEB) item, annotators were first asked whether the context is clear and has a unique interpretation , i.e. clearly specifies what structure to be built next, and there is only a single way of placing that structure on the board (as is typically the case when the board is not empty), or whether the context is unclear, i.e. it is unclear what the Builder should do next, either because more information is required, or no new instruction was given. For empty board (EB) items, annotators could also indicate that the context is clear and has multiple interpretations , i.e. it is clear what structure should be built, even though there are multiple possible placements or orientations within the build region, as is often the case when the board is empty. If the context is clear (with unique or multiple interpretations), a second question asked whether the built structure is correct or incorrect. Details about the annotation interface, the annotation workflow, the annotation tool, data collection setup, etc. are provided in the appendix. We now illustrate the resulting categories of BAP items with a few examples.
Clear context with unique interpretation and correct structure (Figure 5)
It is clear what structure B should build next, only one such structure can be built, and the human Builder actually built that structure (Figure 5(a)). This can also arise when the board is empty (Figure 5(b)), since B is first made to stand at the bottom right corner of the board, and A clearly specifies the locations of the two blocks to be placed, given B’s new orientation.
Clear context with unique interpretation and incorrect structure (Figure 6)
Here, it is clear what structure B should build next (and only one such structure can be built), but B’s actions did not result in that structure, perhaps because B got interrupted by A, or because B made a genuine mistake. In Figure 6(a), A wants B to place two red blocks directly below the middle block of the top of the arch, but B only places one of those blocks. In Figure 6(b), A wants B to place a red block diagonally above (“on top and in front of") the topmost green block, but B places a red block in the wrong location. In the next BAP item from that dialog (Figure 6(c)), A tries to correct B’s mistake, but B now places two red blocks, only one of which should be there in the final structure.
Clear context with multiple interpretations and correct structure (Figure 7)
When the board is empty, it is often clear what structure B should build next, but the placement and orientation of that structure may not have been uniquely specified. In this case, the context is clear, although it can be (correctly) interpreted in multiple ways.
Unclear context (Figure 8)
Finally, some action sequences occur in contexts in which it is unclear what the Builder should do next. In Figure 8(a), it is unclear what exact location A implies by “the side of the red". In Figure 8(b), A only signals they had made an error, but it is unclear what the error is or what its recency is and A does not provide any information that would allow B to correct the error.
((a)) Item 1
((b)) Item 2
Figure 5: Clear context with unique interpretation and correct structure
((a)) Item 1
((b)) Item 2
((c)) Item 3
Figure 6: Clear context with unique interpretation and incorrect structure
Figure 7: Empty board: Clear context with multiple interpretations and correct structure
((a)) Item 1
((b)) Item 2
Figure 8: Unclear context
3.2 Creating a clean test set
To obtain a clean BAP test set, we remove all items with unclear contexts, as the the correctness of the built structure cannot (and should not) be assessed for the BAP task. We also automatically fix/replace incorrect structures that can be fixed in such manner, and remove the items with structures that can’t be fixed. This results in a new, BAP v2 test set that contains only items with clear contexts and correct structures, and thus suitable for fairer BAP evaluation.
Fixing incorrect structures by looking ahead to the next correct structure
If the Builder made a mistake or their action sequence was interrupted by an utterance, the built structure of a BAP item may be incorrect. E.g. in Figure 9(a) B places only one instead of two blocks. However, the next BAP item (Figure 9(b)) has a correct structure, and its structure is the one that should have been built in 9(a) because 9(b) completes the truncated action sequence. We can therefore replace the structure in 9(a) with the one in 9(b). Figure 10 shows a sequence of three BAP items: in 10(a) B misinterprets A and makes an error, in 10(b) B realizes the mistake post A’s correction and proceeds to correct it, but A interrupts them, and in 10(c), the mistake is finally rectified post the interruption. Items 10(a) and 10(b) with incorrect structures should therefore be amended to result in the correct structure built in 10(c).
Manual analysis shows that sequences of one or more incorrect items (with clear contexts) that are immediately followed by a correct item (with clear context) can be fixed by replacing their built structures with the built structure of that subsequent correct item. Those items with incorrect structures that conform to this pattern are thus fixed automatically, and the remaining that can’t are excluded from the final test set.
((a)) Item $i$ has an incorrect built structure
((b)) Item $i+1$ has a correct built structure
Figure 9: An item with an incorrect built structure (9(a)) that can be fixed by the correct built structure in the following item (9(b))
((a)) Item $i$ has an incorrect built structure
((b)) Item $i+1$ has an incorrect built structure
((c)) Item $i+2$ has the correct built structure for items $i$, $i+1$ and $i+2$
Figure 10: A sequence of items with incorrect built structures (10(a), 10(b)) that can be fixed by the following item’s correct built structure (10(c))
Statistics of the clean BAP v2 test set
The BAP v2 test set consists of 1,155 BAP items, all of which have a clear context and a correct built structure. 1,071 of these items have a unique correct interpretation, and 84 of the empty board (EB) items have multiple correct interpretations. This set was obtained from the original BAP test set of 1,616 BAP items by removing 461 items that should not be used for evaluation (418 with unclear contexts and 43 with incorrect structures that could not be fixed), and by fixing the structures of 149 items whose original structures were incorrect.
4 Revisiting BAP Evaluation: Metrics
We now revisit the strict F1 metric and propose a fairer variant that accounts for the fact that some EB items have multiple valid interpretations. Additionally, the current evaluation, relying solely on an aggregated F1 score, lacks detailed insights into model behavior. To address this, we introduce additional metrics. These contributions collectively define an updated BAP v2 evaluation benchmark.
4.1 Fairer F1
The clean v2 test set consists only of BAP items with a clear context and correct structures, but 84 of the EB items (7.3% of the v2 test set) do not have a unique interpretation, since their structure can be placed and oriented in many different ways (see Figure 7 above). The v2 test set can therefore be split up into items with a unique correct interpretation ($\mathcal{U}$) and items with multiple correct interpretations ($\mathcal{M}$). But while the strict F1 score defined in Section 2.4 is appropriate for items in $\mathcal{U}$, it unfairly penalizes Builders on items in $\mathcal{M}$ when their structure is correct, but placed in a different location and/or orientation than the reference structure. A fair evaluation metric needs to account for these allowable changes in location and orientation for items in $\mathcal{M}$, i.e. structures placed on an empty board when the context does not uniquely identify their placement.
Recall from Section 2.3 that a structure $S=\{(\langle x,y,z\rangle,\textsf{c})\}$ is a set of blocks that each have a color c and a location $\langle x,y,z\rangle$, and that the distance $\Delta(S,S^{\prime})$ between two structures $S$ and $S^{\prime}$ is defined as the number of net actions of any action sequence that changes $S$ to $S^{\prime}$ (or vice versa). To fairly compare two structures $S,S^{\prime}$ on the empty board, we first align them by searching for a translation $\mathcal{A}_{T}^{*}$ of blocks in the horizontal plane and a rotation $\mathcal{A}_{R}^{*}$ about the vertical axis in 90-degree intervals that transform $S$ into a structure $S^{*}=\mathcal{A}_{R}^{*}(\mathcal{A}_{T}^{*}(S))$ in the BuildRegionthat is as close to $S^{\prime}$ as possible (i.e. where $\Delta(S^{*},S^{\prime})$ is minimized). We refer to the composed transform $\mathcal{A}_{R}\mathcal{A}_{T}$ as alignment $\mathcal{A}$, and call $\mathcal{A}^{*}=\operatorname*{arg\,min}_{\mathcal{A}}(\Delta(\mathcal{A}(S),S^{\prime}))$ the optimal alignment of $S$ to $S^{\prime}$.
To evaluate a model’s net action set $A_{M}^{\mathrm{net}}$ against a human reference net action set $A_{H}^{\mathrm{net}}$ for an item in $\mathcal{M}$, where $A_{M}$ leads from the (empty) board to a predicted structure $S_{M}$, and $A_{H}$ leads from the (empty) board to a reference structure $S_{H}$, we therefore first identify an optimal alignment of the predicted structure to the reference, $\mathcal{A}^{*}=\operatorname*{arg\,min}_{\mathcal{A}}(\Delta(\mathcal{A}(S_{M}),S_{H}))$, and apply the same transformation $\mathcal{A}^{*}$ to all actions in the model’s net action set $A_{M}^{\mathrm{net}}$, yielding an aligned action sequence $\widetilde{A}_{M}^{\mathrm{net}}$ that can now be fairly compared against $A_{H}^{\mathrm{net}}$ by our original strict F1 metric. That is, a fairer F1 metric computes a strict F1 score for the predicted action sequences $A_{M}^{\mathrm{net}}$ for any item in $\mathcal{U}$, because the reference sequences for these items yield the only correct structures for these items, but for any item in $\mathcal{M}$, it computes a strict F1 score for the optimally aligned predicted action sequences $\widetilde{A}_{M}^{\mathrm{net}}$ because this item has multiple correct interpretations that only differ from each other in location and orientation. Thus, in the example in Figure 7, any predicted orange 3x3 U-shaped structure placed on the ground will receive a fairer F1 score of 1.0.
4.2 Auxiliary metrics
The current BAP evaluation relies solely on an aggregate F1 metric. While it reasonably reflects overall model performance (especially with fairer F1), it remains opaque and lacks detailed insight into specific model capabilities, such as spatial reasoning. This is a key missing aspect in the evaluation framework. To address this, we introduce finer-grained auxiliary metrics to complement the aggregate fairer F1 metric. These metrics aim to enhance evaluation robustness, provide deeper insight into model behavior, enable precise error analysis, better differentiate models, identify reasons behind a model’s performance changes, and ultimately guide the development of better models.
Type, Color, and Location F1
In analogy to the strict F1 score, we define auxiliary metrics that only evaluate certain aspects of the predicted actions. Recall that the strict evaluation assumes that a Builder action $a_{M}^{m}=(t,c,\langle x,y,z\rangle)\in A_{M}^{\textrm{net}}$ is correct if and only if there is an equal reference action $a_{H}^{h}=(t,c,\langle x,y,z\rangle)\in A_{H}^{\textrm{net}}$, and that our strict precision, recall and F1 scores are therefore based on the sizes of Builder and Human net actions and their intersection. To evaluate whether a Builder model correctly predicts action types $t$, colors $c$, or locations $\langle x,y,z\rangle$, we define corresponding evaluation metrics that consider Builder actions correct if and only if there is an equivalent reference action in $A_{H}^{\textrm{net}}$. We consider three different equivalence relations: Type, Color (and type), and Location:
[…]
(Note that the condition for Location equivalence is effectively the same as $t=t^{\prime}~{}\textrm{and}~{}\langle x,y,z\rangle=\langle x^{\prime},y^{\prime},z^{\prime}\rangle$). Accordingly, we can transform both net action sets $A_{M}^{\textrm{net}}$ and $A_{H}^{\textrm{net}}$ into Type, Color or Location multi-sets:
$M^{\textrm{Type}}$ =$=\{t\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
$M^{\textrm{Color}}$ =$=\{(t,c)\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
$M^{\textrm{Loc}}$ =$=\{\langle x,y,z\rangle\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
which yields Type, Color and Location (strict) precision, recall and F1 scores by comparing the sizes of the corresponding multisets and their intersections (we only show Type F1 below, as Color and Location F1 are similarly defined):
Strict Type $\textbf{Strict Type Precision}~{}P_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{M}^{\textrm{Type}}|}$
Strict Type $\textbf{Strict Type Recall}~{}R_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{H}^{\textrm{Type}}|}$
[…]
Intuitively, the Type F1 metric quantifies how good a model is at understanding what type of actions to perform – placements, or removals, or both, and how many of each type. The Color F1 metric narrows down Type F1 a bit more to factor for colors as well. It quantifies how good a model is at understanding the colors of blocks that need to be placed or removed. The Location F1 metric quantifies how good a model is at understanding the specific locations of blocks that need to be placed or removed (without regard to color), i.e., spatial reasoning. Also, as mentioned earlier, these metrics are analogous to the strict F1 metric. They can then be used to define their fairer counterparts in the same way the strict F1 was used to define the fairer F1. Henceforth, we will use Type, Color and Location F1 to refer to their fairer counterparts.
Shape F1
To evaluate how good a Builder model is at understanding the "shape" in which the required blocks need to be placed or removed, we define a metric, Shape F1, that is agnostic to the absolute $\langle x,y,z\rangle$ locations of the required blocks, but only cares that they be placed/removed in the right locations relative to each other. Essentially, it compares net actions modulo their exact placement and orientation. Therefore, similar to the methodology used for fairer F1 (Section 4.1), we optimally align the model’s net actions $A_{M}^{\mathrm{net}}$ against the human reference net actions $A_{H}^{\mathrm{net}}$ (instead of the corresponding resulting built structures themselves like was done for fairer F1), yielding aligned net actions $\hat{A}_{M}^{\mathrm{net}}$ that can now be compared against $A_{H}^{\mathrm{net}}$ by our original strict F1 metric. This is done for all items in the v2 test set (and not just the ones with multiple interpretations like was done for fairer F1). We will see examples of Shape F1 and the other auxiliary metrics later in Sections 7.2 and 8.1.
4.3 Micro vs Macro F1
Current BAP evaluation reports the micro-averaged F1 (micro F1) over all action sequences in the test data (Section 2.4), reflecting a model’s quality in predicting individual actions. However, since the BAP task requires sequence prediction, it’s equally important to assess the model’s quality in predicting entire action sequences using the macro-averaged F1 (macro F1). Intuitively, this is also particularly relevant for Shape F1, which measures the model’s performance on a broader sequence-level property (the "shape" of the net actions). Therefore, hereafter, we will report both micro and macro F1.
4.4 BAP v2 evaluation benchmark
The above metrics enable a fairer, robust, and more insightful evaluation on the clean v2 test set. We use the fairer F1 metric to report both micro and macro F1 scores in Tables 1 and 2 respectively. For each, we include both auxiliary and overall F1 metrics, and report them on the overall/full test set and separately for the EB and NEB subsets as well. This forms the BAP v2 evaluation benchmark. Our approach also enables more accurate evaluation within the EB and NEB categories, offering a clearer understanding of model performance across these two distinct scenarios. This division is important for the following reasons:
  • EB examples are generally the easiest for the BAP task, making EB performance a basic competency or sanity test, and it also reflects a model’s maximum potential quality.
  • EB performance is also an estimate of the ability of a model to begin a game accurately – an elementary yet important trait to measure. Moreover, in an interactive/online setting, from a user (human Architect) experience perspective and considering the potential for subsequent cascading errors, mistakes made in EB scenarios can potentially be costlier than those in NEB, as EB marks the start of a game (even though EB examples in the data are far fewer than NEB). Quantifying EB performance will help provide a sense for how a model can potentially fare in an interactive setting wrt this aspect.
We begin with the micro scores in Table 1 for direct comparison with our prior work. Recall that the baseline model achieves an F1 score of 21.1% under the strict F1 metric. This increases to 27.3% under the fairer F1 metric (the overall F1 on the overall dataset), a notable difference of 6.2 percentage points and a 29.4% improvement. This highlights the importance of the high quality cleaner test data and the fairer F1 metric in providing a fairer reflection of model capability. On the overall dataset, the model achieves reasonably good F1 scores of 65.2% for Type and 62.4% for Color, but only 27.7% for Location F1, which is also very close to the overall F1 of 27.3%. This suggests that spatial reasoning (Location F1) is the key bottleneck for high performance on this task. As expected, the model performs better on EB than NEB across all metrics. Recall that the number of EB items are fewer than NEB items though (and the metrics on the overall dataset will therefore be more influenced by NEB). Next, we turn to the macro scores in Table 2. While all values increase compared to the micro scores, the overall trends remain the same. The overall F1 on the overall test set increases from 27.3% (micro) to 28.1% (macro). The model achieves a Shape F1 of 36.3% (micro) and 41.1% (macro) on the overall test set, with the macro score providing a more intuitive measure in this context, as discussed in Section 4.3.
Hereafter, we will use the v2 evaluation to report model performance.
Dataset Type Color Location Shape Overall
EB 76.4 75.2 53.5 56.0 53.2
NEB 63.8 60.8 24.6 33.8 24.2
Overall 65.2 62.4 27.7 36.3 27.3
Table 1: Micro F1 scores for the baseline model under the v2 benchmark
Dataset Type Color Location Shape Overall
EB 80.9 79.8 61.5 64.9 61.2
NEB 73.1 69.9 25.7 39.0 25.2
Overall 73.8 70.7 28.5 41.1 28.1
Table 2: Macro F1 scores for the baseline model under the v2 benchmark
5 Revisiting BAP Data: Generating Synthetic Data
One key challenge of the BAP task is the limited training data, despite the data augmentation (Section 2.5). We aim to avoid simply collecting more data like the MDC, due to scalability issues and the desire to stay aligned with many realistic scenarios, where data collection for complex situated dialogue tasks, like the MCBT, can be impractical/expensive, warranting exploration of alternative methods. Additionally, we observe that the baseline model struggles primarily with spatial reasoning, as reflected in its Location F1 score. These reasons motivate the need for additional synthetic training data that, while naturally simpler than the MDC, remains rich in spatial relations and referential expressions to help train more robust models. Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. (We, in fact, also illustrate the impact of this data on training a simple LLM and transformer-based model in Section 6.3.)
We design novel Minecraft dialogue and target structure simulators that emulate the MCBT, and use them to generate a set of three synthetic dialog datasets with corresponding target structures . Each dataset is modeled after the MDC and is understandably simpler, while preserving essential elements of realistic human behavior observed in the MDC as much as possible, such as the language, clarification exchanges, Architect A ’s and Builder B ’s planning perspectives, dialog flow, etc. Across these datasets, we vary the level of instruction abstraction, the complexity of spatial relations and referential expressions, the complexity of target structures, etc. Special emphasis is laid on spatial relations as that is a core part of our motivation here. Each dataset is thus unique, contributing to the overall diversity and richness of the overall set.
Note that although our primary focus is on BAP, the simulators and data introduced here are applicable to other MCBT subtasks as well (e.g. the AUG subtask; Section 2.1.2), supporting advancements across the broader MCBT ecosystem. We discuss this in more detail in Section 8.4.
5.1 General dialog simulation framework
5.1.1 Desiderata
Similar to the target structures and game logs in the MDC, the following components of synthetic dialogues need to be generated:
  • The target structure
  • The dialogue/game – a chronological sequence that interleaves the following elements:
    • A dialogue between A and B, consisting of utterances that include instructions, potential clarification exchanges, etc.
    • B’s position and orientation at appropriate points in the game, i.e., when an utterance or action is recorded (Grounded elements of the dialog, e.g. spatial relations, are always wrt this frame of reference.)
    • B’s actions, which reflect the evolving structure being built
5.1.2 High-level Algorithm
As described in Section 2.1, the MCBT task is asynchronous and loosely structured. The MDC dialogs are real human-human game logs collected in a low-stakes, free-flowing setting with minimal constraints on participants’ language and actions. To enable tractable simulation and facilitate simpler synthetic data generation, we make some simplifications and structure the task as follows.
Our dialog simulators generate synthetic dialogs that meet the desiderata, following a general framework. At a high level, it is an iterative procedure that continues until the target structure is built. The target structure may be provided as input or generated dynamically. In the former case, a separate simulator generates target structures. Each iteration produces one dialog segment—A’s instruction followed by B’s actions in response, with optional clarification exchanges and confirmations—leading to the construction of a part of the target structure. We are essentially simulating A and B, including their latent planning processes. Each iteration proceeds as shown below. (We use Figure 11 as an example. It illustrates a BAP item $(\mathcal{H}^{t-1},S^{t-1},D^{t},A^{t},S^{t})$ from one of our synthetic datasets (recall notation from Section 2.2). In this iteration, the dialog and actions following $S^{t-1}$ are simulated to produce the next structure, $S^{t}$ – specifically, simulating $D^{t}$, $A^{t}$, and $S^{t}$.)
  1. 1.
    Architect Planning: First, A’s plan for the next steps is simulated, including which blocks to place or remove and their order if multiple blocks are involved. A reference block is chosen when necessary, enabling A to describe the new block(s) with appropriate spatial relations wrt the reference block. E.g., in Figure 11, A decides to place the red floating block, using the previously placed blue block as the reference.
  2. 2.
    Builder Position and Orientation: A "valid" position and orientation for B is sampled, ensuring a frame of reference for the utterances in the dialog segment. B is positioned at a suitable distance from the action area and oriented to face it while adhering to world constraints. E.g., in Figure 11, the structure $S^{t-1}$ is rotated according to B’s gaze (yaw angle only).  [^4] B’s position and pitch angle is omitted in our figures for clarity as noted in Section 2.2
  3. 3.
    Architect Instruction: An instruction from A is synthesized, incorporating necessary block details (e.g. color), spatial relations, and referential expressions, with template-based lexical and syntactic variations. E.g., in Figure 11, A instructs, "put a floating block …".
  4. 4.
    Optional Clarification Exchange: Some information may be omitted from the instruction with a small probability, allowing B to ask a clarification question. An appropriate exchange is then generated, with B asking and A providing the missing information. E.g., in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
  5. 5.
    Builder Actions and Low-Level Planning: B’s actions to add/remove blocks are then simulated, including any necessary low-level planning for deciding order of actions and accounting for temporary supporting blocks. This updates the built structure. E.g., in Figure 11, B places the red floating block to produce structure $S^{t}$, temporarily adding/removing support for the floating block.  [^5] Temporary supporting blocks are not shown in our figures as noted in Section 2.2.
  6. 6.
    Optional Confirmation from Builder: Optionally, a confirmatory utterance from B, like "done," may be synthesized with a small probability.
Figure 11: A BAP item from one of our synthetic datasets
We implement this general framework in three distinct ways, resulting in three different simulators. Each simulator uniquely instantiates steps 1, 3, and 4, along with independent target structure generation, while steps 2, 5, and 6 remain largely consistent across simulators. For brevity, we provide only a summary of each simulator in the following sections, with detailed descriptions in the appendix.
5.1.3 Spatial Relations
Given the special emphasis on spatial relations in our dataset design, we briefly outline them here. We focus on the following types of spatial relations:
  1. 1.
    Relative to a Reference Block: We can describe the location of a target block in relation to another block, referred to as the reference block. Around the reference block, there are 6 rows, 12 quadrants, and 8 octants where the target block may be positioned. This leads to the following spatial relations (all of them are considered with respect to B’s POV/frame of reference):
    1. (a)
      1D: The target block lies in one of the rows, differing from the reference block in only one dimension. This results in 6 basic spatial relations: left, right, top, bottom, front, and behind (e.g., "to the right of …", "behind …").
    2. (b)
      2D: The target block lies in one of the quadrants, differing in exactly two dimensions. This results in 12 possible combinations of the 1D relations, such as top+right, bottom+left, etc. (e.g., "to the top right of …").
    3. (c)
      3D: The target block lies in one of the octants, differing in all three dimensions. This results in 8 possible combinations of the 1D relations, such as top+right+front (e.g., "one to the left, one above, and one block behind …").
    These 1D/2D/3D relations apply when one needs to specify spatial relations between blocks or substructures/shapes.
  2. 2.
    Using B as a Spatial Anchor: Spatial relations can also be described using B as a spatial anchor directly (e.g., "going up and to the left of you"). These only apply when one needs to specify the direction in which to build a certain substructure/shape, e.g., a diagonal.
We will see concrete examples of these in the following sections.
5.2 Random target structures and dialogs
Dialogs
The dialogs correspond to random target structures generated dynamically during the dialog simulation (Section 5.1.2) which runs for a maximum number of iterations. Blocks are placed/removed randomly, and one at a time in each iteration, with limited planning on A ’s part for block placement. As a result, the target structure is disordered, and A ’s instructions are fairly low-level, addressing only one block at a time. Most of them use spatial relations relative to arbitrary blocks identified by color and location (e.g., "to the right of the red" or "to the right of the bottom-most red block"). Figure 12 shows an example BAP item from this dataset. A identifies the orange block as a reference block and uses a 1D spatial relation "behind" to specify the target yellow block.
Figure 12: A BAP item from the random structure-based dialogs
Target structures
Target structures are generated dynamically during the dialog simulation and correspond to the final built structure at the simulation’s end. Some global constraints are enforced:
  1. 1.
    The structure must have at least one block on the ground.
  2. 2.
    The entire structure must be "connected," i.e., each block shares at least one face or edge with at least one other block.
Figure 13 provides an example.
Figure 13: A random target structure
5.3 Shape-based target structures and dialogs
Target structures in the MDC are not random; they are far more ordered (e.g., a dining table) and often composed of meaningful shapes/concepts (e.g., rows, towers, diagonals, planes, etc.). Architect instructions describe these structures at various levels of abstraction, from high-level references to the whole structure ("flower" or "bell tower"), mid-level references to sub-shapes ("row" or "plane"), down to low-level instructions referencing individual blocks. This observation motivates the need for similarly structured synthetic data as well. We first describe a simulator that can generate these structures (Section 5.3.1), and then outline two types of dialogs simulated based on these structures (Sections 5.3.2 and 5.3.3).
5.3.1 Shape-based structures
We define six unique elementary 3D shapes: rows, diagonals, T-shapes, L-shapes, U-shapes, and planes. These shapes were chosen for their simplicity and occurrence in MDC dialogs and structures. Formal definitions of these shapes are provided in the appendix. We design a simulator that generates target structures by combining multiple randomly sampled elementary shapes into a single composite structure. Sampling is performed with replacement, allowing multiple instances of the same shape type within a structure. For each shape instance, properties such as size, color, orientation, and exact 3D location are randomly sampled within specific constraints (e.g., minimum and maximum size bounds). Some global constraints are also enforced (similar to those for random structures in Section 5.2, with an added constraint for shapes):
  1. 1.
    The structure (i.e. at least one shape) must have at least one block on the ground.
  2. 2.
    The entire structure must be "connected", i.e., each block shares at least one face or edge with at least one other block, and each shape instance shares at least one block face or edge with at least one other shape instance.
The simulator is flexible, allowing parameters such as shape types, number of shapes per structure, etc. to be customized. Figure 14 illustrates two examples of such target structures.
Figure 14: Shaped-based target structures
5.3.2 Blocks-based dialogs for shape-based structures
Target structures
Each target structure consists of three shape instances, sampled from the full set of the six elementary shapes.
Dialogs
The dialog simulation (Section 5.1.2) takes the target structure as input and continues iteratively until the target structure is built. Construction proceeds one shape at a time. A ’s planning is involved and follows a heuristic that selects one or more blocks to place in each iteration until all shapes are completed. Consequently, A ’s instructions reference a single block or multiple ones collectively, with each shape decomposed into multiple instructions. Most of A ’s instructions include spatial relations relative to the last block placed by B (e.g., "now, place two red blocks behind the last block"). Figure 15 illustrates an example BAP item from this dataset. The first blue square has been built, and the second (yellow) shape is under construction. A is instructing B to continue building the latter by placing three more blocks, using a 1D spatial relation "on top of" relative to the last block. (In the previous step, A used a 2D spatial relation, "diagonally in front of and to the right of".)
Figure 15: A BAP item from the blocks-based dialogs for shape-based structures
5.3.3 Shape-based dialogs for shape-based structures
Target structures
For tractability in dialog simulation and to prevent overly complex utterances, each target structure is limited to two shape instances, sampled from only three elementary shapes: rows, diagonals, and planes.
Dialogs
The dialog simulation (Section 5.1.2) takes the target structure as input and proceeds shape by shape, with one complete shape placed per iteration until the target is built. A ’s planning employs simple heuristics to determine the order of shapes and select a target block within each shape for defining spatial relations, which also serves as B ’s starting point for building the shape. A ’s instructions, therefore, describe entire shapes rather than individual blocks, resulting in higher-level, abstractive instructions. Since each target structure consists of two shapes, each placed in one go, a 1D/2D/3D spatial relation is only used to specify the position of the initial block of the second shape relative to the last block placed in the first shape. Additionally, A can also use spatial relations that use B as a spatial anchor (Section 5.1.3), for building either shape. Figure 16 illustrates a BAP item from this dataset. The first shape (a yellow column) has been built, and A is instructing B to build the second shape (a red row) in one go. A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row, along with the 3D spatial relation "one to the left, two underneath, and one block in front of" relative to the last block (the topmost yellow one), to indicate the starting position for the row.
Figure 16: A BAP item from the shapes-based dialogs for shape-based structures
6 Training with Synthetic Data
Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with simple training methods, thereby addressing the challenge of insufficient training data. We first show that simple joint training on the aggregation of the synthetic data and the original BAP (augmented) data already yields a better model (Section 6.1). We then go a step further and show that a simple Curriculum Learning method can further boost performance (Section 6.2). The synthetic data together with the original BAP (augmented) data now constitutes the BAP v2 training dataset, completing the BAP v2 framework. We also illustrate the impact of the synthetic data and training methodologies on training a simple LLM (BERT) and transformer-based model (Section 6.3).
A note on experimental design choices
For the GRU-based models in Sections 6.1 and 6.2, we maintain the same architecture and hyperparameters as the baseline model (Section 2.6), along with the same amount of data augmentation in the BAP training data as in our prior work (Section 2.5). Fine-tuning these aspects for different training methods is beyond the scope of our objectives here and left for future work to further optimize performance. Similar choices made for the BERT-based models are discussed later in Section 6.3.
Key Terminology
Hereafter, we refer to the random structure-based dialogs (Section 5.2) as $D_{\mathrm{r}}$, blocks-based dialogs for shape-based structures (Section 5.3.2) as $D_{\mathrm{bs}}$, and shape-based dialogs for shape-based structures (Section 5.3.3) as $D_{\mathrm{ss}}$. The Minecraft-based original BAP data is denoted as $D_{\mathrm{mc}}$. Each dataset includes train, test, and val splits, with the $D_{\mathrm{mc}}$ train set also incorporating augmented data from Section 2.5. The baseline model (Section 2.6), trained on $D_{\mathrm{mc}}$ alone, is denoted as $M_{\mathrm{mc}}$.
6.1 Joint training on aggregated data
The most straightforward approach to use the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) is to just combine them all together along with the original $D_{\mathrm{mc}}$ data and retrain the baseline model on this aggregated dataset.
6.1.1 Experimental Setup
Our setup closely follows that of Section 2.6, where the baseline $M_{\mathrm{mc}}$ was trained on $D_{\mathrm{mc}}$ alone. We retain the same exact model architecture and hyperparameters as in $M_{\mathrm{mc}}$. Training employs teacher forcing with cross-entropy loss, and decoding is performed via greedy decoding, with a maximum sequence length of 20 actions. We use the fairer F1 metric to report micro F1 scores on the BAP v2 test set for $D_{\mathrm{mc}}$ (Section 4.4) and the test sets of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$.  [^6] For fairer F1 on $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $\mathcal{M}$ consists of all EB items, unlike $D_{\mathrm{mc}}$, where only a subset is considered (Section 4.1). This is because, by design, all EB items in $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ have multiple correct interpretations. Results on the full v2 evaluation benchmark are reported for the baseline and best performing models in this work later in Section 7.1.
How much of each synthetic dataset to use during train time?
An important consideration is determining the optimal amount of each synthetic dataset to use alongside $D_{\mathrm{mc}}$ in the training dataset. We adopt an empirical yet systematic approach to address this, yielding a data mix that optimizes $D_{\mathrm{mc}}$ performance while also being highly performant on the synthetic data as much as possible. Details can be found in the appendix. Table 3 presents the final data statistics. Similar to the methodology used for data splits for the MDC  (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), for each of the synthetic datasets, the game logs are split into disjoint test, training, and validation sets such that training target structures do not appear during test or validation. We now have a total of 52419 items in the aggregated training dataset.
Dataset Train Val Test
$D_{\mathrm{bs}}$ 9890 1186 1181
$D_{\mathrm{ss}}$ 11868 1000 1000
$D_{\mathrm{r}}$ 15825 1161 1089
$D_{\mathrm{mc}}$ 14836 1331 1616
Table 3: Data statistics (#items) for training, validation, and test splits across datasets.
6.1.2 Experiments and Results
Table 4 presents the results. We denote the model trained on the aggregated dataset as $M_{\mathrm{agg}}$, and include comparisons to the baseline model ($M_{\mathrm{mc}}$– trained on $D_{\mathrm{mc}}$), and models trained on the three synthetic datasets separately ($M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$– trained on $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ respectively). F1 performance is reported (as percentages) on the respective test sets of all four datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$).
$M_{\mathrm{agg}}$ achieves an F1 score of 30.3% on $D_{\mathrm{mc}}$, up from the baseline model $M_{\mathrm{mc}}$ ’s 27.3%, a notable 3-percentage point improvement. Additionally, $M_{\mathrm{agg}}$ is highly performant across the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}$ by large margins on these. For a sense of what high performance means on these datasets, we reference the diagonal entries in rows for models $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, representing performance of these models on the same test data distribution as train: 83.9% for $D_{\mathrm{bs}}$, 75.4% for $D_{\mathrm{ss}}$, and 64.4% for $D_{\mathrm{r}}$. In light of these numbers, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. In contrast, $M_{\mathrm{agg}}$ is more robust, and retains much of the diagonal-entry performance on the synthetic datasets while also significantly improving on $D_{\mathrm{mc}}$. It is also better on the synthetic datasets than on $D_{\mathrm{mc}}$ by large margins (as it should be), unlike $M_{\mathrm{mc}}$. Lastly, the models trained solely on synthetic datasets perform poorly on $D_{\mathrm{mc}}$, as does $M_{\mathrm{mc}}$ on synthetic data, indicating that joint training on the aggregated dataset is essential for robust performance across both real and synthetic data.
Model Train Data Test Data
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
$M_{\mathrm{mc}}$ $D_{\mathrm{mc}}$ 35.2 19.1 12.4 27.3
$M_{\mathrm{bs}}$ $D_{\mathrm{bs}}$ 83.9 12.9 44.9 6.3
$M_{\mathrm{ss}}$ $D_{\mathrm{ss}}$ 11.3 75.4 7.9 9.1
$M_{\mathrm{r}}$ $D_{\mathrm{r}}$ 43.8 10.1 64.4 7.7
$M_{\mathrm{agg}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 83.4 68.8 63.6 30.3
Table 4: Models trained on datasets separately ($M_{\mathrm{mc}}$, $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$) and together ($M_{\mathrm{agg}}$)
6.2 Curriculum Learning (CL)
Another key observation from Table 4 is that performance of the $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, and $M_{\mathrm{mc}}$ models decreases in the order $M_{\mathrm{bs}}$ > $M_{\mathrm{ss}}$ > $M_{\mathrm{r}}$ > $M_{\mathrm{mc}}$, suggesting that the complexity/difficulty of examples increases in the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$, with the most complex $D_{\mathrm{mc}}$ examples naturally appearing at the end. This observation raises the possibility of using this ordering to improve over the simple joint training approach (where data is randomly ordered) when training on the aggregated dataset, so as to achieve even better performance on the complex $D_{\mathrm{mc}}$. Curriculum Learning (CL) is a natural next step to explore in this direction. CL is a training strategy that sequences training data from simpler to more complex examples, similar to human learning. Introduced by Bengio et al. (2009), CL begins with easier examples, gradually increasing in difficulty. This progression allows the model to first learn basic patterns and representations, building a foundation before tackling more challenging cases, which can enhance generalization and improve performance on complex tasks.
6.2.1 CL Mechanics
Difficulty Metrics
A key prerequisite in CL is establishing difficulty metric(s) that quantify how difficult a training example is for a model to learn, allowing the data to be sorted accordingly during training. In this work, we explore three simple, natural, and intuitive difficulty metrics, along with their combinations (e.g., sorting first by one metric and then sorting (breaking ties) by another). The first metric is the dataset $D$ an example belongs to ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$). As mentioned before, the datasets can be ordered as $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$ in terms of difficulty. Additionally, we consider the lengths of the action (output) and input sequences: the action (output) sequence length $L_{A}$ reflects the number of actions in the gold action sequence, while the input sequence length $L_{I}$ measures the number of tokens in the game history fed to the model. These are also intuitive as longer sequences are harder to generate as output and process as input.
CL Algorithm
We make use of a simple CL algorithm, specifically using the Curriculum Arrangement algorithm within the CL framework proposed by Xu et al. (2020). Details of the method are provided in the appendix.
6.2.2 Experiments and Results
Overall experimental setup remains similar to the batch training one (Section 6.1.1). We experiment with each of the three difficulty metrics individually and the four possible combinations of them, making a total of seven metrics – $D$, $L_{A}$, $L_{I}$, $(D,L_{A})$, $(D,L_{I})$, $(L_{I},L_{A})$, and $(D,L_{I},L_{A})$ (ordering within a tuple denotes the sorting order). We also experiment with the CL algorithm’s hyperparameter, $N$, trying six values – 5, 10, 15, 20, 25, and 30. (This controls the number of buckets/difficulty levels – see appendix.) This results in a grid of $7\times 6=42$ experiments.
Table 5 presents the best model, $M_{\mathrm{agg+cl}}$, selected based on $D_{\mathrm{mc}}$ performance. For comparison, we also include the baseline $M_{\mathrm{mc}}$, and the aforementioned $M_{\mathrm{agg}}$ model (Table 4). On $D_{\mathrm{mc}}$, $M_{\mathrm{agg+cl}}$ achieves an F1 of 33.0%, a notable 2.7-point improvement over $M_{\mathrm{agg}}$, and an even more significant 5.7-point improvement over the baseline $M_{\mathrm{mc}}$. It also continues to be highly performant on the synthetic data similar to $M_{\mathrm{agg}}$. Thus, Curriculum Learning on the aggregated data provides a meaningful performance boost on the BAP task over simple joint training. Hereafter, we refer to $M_{\mathrm{agg+cl}}$ as the best model.
Model Train Data Test Data
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
$M_{\mathrm{mc}}$ $D_{\mathrm{mc}}$ 35.2 19.1 12.4 27.3
$M_{\mathrm{agg}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 83.4 68.8 63.6 30.3
$M_{\mathrm{agg+cl}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 84.0 68.8 60.3 33.0
Table 5: Curriculum learning ($M_{\mathrm{agg+cl}}$) compared to models $M_{\mathrm{mc}}$ and $M_{\mathrm{agg}}$
6.3 Transformers and LLMs (BERT)
As noted in Section 1, modeling is not the primary focus of this work. However, solely to test the robustness of our synthetic data and training methodologies and show their applicability to relatively contemporary model architectures, we experiment with a very simple alternative version of the baseline model (Section 2.6) that incorporates a Transformer architecture (Vaswani et al., 2017) and an LLM (BERT; Devlin et al. (2019)). Specifically, we replace the GRU-based encoder (and GloVe embeddings) with BERT and substitute the GRU decoder with a Transformer, keeping the rest of the architecture unchanged (Figure 4). (A projection layer is added atop BERT for dimensionality reduction.) The simplicity also allows us to assess whether/to what extent simpler LLMs like BERT and vanilla Transformers alone can already effectively handle the BAP task. Future work can build on the insights we obtain to explore more sophisticated LLMs and architectures while leveraging the BAP v2 framework.
Model details
To control the representation dimensionality and complexity of the decoder transformer, we add a simple feedforward projection layer atop BERT, compressing its output (as otherwise the dimensionality of the decoder transformer would be forced to be the same as the BERT encoder and make it overly complex than need be). The BERT encoder and transformer decoder are connected in standard encoder-decoder fashion but via the projection layer. BERT’s token embeddings are compressed by the projection layer, and the transformer decoder then attends to these representations at each decoding time-step, and the representation dimensionality of the decoder matches the output size of the projection layer. This also matches the embedding dimensionality for action vectors fed into the decoder. Positional embeddings are also used for the decoder input. The output of the decoder’s last layer conditions the CNN-based action predictor (replacing the GRU hidden state from the baseline model).
6.3.1 Experiments and Results
We conduct similar experiments with this model like with the GRU-based model, i.e., training on $D_{\mathrm{mc}}$ only, joint training on the aggregated dataset, and CL. The experimental setup and the CL experiments are same as before (Sections 6.1.1 and 6.2.2). We tested various sizes of BERT models (Turc et al., 2019), keeping BERT frozen, and found that $\text{BERT}_{\text{BASE}}$ (110.1M parameters) performed best. In the transformer decoder, five layers yielded optimal results, and a projection layer output size of 256 proved most effective. For tractability of experimentation, we only experimented with the number of transformer layers and the projection layer’s output size, keeping other architectural hyperparameters identical to the GRU-based models. Details specific to training regime, corresponding hyperparameters, etc. can be found in the appendix. Results are shown in Table 6. (Note: We also experimented with finetuning BERT, but found the frozen scenario to perform relatively better. However, as with the frozen case below, synthetic data and CL improved performance in the finetuning scenario as well, demonstrating the robustness of our approach. Results are provided in the appendix.)
Robustness of our approach
Overall trends align with those observed for the GRU-based models in Table 5. The synthetic data, joint training on the aggregated dataset, and CL, all also improve the BERT-based model’s performance on both synthetic and real data, highlighting the robustness of our approach. $M_{\mathrm{agg}}^{\prime}$ achieves an F1 score of 29.4% on $D_{\mathrm{mc}}$, up from $M_{\mathrm{mc}}^{\prime}$ ’s 23.0%, a notable 6.4-percentage point improvement. It is also highly performant across the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}^{\prime}$ by large margins on these. On $D_{\mathrm{mc}}$, CL again yields the best-performing model, as with the GRU-based models. $M_{\mathrm{agg+cl}}^{\prime}$ achieves an F1 of 32.3%, a notable 2.9-point improvement over $M_{\mathrm{agg}}^{\prime}$, and an even more significant 9.3-point improvement over $M_{\mathrm{mc}}^{\prime}$. It also achieves significant gains in performance over $M_{\mathrm{agg}}^{\prime}$ on the synthetic data as well.
Comparison to GRU-based models
The lower performance of the BERT-based models than their respective GRU-based counterparts’ is more pronounced for performance on the synthetic data than the Minecraft $D_{\mathrm{mc}}$ data. $M_{\mathrm{mc}}^{\prime}$ achieves an F1 of 23.0% on $D_{\mathrm{mc}}$, 4.3 points lower than $M_{\mathrm{mc}}$’s 27.3%. This gap between the GRU-based and BERT-based counterparts reduces significantly as we move on to look at $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg+cl}}^{\prime}$. It goes down to 0.9 and 0.7 points respectively. $M_{\mathrm{agg+cl}}^{\prime}$ is therefore not too far below $M_{\mathrm{agg+cl}}$. The gap in synthetic data performance between $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg}}$ is fairly significant. But CL reduces this quite a bit – $M_{\mathrm{agg+cl}}^{\prime}$ is very close to $M_{\mathrm{agg+cl}}$ on $D_{\mathrm{ss}}$, although still behind on the other two. Two key factors contribute to the relatively lower performance of the BERT-based models compared to their GRU counterparts:
  • As mentioned earlier, the sizes of all four datasets were optimized for the GRU-based models. These might need separate tuning for the BERT-based models (and will probably need to be scaled up also as transformers are more complex and data-hungry than GRUs).
  • For tractability of experimentation, we only experimented with the number of transformer layers and the projection layer’s output size, keeping other architectural hyperparameters identical to the GRU-based models.
These choices were intentional, as our goal here (as stated in the beginning of this section) was primarily to demonstrate the robustness of our synthetic data and training approach, and not to optimize dataset sizes or architectural hyperparameters for maximizing performance. We leave such fine-tuning for future work.
Model Train Data Test Data
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
$M_{\mathrm{mc}}^{\prime}$ $D_{\mathrm{mc}}$ 29.1 17.8 12.0 23.0
$M_{\mathrm{agg}}^{\prime}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 70.2 55.0 36.8 29.4
$M_{\mathrm{agg+cl}}^{\prime}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 75.9 67.4 45.3 32.3
Table 6: BERT- and transformer-based models
7 Model Analysis
We now analyze the performance and behavior of the best model $M_{\mathrm{agg+cl}}$ (Section 6.2.2) relative to the baseline $M_{\mathrm{mc}}$ (Section 2.6). In Section 7.1, we do so using the BAP v2 evaluation benchmark introduced in Section 4.4. Section 7.2 complements this with a qualitative evaluation of outputs from both models.
7.1 Quantitative evaluation
Dataset Model Type Color Location Shape Overall
EB Baseline ($M_{\mathrm{mc}}$) 76.4 75.2 53.5 56.0 53.2
Best ($M_{\mathrm{agg+cl}}$) 81.4 80.0 61.8 65.5 61.6
NEB Baseline ($M_{\mathrm{mc}}$) 63.8 60.8 24.6 33.8 24.2
Best ($M_{\mathrm{agg+cl}}$) 66.0 63.2 30.2 39.4 29.6
Overall Baseline ($M_{\mathrm{mc}}$) 65.2 62.4 27.7 36.3 27.3
Best ($M_{\mathrm{agg+cl}}$) 67.7 65.0 33.6 42.2 33.0
Table 7: Micro F1 scores for the baseline and best models under the v2 benchmark
Dataset Model Type Color Location Shape Overall
EB Baseline ($M_{\mathrm{mc}}$) 80.9 79.8 61.5 64.9 61.2
Best ($M_{\mathrm{agg+cl}}$) 83.1 81.7 69.0 72.6 68.6
NEB Baseline ($M_{\mathrm{mc}}$) 73.1 69.9 25.7 39.0 25.2
Best ($M_{\mathrm{agg+cl}}$) 73.7 71.2 32.1 46.5 31.5
Overall Baseline ($M_{\mathrm{mc}}$) 73.8 70.7 28.5 41.1 28.1
Best ($M_{\mathrm{agg+cl}}$) 74.5 72.0 35.0 48.6 34.5
Table 8: Macro F1 scores for the baseline and best models under the v2 benchmark
Tables 7 and 8 present all metrics obtained using the v2 benchmark, augmenting Tables 1 and 2 (Section 4.4) with results for the best model. Overall, the best model outperforms the baseline across all metrics. Key takeaways are discussed below.
We begin with the micro scores in Table 7. On the overall dataset, the best model outperforms the baseline on all three F1 metrics: Type, Color, and Location. As noted in Section 4.4, Location F1 is the primary bottleneck to achieving high performance in this task. The best model achieves an overall F1 of 33%, a notable 5.7 percentage-point improvement over the baseline’s 27.3%. This increase in overall F1 is primarily driven by the improvement in Location F1, which similarly rises from 27.7% to 33.6%. These results validate the motivation and impact of our synthetic data. The best model outperforms the baseline in both EB and NEB scenarios. However, as discussed in Section 4.4 regarding the importance of performance on EB, there remains room for improvement: the best model reaches only 61.6%, in spite of the increase from the baseline’s 53.2%. This insight can inform further error analysis and targeted improvements in future models for this task.
Turning to the macro scores in Table 8, we observe an increase across all metrics compared to the micro scores, but the same trends persist. The best model achieves an overall F1 of 34.5% on the overall test set, a notable 6.4 percentage-point improvement over the baseline’s 28.1%. For Shape F1 on the overall test set, the best model scores 42.2% (micro) and 48.6% (macro), up from the baseline’s 36.3% and 41.1%, marking significant gains of 5.9 and 7.5 percentage points, respectively.
7.2 Qualitative evaluation
We present examples of output from both the baseline and best models, demonstrating the substantial improvement of the latter. For each example, we also discuss the overall F1 metric and relevant auxiliary metrics. A complete table of scores for all metrics across examples is provided in the appendix (Table 10).
Figure 17: Example 1
Example 1
As shown in Figure 17, copying substructures is a common strategy in the MDC. Here, A instructs B to build an arrowhead, asking B to "mirror" the completed half on the opposite side. The baseline model only captures the "on the top" aspect of this instruction, while the best model fully understands and completes the arrowhead. This example also involves placing a floating block. The best model accomplishes this too, by placing and then removing a supporting block (not shown in the figure). It achieves an overall F1 score of 1.0, compared to the baseline model’s 0. However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need for a single purple block to be placed (in terms of net actions).
Figure 18: Example 2
Example 2
As shown in Figure 18, A gives a complex instruction requiring blocks to meet several highly specific criteria, including a criterion involving the builder’s position and orientation. The baseline model only grasps the "two other orange blocks" aspect but poorly interprets their placement. In contrast, the best model fully understands the instruction. Similar to the previous example, this example also involves placing floating blocks, which the best model accomplishes successfully. It achieves an overall F1 score of 1.0, compared to the baseline model’s 0. However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need for two orange blocks to be placed (in terms of net actions).
Figure 19: Example 3
Example 3
As shown in Figure 19, this example mirrors the previous one but focuses on block removals instead of placements. The initial structure, $S_{H}^{t-1}$, consists of three steps extending from a base square. A instructs B to remove a very specific block, using terms like "center" and "second highest step" to describe its location. The baseline model incorrectly selects the third highest step instead of the second, while the best model interprets the instruction accurately. The latter achieves an overall F1 score of 1.0, compared to the former’s 0. However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need to remove one green block.
8 Discussion
We now discuss four topics in detail: an additional key challenge in BAP evaluation and its implications for future work (Section 8.1), the importance of evaluation on the synthetic data (Section 8.2), relevant concurrent work (Section 8.3), and broader implications of our work that go beyond BAP (Section 8.4).
8.1 An additional key challenge for BAP evaluation
Through manual inspection of random samples during the qualitative analysis of model outputs (Section 7.2), we identified an additional challenge for fair evaluation in the BAP task, separate from the issue addressed by the fairer F1 metric in Section 4.1. Specifically, and rather unsurprisingly, the F1 metric lacks the sensitivity to distinguish nuanced differences in error quality. It is possible for two models to make different types of errors, with one type being qualitatively "better" than the other – yet the F1 metric may not capture this distinction effectively. This limitation can hinder fair evaluation and reduce our ability to distinguish between models based on their F1 scores. Below, we illustrate this challenge through qualitative examples. Across all examples, the best model $M_{agg+cl}$ consistently performs qualitatively better than the baseline $M_{mc}$ and is arguably closer to the human reference. However, the F1 metric lacks the sensitivity to reflect this substantial quality gap. In some cases, the auxiliary metrics proposed in Section 4.2 offer some help. A full table of scores for all metrics across examples is provided in the appendix (Table 11). It’s worth noting that such examples were frequent enough in our random inspection and the ones discussed here are fairly representative of the problem, not cherry-picked to favor the best model.
This analysis underscores the inherent challenges in automated evaluation for the BAP task, highlighting the need for further research in this area. The examples also illustrate the difficulty of creating a clear, unambiguous rubric for grading model output on an absolute scale, making human evaluation equally complex. (In fact, we initially considered human evaluation but ultimately dismissed it due to these challenges.)
Figure 20: Example A
Example A
As shown in Figure 20, A instructs B to place a yellow block meeting multiple highly specific criteria, including one involving the builder’s position and orientation. The best model largely understands these criteria and places the correct block, though it also adds an extraneous one on each side. In contrast, the baseline model performs poorly, producing seemingly random placements.
The F1 metric rates the baseline model as 0 and the best model as 0.5, indicating an improvement. However, this score does not fully capture the qualitative gap between the two models, especially in spatial reasoning. While both models produce two false positives, the errors made by the best model demonstrate a clearer understanding of the task, an aspect that the F1 metric fails to reflect adequately.
Figure 21: Example B
Example B
As shown in Figure 21, A and B are building a 2D giraffe, and A instructs B to place three orange blocks diagonally to form the neck, opposite the tail. The baseline model places three orange blocks, but arranges them horizontally and closer to the tail. While the best model recognizes the need for a diagonal structure, it places it in the wrong location and orientation.
Both models receive an F1 score of 0, with identical scores of 1, 1, and 0 for type, color, and location F1, respectively. However, shape F1 distinguishes the models more effectively: the baseline achieves a shape F1 of 0.33, while the best model achieves a perfect 1.
Figure 22: Example C
Example C
As shown in Figure 22, A instructs the addition of two rows of orange blocks directly below each hanging row on either side of the arch. The baseline model places two orange blocks on top, while the best model attaches blocks to each hanging row—still incorrect but a noticeably closer interpretation of A’s intent.
Here, too, the F1 metric scores both models as 0. Unlike the previous example, however, none of the auxiliary metrics distinguish between the models, as both receive identical scores.
Figure 23: Example D
Example D
As shown in Figure 23, A initially directs B to place a red block on the left, but then corrects this to the analogous position on the right, requiring removal of the original block and placement of a new one. The baseline model correctly removes the original red block but fails to place a new one, instead removing the top green block randomly. The best model performs only the correct removal and places a new red block to the right, though not in the exact position specified—nonetheless, it is clearly a closer interpretation of A’s intent.
In this case, the F1 metric assigns both models a non-zero score of 0.5, as both have 1 true positive, 1 false positive, and 1 false negative. However, the type and color F1 metrics penalize the baseline model, scoring it 0.5 in each, while the best model achieves a perfect score of 1.
8.2 Evaluation on synthetic data
Going forward, evaluation on the synthetic data will also be important for making more meaningful progress on the BAP task. First, the more structured and simpler nature of the synthetic data allows us to assess basic model competencies. Second, as noted in Section 6.1.2, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. Evaluating on the synthetic data will help build models that do not show such counterintuitive behavior (Section 8.3 provides further evidence supporting this argument), and complements evaluation on $D_{\mathrm{mc}}$. In fact, future models should aim to surpass previous performance on $D_{\mathrm{mc}}$ while also achieving strong results on synthetic data, ideally improving on both.
8.3 Concurrent work
Concurrent to our work are the recently published (at EMNLP, Nov 12-16, 2024) works of Chaturvedi, Thompson, and Asher (2024) and Kranti, Hakimov, and Schlangen (2024), both exploring SOTA LLMs for the BAP task. The former finetunes the Llama-3-8B model (Dubey et al., 2024) on the original BAP data, resulting in the Nebula model, which predicts the next action sequence in a text-to-text fashion using all preceding conversation and action sequences as context. The latter employs in-context learning with GPT-4. Nebula achieves an F1 score of 39.2% and GPT-4 scores 39.0% (under the now legacy BAP evaluation). Nebula is marginally better, although both are very close and mark improved performance on the task.
As mentioned in Section 1, modeling is not the primary focus of our work. Instead, we focus on the BAP task as a whole, addressing key challenges in data and evaluation to propose BAP v2, an upgraded version of the task itself. We also illustrate the impact of this data on training a simple LLM and transformer-based model. This foundation enables future work to make more efficient and meaningful progress on the task, and sets the stage for more sophisticated models/LLMs. These works are therefore complementary to ours. A slightly more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in the appendix, given some other relevant aspects of their work. We, still, however, evaluate Nebula using our new BAP v2 benchmark to further demonstrate the impact and robustness of our evaluation, and to provide a fairer assessment of its performance. We also evaluate it on our synthetic data to further strengthen our argument about the importance of such evaluation as discussed in Section 8.2. Future work can explore finetuning models such as Llama-3-8B (Nebula) and other SOTA LLMs using our synthetic data and training methodologies like CL.
BAP v2 evaluation
Similar to the 29.4% improvement we observed for our baseline model $M_{\mathrm{mc}}$ under the v2 evaluation on $D_{\mathrm{mc}}$, Nebula’s performance also increases from 39.2% to 50.1%, reflecting a 27.8% gain. This further underscores the impact and robustness of our v2 evaluation, offering a fairer reflection of Nebula’s performance. Detailed tables for the full v2 benchmark are provided in the appendix (Tables 12 and 13).
Evaluation on the synthetic data
Table 9 shows Nebula’s performance on both the synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and the original $D_{\mathrm{mc}}$ test sets. Evaluating Nebula on the synthetic data reveals behavior consistent with our baseline model $M_{\mathrm{mc}}$, as reported in Section 6.1.2 and further discussed in Section 8.2. It underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. This shows that models trained solely on $D_{\mathrm{mc}}$ may not be as robust, and reinforces our argument from Section 8.2 about the importance of synthetic data evaluation: it will help build models that don’t show such counter-intuitive behavior.
Model Test Data
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
Nebula 51.5 33.1 30.9 50.1
Table 9: Nebula’s performance across datasets (and using the v2 evaluation on $D_{\mathrm{mc}}$)
8.4 Broader implications of our work
While our focus in this work is on the Builder Action Prediction (BAP) subtask within the Minecraft Collaborative Building Task (MCBT) setting, the contributions have broader applications and takeaways as well. Below, we discuss their relevance for other subtasks within MCBT, for research building directly on the MCBT, and for similar domains and tasks in the field. This highlights the potential impact of our work and directions for future research.
Within MCBT
Although our primary focus is on BAP, the synthetic data and training methodologies introduced here are applicable to other MCBT subtasks constrained by the limited size of the MDC, such as the Architect Utterance Generation (AUG) subtask (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) summarized briefly in Section 2.1.2. AUG and BAP represent two key pillars for building fully interactive Architect and Builder agents. Additionally, other subtasks, such as deciding when the Builder should act versus speak or generating clarification questions during dialogue, could also benefit from our synthetic datasets. Our simulators and data were carefully designed to emulate the nuances of the MCBT itself (Section 5.1), making them directly usable or adaptable to these subtasks. The simulators’ parameterized nature enables generating tailored data distributions to this end, supporting further advancements across the broader MCBT ecosystem.
Beyond MCBT and Related Works
Potential applications of our work also go beyond the MCBT. Several works directly build upon or are inspired by the MCBT/MDC, including IGLU (Mohanty et al., 2024), Bonn et al. (2020), Thompson, Hunter, and Asher (2024), and Bonial et al. (2021). Therefore, future work on them can also potentially benefit from the synthetic data.
For Other Similar Domains and Tasks
Our findings also offer broader takeaways for research on tasks involving grounded, multimodal task-oriented dialogue systems operating in dynamic, spatially complex environments. First, in many realistic scenarios, data collection for complex situated dialogue tasks is often impractical or expensive. Our work illustrates that synthetic data generation can be a viable and scalable alternative to simply collecting more data. Second, we demonstrate that rich synthetic data can be generated for such tasks in game-based or simulated environments like Minecraft by carefully emulating task and game-specific nuances. This approach produces full embodied task oriented dialogues combining utterances, environment actions, and other game state information, offering a viable strategy for similar game-based/simulated and grounded dialog tasks –– an underexplored area as highlighted in Section 9.2. Finally, techniques like curriculum learning demonstrate potential and can be explored further in similar domains/tasks in combination with synthetic data.
9 Related Work
9.1 Task and Dataset
The MDC and BAP (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Jayannavar, Narayan-Chen, and Hockenmaier, 2020) continue to remain a challenging and relevant testbed for grounded language instruction following. We discuss a few more recent works to highlight this.
Vision-and-Language Navigation (VLN) (Anderson et al., 2018), and its dialog counterpart, Cooperative Vision-and-Dialog Navigation (CVDN) (Thomason et al., 2019), explore instruction following in photorealistic navigation settings. However, Minecraft dialogues are more complex due to their asynchronous communication, longer turn sequences, and dynamic environment changes during construction. Unlike navigation, where referring expressions target existing objects, construction often involves instructions for objects yet to be built. And although more recent navigation tasks require real vision, their underlying world state space (as defined by fixed viewpoints and the underlying navigation graph) is just as highly discretized. Our task does not require vision, but poses an arguably more challenging planning problem, since its action space is much larger (7623 possible actions vs. six actions in the vision-language navigation work). Tasks involving vision and language-based task completion beyond navigation have also been explored. ALFRED (Shridhar et al., 2020) is a benchmark for mapping natural language instructions and egocentric vision to action sequences for household tasks in simulated environments. However, similar to VLN and unlike the MDC, it lacks dialogue and relies solely on planner-based demonstrations instead of human ones. Efforts to combine vision and dialogue navigation with task completion include CEREALBAR (Suhr et al., 2019), which operates in a game-based setting. Yet, similar to CVDN, it uses turn-based dialogue, contrasting with the free-form, asynchronous communication of the MDC.
This limitation is addressed by the more recent benchmark TEACh (Padmakumar et al., 2022), which includes free-form dialogue, similar to the MDC, and features over 3000 human-human interactive dialogues for completing household tasks in simulation and with object state changes. However, unlike the MDC, the instruction giver in TEACh is provided with a predefined plan of steps, while the Architect in the MDC independently devises a unique plan for constructing the target structure. This results in greater creativity and diversity in the MDC, with multiple possible plans for the same structure. Although the TEACh dataset is larger (#dialogs), its dialogues and utterances are shorter. Additionally, the MDC allows for greater abstraction in instructions, such as referring to substructures, individual blocks, or shapes, whereas TEACh focuses on well-defined objects in the environment. Similar to VLN and CVDN, referring expressions in TEACh refer to objects that exist in the world, but construction instructions like those in the MDC frequently refer to objects that need to be built.
Minecraft, in recent years, has garnered a lot of interest as an AI experimentation platform. Our previous work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) was among the first to use it to study grounded task-oriented dialogue, and introduced the MCBT and MDC. Facebook’s CraftAssist (Gray et al., 2019; Jernite et al., 2019; Szlam et al., 2019) is another such example, enabling two-way human-bot interactions where a human architect instructs an automated builder to build complex structures. Their data includes synthetically generated and crowdsourced instructions paired with logical tree structures consisting of action primitives, unlike the MDC, which features human-human dialogues with greater ambiguity, variety, and noisier Builder actions. MineDojo (Fan et al., 2022) focuses on creating versatile agents for diverse tasks via an internet-scale knowledge base, contrasting with tasks like IGLU (Mohanty et al., 2024; Kiseleva et al., 2022b, a) and MCBT, which prioritize grounded natural language dialogue and clarification for interactive agents.
Among Minecraft-related efforts, IGLU is most closely related to the MDC/MCBT and was directly inspired by it. The IGLU dataset, unlike the MDC, is single turn only, and not dialogue based. It comprises 8136 instruction-action pairs, including 1056 ambiguous instructions with corresponding clarification questions. The task setting is similar to the MCBT but incorporates some simplifications which make it relatively less challenging. E.g., cardinal directions (e.g., North, South) are displayed in the voxel world, which simplifies instruction giving/following through the use of absolute rather than relative references (e.g., left, right). IGLU also introduces a very small multi-turn dataset (127 dialogues) based on this same task setting. It features strict turn-taking and clearly demarcated instructions and clarification questions. This contrasts with the MDC’s free-form, asynchronous dialogues, which are less structured and whose utterances have more diversity.
Some other works also build on or are inspired by the MCBT/MDC. For instance, Bonn et al. (2020), Thompson, Hunter, and Asher (2024), and Bonial et al. (2021) provide additional linguistic annotations for the MDC.
9.2 Synthetic Training Data
Synthetic data has proven beneficial for various tasks, including those mentioned in Section 9.1. VLN agents, constrained by limited human instruction data and diversity in training environments, often struggle with complex language grounding and spatial language understanding. Kamath et al. (2023) address this by using synthetic data, achieving SOTA performance on the RxR dataset (Ku et al., 2020), while Wang et al. (2023) are similar in spirit and achieve SOTA on the aforementioned CVDN and other benchmarks. Similarly, Kang et al. (2023) demonstrate strong performance gains in low-data regimes for Visual Dialog Das et al. (2017) by leveraging synthetic data. Synthetic data has also been effective for dialogue systems, particularly those focused on text-only dialogues (Kim et al., 2022; Bao et al., 2023; Zhan et al., 2023). Zhan et al. (2024) is an example of extending this to multi-modal dialogue agents, demonstrating the effectiveness of synthetic visual descriptions in enhancing agents’ grounding capabilities.
However, fewer efforts have addressed synthesizing task-oriented embodied dialogues (involving both utterances and environment actions). Padmakumar et al. (2023) were the first to design a framework to do so. They extend agenda-based dialogue simulation (Schatzmann and Young, 2009) to a multimodal embodied agent framework, and demonstrate the impact of the synthetic dialogues on the TEACh task (see Section 9.1 for a comparison of TEACh to MDC). Our work serves to provide another useful example in this general domain, further advancing synthetic data generation for such complex tasks. Dan, Han, and Roth (2021) is another closely related work that demonstrates the potential of synthetic data in the Blocks World domain. They use simpler synthetic data for the benchmarks of Bisk, Yuret, and Marcu (2016), which require understanding single-shot instructions that transform one world state to another using simulated 3D blocks. Blocks are viewed from a fixed bird’s eye perspective, initialized randomly in the initial world state, and uniquely identifiable. The varying Builder perspective and lack of easily identifiable referents, along with the need to understand utterances in a dialogue context, make the BAP task a much more challenging problem.  [^7] Also, unlike traditional Blocks World, Minecraft allows blocks to float (requiring non-monotonic action sequences where placement is followed by removal), or attach to any side of an existing block. Our work on synthetic data thus serves as a stronger demonstration of its potential for richer, embodied task-oriented dialogue tasks that extend beyond simpler Blocks World settings.
10 Conclusion and Future Work
In this work, we revisited the challenging Builder Action Prediction (BAP) subtask within the Minecraft Collaborative Building Task (MCBT) setting, addressing key shortcomings in evaluation and training data. We introduced BAP v2, an upgraded version of the task, comprising two core contributions: (1) an enhanced evaluation benchmark that includes a cleaner test set and fairer, more insightful metrics, and (2) diverse synthetic training datasets generated from novel Minecraft dialogue and target structure simulators. These contributions together provide a robust foundation for future research on the BAP task. Additionally, our synthetic data demonstrates the potential to train more performant and robust models, even with relatively simple training methods and architectures. Although modeling was not the primary focus of this work, we also illustrated the impact of the synthetic data and training methodologies on training a simple LLM and transformer-based model, thus validating the robustness of our approach, and setting the stage for more sophisticated models/LLMs going forward.
Future work can therefore build on our work to explore more advanced LLMs and architectures while leveraging the BAP v2 framework, alongside training methodologies such as Curriculum Learning (CL). Further improvements could stem from experimenting with more sophisticated CL approaches or entirely new training paradigms to get even more out of the synthetic data. Additionally, as highlighted in Section 8.1, BAP evaluation remains a nuanced challenge, warranting deeper investigation. Finally, Section 8.4 outlined the broader implications of our work, while also highlighting potential directions for future research.
11 Appendix: BAP v2 metrics for Sections 7.2 and 8.1
Example ID Model Type Color Location Shape Overall
Example 1 Baseline 1.0 1.0 0.0 0.0 0.0
Best 1.0 1.0 1.0 1.0 1.0
Example 2 Baseline 1.0 1.0 0.0 0.5 0.0
Best 1.0 1.0 1.0 1.0 1.0
Example 3 Baseline 1.0 1.0 0.0 0.0 0.0
Best 1.0 1.0 1.0 1.0 1.0
Table 10: BAP v2 metrics by Example ID for Section 7.2 (range $[0,1]$)
Example ID Model Type Color Location Shape Overall
Example A Baseline 0.67 0.67 0.0 0.0 0.0
Best 0.5 0.5 0.5 0.5 0.5
Example B Baseline 1.0 1.0 0.0 0.33 0.0
Best 1.0 1.0 0.0 1.0 0.0
Example C Baseline 0.5 0.5 0.0 0.0 0.0
Best 0.5 0.5 0.0 0.0 0.0
Example D Baseline 0.5 0.5 0.5 0.0 0.5
Best 1.0 1.0 0.5 0.0 0.5
Table 11: BAP v2 metrics by Example ID for Section 8.1 (range $[0,1]$)
12 Appendix: Concurrent work
Dataset Type Color Location Shape Overall
EB 92.8 92.8 78.3 84.8 78.3
NEB 79.0 77.4 47.2 59.0 46.4
Overall 80.6 79.1 50.7 61.9 50.1
Table 12: Micro F1 scores for the Nebula model under the v2 benchmark
Dataset Type Color Location Shape Overall
EB 96.8 96.8 84.5 90.5 84.5
NEB 81.7 80.9 47.1 64.1 46.5
Overall 82.9 82.2 50.1 66.3 49.6
Table 13: Macro F1 scores for the Nebula model under the v2 benchmark
[…]

### v2

@@
 Although modeling is not the primary focus of this work, we also illustrate the impact of our data and training methodologies on a simple LLM- and transformer-based model, thus validating the robustness of our approach, and setting the stage for more advanced architectures and LLMs going forward.
-issue: 1
 1 Introduction
@@
 The latter defined the challenging Builder Action Prediction (BAP) task, aimed at predicting the sequence of actions (block placements and/or removals) a human Builder performed at a specific point in a human-human game.
-Despite the task’s inherent challenges and limited training data, we demonstrate that models trained with a suitable amount of game history, enriched world representations, and a sufficient amount of synthetic, diversified data (via simple data augmentation techniques) yield promising initial results.
+Despite the task’s inherent challenges and limited training data, we demonstrate that models trained with a suitable amount of game history, enriched world representations, and a sufficient amount of synthetic, diversified data (via data augmentation techniques) yield promising initial results.
 We recap relevant prior work on the Minecraft Builder in Section 2.
@@
 Instead of merely collecting more data (which can be impractical/expensive in realistic scenarios for such complex situated dialogue tasks), we create additional diverse synthetic training data generated from novel Minecraft dialogue and target structure simulators that emulate the MCBT (Section 5).
-This data, though naturally simpler than it, is rich in spatial relations and referential expressions and we show that it can be successfully used to train better models, even with simple training methods (Section 6).
+This data, though naturally simpler than it, is rich in spatial relations and referential expressions and we show that it can be successfully used to train better models, even with straightforward training methods (Section 6).
 We demonstrate that jointly training on combined synthetic and original BAP (augmented) data already improves performance.
@@
 We conclude and discuss future work in Section 10.
+We plan to release code and data involved in this work.
 2 Prior Work on the Minecraft Builder
@@
 Although both of these tasks consider only a single turn, and assume the context of a game and dialogue between two human players, they are important first steps towards the creation of agents that can complete an entire game in the MCBT with a human counterpart.
+We note that compared to BAP, Shi, Feng, and Lipani (2022) propose a slightly different building task involving predicting a single action instead of a sequence of actions.
+Thus, their setup is not directly comparable to ours.
 2.2 The Builder Action Prediction (BAP) Task
@@
 2.5 Data Augmentation
-Since the small size of the training set (3,709 examples) is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B) (details in appendix).
+Since the small size of the training set (3,709 examples) is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B).
 Empirically, we found that increasing the training data to 14,836 (4x) items gave the best performance for our baseline model (Section 2.6).
@@
 This yields an F1 score of 21.2% on the BAP test set.
-More details about the model are provided in the appendix.
 3 Revisiting BAP Evaluation: A Cleaner Test Set
@@
 If the context is clear (with unique or multiple interpretations), a second question asked whether the built structure is correct or incorrect.
-Details about the annotation interface, the annotation workflow, the annotation tool, data collection setup, etc.
-are provided in the appendix.
+Details about the annotation setup are provided in the appendix (Section 14).
 We now illustrate the resulting categories of BAP items with a few examples.
@@
     Optional Clarification Exchange: Some information may be omitted from the instruction with a small probability, allowing B to ask a clarification question.
-An appropriate exchange is then generated, with B asking and A providing the missing information.
+An appropriate exchange is then generated, with B asking and A providing the missing information, with template-based lexical and syntactic variations.
 E.g., in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
@@
 Each simulator uniquely instantiates steps 1, 3, and 4, along with independent target structure generation, while steps 2, 5, and 6 remain largely consistent across simulators.
-For brevity, we provide only a summary of each simulator in the following sections, with detailed descriptions in the appendix.
+For brevity, we provide only a summary of each simulator in the following sections and provide more details in the appendix (Section 11).
 5.1.3 Spatial Relations
@@
 Dialogs
-The dialogs correspond to random target structures generated dynamically during the dialog simulation (Section 5.1.2) which runs for a maximum number of iterations.
+The dialogs correspond to random target structures generated dynamically during the dialog simulation (Section 5.1.2) which runs for a (bounded) random number of iterations.
 Blocks are placed/removed randomly, and one at a time in each iteration, with limited planning on A ’s part for block placement.
@@
 A identifies the orange block as a reference block and uses a 1D spatial relation "behind" to specify the target yellow block.
+We provide more details about the simulator in the appendix (Section 11).
 Figure 12: A BAP item from the random structure-based dialogs
@@
 5.3.1 Shape-based structures
-We define six unique elementary 3D shapes: rows, diagonals, T-shapes, L-shapes, U-shapes, and planes.
+We define six unique elementary shapes: rows, diagonals, T-shapes, L-shapes, U-shapes, and planes.
 These shapes were chosen for their simplicity and occurrence in MDC dialogs and structures.
-Formal definitions of these shapes are provided in the appendix.
+(Formal definitions of the shapes are provided in the appendix (Section 11.4)).
 We design a simulator that generates target structures by combining multiple randomly sampled elementary shapes into a single composite structure.
@@
 A is instructing B to continue building the latter by placing three more blocks, using a 1D spatial relation "on top of" relative to the last block.
-(In the previous step, A used a 2D spatial relation, "diagonally in front of and to the right of".)
+(In the previous step, A used a 2D spatial relation, "diagonally in front of and to the right of".) We provide more details about the simulator in the appendix (Section 11).
 Figure 15: A BAP item from the blocks-based dialogs for shape-based structures
@@
 Additionally, A can also use spatial relations that use B as a spatial anchor (Section 5.1.3), for building either shape.
+B’s low-level planning also uses heuristics to determine the order of blocks within a shape, mimicking human Builders.
+Multiple orderings are possible, with some being more optimal or natural than others.
+E.g., 2D shapes like planes can be built either by column or by row.
+Additionally, B may also adopt a zigzag pattern, alternating the starting side for each row or column.
 Figure 16 illustrates a BAP item from this dataset.
@@
 A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row, along with the 3D spatial relation "one to the left, two underneath, and one block in front of" relative to the last block (the topmost yellow one), to indicate the starting position for the row.
-Figure 16: A BAP item from the shapes-based dialogs for shape-based structures
+We provide more details about the simulator in the appendix (Section 11).
+Figure 16: A BAP item from the shape-based dialogs for shape-based structures
 6 Training with Synthetic Data
-Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with simple training methods, thereby addressing the challenge of insufficient training data.
-We first show that simple joint training on the aggregation of the synthetic data and the original BAP (augmented) data already yields a better model (Section 6.1).
+Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with straightforward training methods, thereby addressing the challenge of insufficient training data.
+We first show that just joint training on the aggregation of the synthetic data and the original BAP (augmented) data already yields a better model (Section 6.1).
 We then go a step further and show that a simple Curriculum Learning method can further boost performance (Section 6.2).
@@
 6.1.1 Experimental Setup
-Our setup closely follows that of Section 2.6, where the baseline $M_{\mathrm{mc}}$ was trained on $D_{\mathrm{mc}}$ alone.
+Our setup closely follows the one used for the baseline model $M_{\mathrm{mc}}$ when it was trained on $D_{\mathrm{mc}}$ alone (Section 2.6).
 We retain the same exact model architecture and hyperparameters as in $M_{\mathrm{mc}}$.
 Training employs teacher forcing with cross-entropy loss, and decoding is performed via greedy decoding, with a maximum sequence length of 20 actions.
+Further details are provided in the appendix (Section 12.3).
 We use the fairer F1 metric to report micro F1 scores on the BAP v2 test set for $D_{\mathrm{mc}}$ (Section 4.4) and the test sets of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$.
@@
 We adopt an empirical yet systematic approach to address this, yielding a data mix that optimizes $D_{\mathrm{mc}}$ performance while also being highly performant on the synthetic data as much as possible.
-Details can be found in the appendix.
+More details are provided in the appendix (Section 12.2).
 Table 3 presents the final data statistics.
@@
 For a sense of what high performance means on these datasets, we reference the diagonal entries in rows for models $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, representing performance of these models on the same test data distribution as train: 83.9% for $D_{\mathrm{bs}}$, 75.4% for $D_{\mathrm{ss}}$, and 64.4% for $D_{\mathrm{r}}$.
-In light of these numbers, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
+In light of these numbers, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
 In contrast, $M_{\mathrm{agg}}$ is more robust, and retains much of the diagonal-entry performance on the synthetic datasets while also significantly improving on $D_{\mathrm{mc}}$.
@@
 Another key observation from Table 4 is that performance of the $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, and $M_{\mathrm{mc}}$ models decreases in the order $M_{\mathrm{bs}}$ > $M_{\mathrm{ss}}$ > $M_{\mathrm{r}}$ > $M_{\mathrm{mc}}$, suggesting that the complexity/difficulty of examples increases in the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$, with the most complex $D_{\mathrm{mc}}$ examples naturally appearing at the end.
-This observation raises the possibility of using this ordering to improve over the simple joint training approach (where data is randomly ordered) when training on the aggregated dataset, so as to achieve even better performance on the complex $D_{\mathrm{mc}}$.
+This observation raises the possibility of using this ordering to improve over the naive joint training approach (where data is randomly ordered) when training on the aggregated dataset, so as to achieve even better performance on the complex $D_{\mathrm{mc}}$.
 Curriculum Learning (CL) is a natural next step to explore in this direction.
@@
 A key prerequisite in CL is establishing difficulty metric(s) that quantify how difficult a training example is for a model to learn, allowing the data to be sorted accordingly during training.
-In this work, we explore three simple, natural, and intuitive difficulty metrics, along with their combinations (e.g., sorting first by one metric and then sorting (breaking ties) by another).
+In this work, we explore three straightforward, natural, and intuitive difficulty metrics, along with their combinations (e.g., sorting first by one metric and then sorting (breaking ties) by another).
 The first metric is the dataset $D$ an example belongs to ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$).
@@
 CL Algorithm
-We make use of a simple CL algorithm, specifically using the Curriculum Arrangement algorithm within the CL framework proposed by Xu et al. (2020). Details of the method are provided in the appendix.
+We make use of a simple CL algorithm, specifically using the Curriculum Arrangement algorithm within the CL framework proposed by Xu et al.
+(2020).
+Details of the method are provided in the appendix (Section 12.1).
 6.2.2 Experiments and Results
-Overall experimental setup remains similar to the batch training one (Section 6.1.1).
+Overall experimental setup remains similar to the one used for joint training (Section 6.1.1).
 We experiment with each of the three difficulty metrics individually and the four possible combinations of them, making a total of seven metrics – $D$, $L_{A}$, $L_{I}$, $(D,L_{A})$, $(D,L_{I})$, $(L_{I},L_{A})$, and $(D,L_{I},L_{A})$ (ordering within a tuple denotes the sorting order).
-We also experiment with the CL algorithm’s hyperparameter, $N$, trying six values – 5, 10, 15, 20, 25, and 30.
-(This controls the number of buckets/difficulty levels – see appendix.) This results in a grid of $7\times 6=42$ experiments.
+We also experiment with the CL algorithm’s parameter, $N$, trying six values – 5, 10, 15, 20, 25, and 30.
+(This controls the number of buckets/difficulty levels – see appendix (Section 12.1).) This results in a grid of $7\times 6=42$ experiments.
 Table 5 presents the best model, $M_{\mathrm{agg+cl}}$, selected based on $D_{\mathrm{mc}}$ performance.
@@
 It also continues to be highly performant on the synthetic data similar to $M_{\mathrm{agg}}$.
-Thus, Curriculum Learning on the aggregated data provides a meaningful performance boost on the BAP task over simple joint training.
+Thus, Curriculum Learning on the aggregated data provides a meaningful performance boost on the BAP task over naive joint training.
 Hereafter, we refer to $M_{\mathrm{agg+cl}}$ as the best model.
@@
 Model details
-To control the representation dimensionality and complexity of the decoder transformer, we add a simple feedforward projection layer atop BERT, compressing its output (as otherwise the dimensionality of the decoder transformer would be forced to be the same as the BERT encoder and make it overly complex than need be).
+To control the representation dimensionality and complexity of the decoder transformer, we add a feedforward projection layer atop BERT, compressing its output (as otherwise the dimensionality of the decoder transformer would be forced to be the same as the BERT encoder and make it overly complex than need be).
 The BERT encoder and transformer decoder are connected in standard encoder-decoder fashion but via the projection layer.
@@
 We conduct similar experiments with this model like with the GRU-based model, i.e., training on $D_{\mathrm{mc}}$ only, joint training on the aggregated dataset, and CL.
-The experimental setup and the CL experiments are same as before (Sections 6.1.1 and 6.2.2).
-We tested various sizes of BERT models (Turc et al., 2019), keeping BERT frozen, and found that $\text{BERT}_{\text{BASE}}$ (110.1M parameters) performed best.
+The experimental setup and the CL experiments are the same as for the GRU scenario (Sections 6.1.1 and 6.2.2).
+Further details specific to the BERT scenario are provided in the appendix (Section 12.4).
+We tested various sizes of BERT models (Turc et al., 2019), and keeping BERT frozen versus finetuning it.
+We found that frozen $\text{BERT}_{\text{BASE}}$ (110.1M parameters) performed the best.
 In the transformer decoder, five layers yielded optimal results, and a projection layer output size of 256 proved most effective.
-For tractability of experimentation, we only experimented with the number of transformer layers and the projection layer’s output size, keeping other architectural hyperparameters identical to the GRU-based models.
-Details specific to training regime, corresponding hyperparameters, etc.
-can be found in the appendix.
+Other hyperparameters are provided in the appendix (Section 12.4).
+For tractability of experimentation, we only experimented with the transformer’s and projection layer’s hyperparameters, keeping other architectural hyperparameters identical to the GRU-based models (e.g.
+those of the CNNs).
 Results are shown in Table 6.
-(Note: We also experimented with finetuning BERT, but found the frozen scenario to perform relatively better.
-However, as with the frozen case below, synthetic data and CL improved performance in the finetuning scenario as well, demonstrating the robustness of our approach.
-Results are provided in the appendix.)
+(Note: We report only the best-performing BERT-based model, which was from the frozen scenario as mentioned earlier.
+However, similar to the results for the frozen case discussed below, synthetic data, joint training, and CL also improved model performance in the finetuning scenario, validating our approach.)
 Robustness of our approach
@@
 These might need separate tuning for the BERT-based models (and will probably need to be scaled up also as transformers are more complex and data-hungry than GRUs).
-  • For tractability of experimentation, we only experimented with the number of transformer layers and the projection layer’s output size, keeping other architectural hyperparameters identical to the GRU-based models.
+  • For tractability of experimentation, we only experimented with the transformer’s and projection layer’s hyperparameters, keeping other architectural hyperparameters identical to the GRU-based models (e.g.
+those of the CNNs).
 These choices were intentional, as our goal here (as stated in the beginning of this section) was primarily to demonstrate the robustness of our synthetic data and training approach, and not to optimize dataset sizes or architectural hyperparameters for maximizing performance.
@@
 First, the more structured and simpler nature of the synthetic data allows us to assess basic model competencies.
-Second, as noted in Section 6.1.2, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
+Second, as noted in Section 6.1.2, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
 Evaluating on the synthetic data will help build models that do not show such counterintuitive behavior (Section 8.3 provides further evidence supporting this argument), and complements evaluation on $D_{\mathrm{mc}}$.
@@
 These works are therefore complementary to ours.
-A slightly more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in the appendix, given some other relevant aspects of their work.
-We, still, however, evaluate Nebula using our new BAP v2 benchmark to further demonstrate the impact and robustness of our evaluation, and to provide a fairer assessment of its performance.
-We also evaluate it on our synthetic data to further strengthen our argument about the importance of such evaluation as discussed in Section 8.2.
-Future work can explore finetuning models such as Llama-3-8B (Nebula) and other SOTA LLMs using our synthetic data and training methodologies like CL.
+A more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in the appendix (Section 15), given some other relevant aspects of their work.
+Here, we evaluate Nebula using the v2 benchmark to further demonstrate the impact and robustness of our evaluation while providing a fairer assessment of its performance.
+We observe a 27.8% gain in performance (compared to the legacy evaluation), similar to the gain for our baseline $M_{\mathrm{mc}}$.
+We also discuss how the v2 evaluation now enables more accurate benchmarking of SOTA LLMs on the BAP task and highlights the significant room for improvement, further motivating the task itself.
+Additionally, it helps us confirm that spatial reasoning remains the key bottleneck for high performance, motivating future efforts leveraging our data and methods, which have shown effectiveness in addressing this challenge.
+We also evaluate Nebula on our synthetic data and find that, like the baseline $M_{\mathrm{mc}}$ (and the BERT-based $M_{\mathrm{mc}}^{\prime}$), it underperforms significantly despite being trained on the far more complex $D_{\mathrm{mc}}$.
+This suggests that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust, reinforcing our argument of Section 8.2 about the importance of the synthetic data evaluation.
+Future work should explore finetuning models such as Llama-3-8B (Nebula) and newer LLMs using our synthetic data and training methodologies like CL.
 BAP v2 evaluation
@@
 This further underscores the impact and robustness of our v2 evaluation, offering a fairer reflection of Nebula’s performance.
-Detailed tables for the full v2 benchmark are provided in the appendix (Tables 12 and 13).
+With this fairer evaluation now possible, we can more accurately gauge current performance on the BAP task and the remaining gap for improvement.
+While the 50.1% score marks improved performance, it also underscores that even with SOTA LLMs, significant room for improvement remains, thus reinforcing the task’s challenging nature and its importance for future research directions.
+Detailed tables for the full v2 benchmark are provided in the appendix (Tables 13 and 14).
+The observed trends generally align with those seen in our models such as the model’s relative strengths and weaknesses.
+They also shed some light on key areas for future improvement.
+For example, spatial reasoning (Location F1) is the key bottleneck for high performance on this task as noted in Section 4.4, and this continues to remain the case.
+Therefore, as demonstrated with $M_{\mathrm{mc}}$, our synthetic data and training methodologies can potentially help address this challenge, motivating future work.
 Evaluation on the synthetic data
-Table 9 shows Nebula’s performance on both the synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and the original $D_{\mathrm{mc}}$ test sets.
-Evaluating Nebula on the synthetic data reveals behavior consistent with our baseline model $M_{\mathrm{mc}}$, as reported in Section 6.1.2 and further discussed in Section 8.2.
-It underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
-This shows that models trained solely on $D_{\mathrm{mc}}$ may not be as robust, and reinforces our argument from Section 8.2 about the importance of synthetic data evaluation: it will help build models that don’t show such counter-intuitive behavior.
+Table 9 shows Nebula’s performance (micro F1) on both the synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and the original $D_{\mathrm{mc}}$ test sets.
+Evaluating Nebula on the synthetic data reveals behavior consistent with our baseline model $M_{\mathrm{mc}}$ (and the BERT-based $M_{\mathrm{mc}}^{\prime}$), as reported in Section 6.1.2 and further discussed in Section 8.2 – it underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
+This shows that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust, and reinforces our argument from Section 8.2 about the importance of the synthetic data evaluation: it will help build models that don’t show such counter-intuitive behavior.
 Model Test Data
@@
 9.1 Task and Dataset
-The MDC and BAP (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Jayannavar, Narayan-Chen, and Hockenmaier, 2020) continue to remain a challenging and relevant testbed for grounded language instruction following.
-We discuss a few more recent works to highlight this.
+We discuss related work to illustrate that the MDC and BAP remain a challenging and relevant testbed for grounded language instruction following.
 Vision-and-Language Navigation (VLN) (Anderson et al., 2018), and its dialog counterpart, Cooperative Vision-and-Dialog Navigation (CVDN) (Thomason et al., 2019), explore instruction following in photorealistic navigation settings.
@@
 This results in greater creativity and diversity in the MDC, with multiple possible plans for the same structure.
-Although the TEACh dataset is larger (#dialogs), its dialogues and utterances are shorter.
+Although the TEACh dataset is larger (by number of #dialogs), its dialogues and utterances are shorter.
 Additionally, the MDC allows for greater abstraction in instructions, such as referring to substructures, individual blocks, or shapes, whereas TEACh focuses on well-defined objects in the environment.
 Similar to VLN and CVDN, referring expressions in TEACh refer to objects that exist in the world, but construction instructions like those in the MDC frequently refer to objects that need to be built.
-Minecraft, in recent years, has garnered a lot of interest as an AI experimentation platform.
+Minecraft has garnered a lot of interest as an AI experimentation platform in recent years.
 Our previous work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) was among the first to use it to study grounded task-oriented dialogue, and introduced the MCBT and MDC.
@@
 [^7] Also, unlike traditional Blocks World, Minecraft allows blocks to float (requiring non-monotonic action sequences where placement is followed by removal), or attach to any side of an existing block.
-Our work on synthetic data thus serves as a stronger demonstration of its potential for richer, embodied task-oriented dialogue tasks that extend beyond simpler Blocks World settings.
+Our work on synthetic data thus serves as a demonstration of the potential for richer, embodied task-oriented dialogue tasks that extend beyond simpler Blocks World settings.
 10 Conclusion and Future Work
@@
 Finally, Section 8.4 outlined the broader implications of our work, while also highlighting potential directions for future research.
-11 Appendix: BAP v2 metrics for Sections 7.2 and 8.1
+11 Appendix: Generating Synthetic Data
+Sections 11.111.2, and 11.3 provide additional details (beyond those covered in Section 5) on how each of the three dialog simulators implements the general simulation framework (Section 5.1.2) – specifically, steps 1 (Architect Planning), 3 (Architect Instruction), and 4 (Optional Clarification Exchange).
+For details on the elementary shapes used for generating shape-based target structures (Section 5.3.1) see Section 11.4.
+Note
+All distances are measured using Manhattan Distance.
+For Steps 3 and 4, we outline only the key elements of our template-based generation framework.
+Further details will be included in our eventual code and data release.
+Thus, we omit detailed lexical and syntactic variations of utterances, providing only a few representative examples.
+11.1 Random target structures and dialogs
+Architect Planning
+Next Action:
+The first four actions in the dialog are placements.
+Subsequent actions are sampled with a 90% probability for placement and 10% for removal.
+For placements, the algorithm selects a color uniformly randomly from the six colors and a location from a set of candidate positions.
+For removals, a block from the built structure is chosen at random.
+Candidate positions are defined as follows.
+If the board is empty (start of dialog), candidate positions are limited to the ground.
+But if at least one block exists, a candidate position must be within the "connected neighborhood" of an existing block.
+The "connected neighborhood" includes all locations within a Manhattan distance of 2 that share a face or edge with an existing block (but not a sole corner).
+This forms a 3x3x3 cube around the block, excluding the 8 diagonal corners but including the 6 adjacent and 12 2D diagonal locations.
+Reference Block:
+If needed, the algorithm selects a reference block as the closest existing block to the chosen next block.
+However, if the block from the last action that took place has the same distance, it takes precedence.
+A reference block can either be an existing block in the built structure (the last added one or a randomly selected one) or the last block that was removed from it.
+There are two special cases where a reference block is not required or cannot be assigned.
+These are discussed later in Section Special Cases with No Reference Block.
+Architect Instruction
+Next Action
+The instruction specifies the color of the block, and when necessary, includes an optional adjective such as "floating" if temporary supporting blocks are required.
+For example, in Figure 11, "blue" and "floating" are used, while in Figure 12, "purple" and "yellow" are used.
+Reference Block
+A phrase is needed to refer to the reference block, determined by the following priority rules (from highest to lowest):
+  1. 1.
+    If the reference block is the last placed/removed block, the phrase used is "the last block you added/removed" etc.
+If the last action was a placement, the algorithm may omit the reference block in the instruction (ellipsis) with a small probability.
+  2. 2.
+    If the reference block’s color is unique in the built structure, it is referred to as "the [COLOR] block," where [COLOR] represents the block’s color.
+  3. 3.
+    If the color is not unique, the block is disambiguated using its spatial position relative to other blocks of the same color, allowing references such as "the leftmost [COLOR] block," etc.
+For example, in Figure 11, the reference block is described as "the last block" and "the last block you placed." In Figure 12, it is referred to as "the closest yellow block" and "the orange one."
+Optional Clarification Exchange
+The elements of the instruction that may be omitted include the location (i.e., the spatial relation + reference block) or the color of the next action.
+B’s clarification questions, such as "Where?" or "What color?" are then generated accordingly, followed by A’s response.
+For example, in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
+Special Cases with No Reference Block
+There are two special cases where no reference block (and consequently no spatial relation involving it) is used:
+Removal of the Last Placed Block
+If the next action is a removal and the block to be removed is the last placed block, the following adjustments are made while keeping everything else unchanged: 1) No reference block is needed.
+2) A’s instruction will be a simple command such as "remove that block".
+3) No clarification exchanges are generated, as there is nothing to omit in the instruction regarding the location or color of the block to be removed.
+Start of Dialog
+At the start of the dialog (empty board), the following modifications are made: 1.
+The first action is the of a block on the ground (randomly sampled from target structure blocks on the ground).
+2.
+Since there are no prior blocks, A’s instruction will be a simple directive, such as "First start by placing an orange block on the ground".
+11.2 Blocks-based dialogs for shape-based structures
+Architect Planning
+Recall that a target structure is provided as input to the dialog simulation.
+In this iteration of the simulation, blocks to be placed next are selected from those that remain unplaced, given the current built structure and the target.
+The algorithm first selects a single block to place.
+Based on this block and the reference block, it determines whether multiple blocks can be placed simultaneously while maintaining the same spatial relation.
+(We give an example of this later.) If so, multiple blocks are added; otherwise, only a single block is placed.
+Next Block
+When deciding the next block to place, the overall goal is to ensure a natural (human-like), ordered, and simple building process.
+More specifically, the following principles help guide this decision.
+The structure should be constructed shape by shape, and it should be ensured that the next block is connected to the existing built structure.
+Out of the possible candidate next blocks, those with the same color and within the connected neighborhood (defined in Section Next Action:) of the last action should be given higher priority.
+The building process should also aim to minimize branching within the structure (i.e., building as linearly as possible) so as to reduce the number of instructions and spatial relations that A needs to provide.
+Lastly, temporary supporting blocks should be used sparingly and only when necessary.
+We implement these principles via a few heuristics.
+Reference Block
+The approach is similar to the one in Section 11.1, with one special case where a reference block cannot be assigned (see Section Special Cases with No Reference Block).
+Multiple Blocks (if any)
+Next, we check whether multiple blocks (i.e., apart from the next block) can be placed "greedily" using the same reference block and spatial relation.
+For instance, if the next block is a yellow block placed to the left of the reference block, and another yellow block exists further to the left, A may plan to place both simultaneously for efficiency.
+To this end, the algorithm computes the direction of the next block relative to the reference block and checks if additional unplaced blocks of the same color exist along that direction.
+It then selects the longest contiguous sequence of such blocks connected to the next block.
+Architect Instruction
+The approach is similar to the one in Section 11.1 but additionally accounts for the number of blocks when multiple blocks are placed.
+For example, in Figure 16, the instruction specifies "three yellow blocks."
+Optional Clarification Exchange
+The approach is similar to the one in Section 11.1.
+Special Cases with No Reference Block
+There is only one special case, which is the same as the "Start of Dialog" one described in Section Special Cases with No Reference Block.
+11.3 Shape-based dialogs for shape-based structures
+Architect Planning
+Recall that a target structure consisting of two shape instances is provided as input to the dialog simulation.
+Next Shape and Its Starting Block
+Shapes are built sequentially.
+If both shape instances are on the ground, one is chosen randomly as the first to be built, otherwise the algorithm starts with the shape on the ground.
+Shapes are naturally built from the bottom up.
+The algorithm first selects a block within the shape to define spatial relations needed to specify where this shape should be built (e.g.
+wrt the reference block).
+This block also serves as the starting point for B when building the shape (the order of the remaining blocks is decided by B’s planning in Step 5 of the simulation algorithm of Section 5.1.2).
+Potential starting blocks are always located in the bottom corners of the shape.
+For the first shape, the starting block is the one farthest from the second shape.
+The starting block for the second shape is the block with the shortest distance to the last placed block in the first shape.
+Reference Block
+For the second shape being placed, the reference block is the last placed block of the first shape. No reference block is used for the first shape.
+Architect Instruction
+Next Shape
+A shape’s size can be described in multiple ways depending on its dimensions.
+For 1D shapes such as rows, the description includes a single number, such as "four blocks long" or "three blocks tall." For 2D shapes like diagonals and planes, various formats can be used, such as "3x3", "two blocks long and three tall", etc.
+For example, in Figure 16, the instructions specify "4 yellow blocks" and "five block long."
+The instruction also includes the shape’s color and name.
+We provide examples here for three shapes – rows, diagonals and planes.
+The primary shape names are "row," "diagonal," and "plane," though synonyms can be used.
+A horizontal row may be called a "line," while a vertical row may be referred to as a "column" or "tower." Similarly, a horizontal plane may be described as a "layer," whereas a vertical plane may be called a "wall." A horizontal diagonal is typically referred to as a "diagonal" or "diagonal line," but a vertical diagonal can also be described as a "staircase" or "stairway." For example, in Figure 16, the terms "pillar" and "line" are used for shape names, while "yellow" and "red" are used for shape colors.
+Reference Block
+This applies only to the second shape being placed.
+Since the reference block is the last placed block, we use phrases such as "the last block you placed," similar to the approach in Sections 11.1 and 11.2.
+Optional Clarification Exchange
+The approach is similar to the one in Sections 11.1 and 11.2, with two additional elements that can be omitted in an instruction.
+Recall that this data also includes spatial relations that use B as a spatial anchor (Section 5.1.3).
+For example, in Figure 16, A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row.
+Such directional spatial relations may be omitted from the instruction.
+Additionally, the dimensions/size of a shape may also be omitted.
+11.4 Shape-based target structures
+The six elementary shapes used to generate the structures are defined as follows:
+Rows
+consist of at least 3 blocks, and can point in three different directions (X, Y and Z), aligning to each of the $x$, $y$, and $z$ axes. Those that align to the $y$ axis are columns.
+Diagonals
+consist of at least 3 blocks, and are planar (along any of the $xy$, $yz$, or $xz$ planes). Horizontal diagonals lie along the $xz$ plane, vertical along either $xy$ or $yz$.
+T-shapes
+($\geq 3\times 3$ blocks) are composed of two orthogonal rows, in which one end of one row intersects the midpoint of another.
+The orientations of these rows determine the orientation of the T.
+Horizontal T-shapes lie completely along the $xz$ plane, while vertical ones span either $xy$ or $yz$.
+Additionally, vertical T-shapes can be up (facing right-side up) or down (upside-down).
+L-shapes
+($\geq 2\times 2$ blocks) are similar to T-shapes in construction in that they consist of two orthogonal rows attached at their ends.
+Orientations are similar to those of T-shapes, including horizontal L-shapes as well as vertical ones that can point either up or down.
+U-shapes
+are symmetrical structures built out of rows: one row (with a length $\geq 3$) constitutes its base, while two orthogonal rows that are parallel to each other (with equal lengths $\geq 2$) constitute the sides.
+U-shape orientations are similar to those of T- and L-shapes, including horizontal and vertical (up and down).
+Planes
+are 2D shapes ($\geq 3\times 2$ blocks) that are either horizontal along the $xz$ plane or vertical along the $xy$ or $yz$ planes.
+12 Appendix: Experimental Details
+12.1 Curriculum Learning (CL) Algorithm
+We use the Curriculum Arrangement algorithm from the CL framework proposed by Xu et al.
+(2020).
+Specifically, we sort the training dataset $D$ by the difficulty metric(s) and divided into $N$ equal-sized buckets $\{C_{i}:i=1,2,\dotsc,N\}$, corresponding to increasing difficulty levels from $C_{1}$ (easiest) to $C_{N}$ (hardest).
+The curriculum consists of $N$ stages $\{S_{i}:i=1,2,\dotsc,N\}$.
+In each stage $S_{i}$, a $1/N$ fraction of examples is sampled from each preceding bucket $\{C_{j}:j=1,2,\dotsc,i\}$, with the proportion of difficult examples gradually increasing across stages from 0 to eventually match the original dataset distribution.
+Within each stage, examples are still shuffled to maintain local stochastics, and there is no overlap between stages to prevent overfitting.
+At each stage, the model is trained for one epoch.
+After reaching $S_{N}$, the model should be ready for the original distribution in $D$.
+Hence, an additional stage $S_{N+1}$ is introduced, covering the entire dataset $D$, and training continues until convergence.
+12.2 How much of each synthetic dataset to use during train time?
+We provide further details of the procedure in Section 6.1.
+We first train models on the three synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) separately to study how increasing the training data affects performance.
+Focusing on the saturation regions of the learning curves—where performance plateaus and additional data has diminishing returns—we identify various "candidate" data amounts (#items) that yield near-peak performance for each dataset.
+Using these data amounts, we construct a grid of combinations across the three datasets, where each grid point corresponds to a different data mix and proportion of each dataset.
+For each grid point, we train a model on the combined $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ and $D_{\mathrm{mc}}$ datasets.
+The combination that achieves the best performance on $D_{\mathrm{mc}}$ is selected.
+Thus, the approach yields a data mix that optimizes $D_{\mathrm{mc}}$ performance while also being highly performant on the synthetic data as much as possible.
+Also, in the best data mixes, the #items per dataset generally followed the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$, which matches their difficulty ordering (as explored for CL in Section 6.2.1).
+12.3 Experimental Setup for GRU-based Models
+We provide further details of the experimental setup for models $M_{\mathrm{agg}}$ and $M_{\mathrm{agg+cl}}$ (apart from those in Section 6.1.1), again closely following the setup used for the baseline model $M_{\mathrm{mc}}$ when it was trained on $D_{\mathrm{mc}}$ alone (Section 2.6).
+Training
+We train models using AdamW Loshchilov and Hutter (2019) and weight decay regularization with a weight decay factor of 0.1.
+The learning rate is set to 0.0001, and the batch size is 1.
+Early stopping is applied when the loss on the held-out development set increases monotonically for ten epochs.
+The development set for $M_{\mathrm{agg}}$ and $M_{\mathrm{agg+cl}}$, used for early stopping, is constructed from the validation splits of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ (Table 3) to match the dataset distribution in the training set.
+Model Hyperparameters
+We continue to use the same hyperparameter values as those of the best model in Jayannavar, Narayan-Chen, and Hockenmaier (2020).
+12.4 Experimental Setup for BERT-based Models
+We provide further details for the experimental setup of $M_{\mathrm{mc}}^{\prime}$, $M_{\mathrm{agg}}^{\prime}$, and $M_{\mathrm{agg+cl}}^{\prime}$ (apart from those mentioned in Section 6.3.1).
+Training
+The training setup is similar to Section 12.3, with the following differences: a learning rate of 0.0001 for $M_{\mathrm{mc}}^{\prime}$ and 0.0005 for $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg+cl}}^{\prime}$, a batch size of 16, and a learning rate schedule consisting of linear warmup followed by inverse square root decay with 4000 warmup steps.
+Model Hyperparameters
+The transformer decoder consists of 5 layers with a representation dimensionality of 256, 2 attention heads, and an inner feedforward layer dimensionality of 512.
+The model uses uncased BERT with a maximum input length of 512 tokens, the highest allowed.
+Other hyperparameters are similar to Section 12.3.
+13 Appendix: Additional Results
+This section contains the BAP v2 metrics for the examples in Sections 7.2 and 8.1.
 Example ID Model Type Color Location Shape Overall
@@
 Table 11: BAP v2 metrics by Example ID for Section 8.1 (range $[0,1]$)
-12 Appendix: Concurrent work
+14 Appendix: Annotation Setup
+Our annotation team comprises one expert annotator (an author) and 10 non-expert annotators, who are undergraduate and graduate computer science students.
+The non-experts received elaborate training from the expert annotator before the final data collection.
+The NEB (non-empty board) subset is annotated by the 10 non-expert annotators, with each BAP item annotated by two annotators.
+Disagreements were resolved by the expert annotator, and the workload was evenly distributed among the annotators (in terms of the number of dialogs per annotator).
+The smaller and easier-to-annotate EB (empty board) subset was exclusively annotated by the expert annotator.
+Annotators were instructed to annotate one dialog at a time, processing items within a dialog in chronological order.
+This ensures they accurately understand the evolving game context and remain unbiased by future "unseen" events in a dialog.
+15 Appendix: Extended Discussion on Concurrent Work
+Apart from the discussion on concurrent work and Nebula’s evaluation in Section 8.3, a slightly more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in this section given some other relevant aspects of their work.
+(The detailed tables for Nebula’s performance on the full v2 benchmark mentioned in Section 8.3 are also provided here (Tables 13 and 14).)
+15.1 Evaluation
+Chaturvedi, Thompson, and Asher (2024) also identify some of the issues with BAP evaluation on $D_{\mathrm{mc}}$ that we do, and hence propose evaluation on synthetic data.
+They design simpler, non-dialog-based data with corresponding evaluation metrics.
+The data is single-turn and evaluates a model’s ability to follow an instruction in two scenarios: when the board is empty and the instruction is to build a single shape, or when a single shape is already present and the instruction involves placing or removing one block.
+While this data is intentionally simple to test elementary Builder capabilities, it overly simplifies the task and therefore deviates significantly from the target MCBT/BAP.
+It lacks a dialog component, game history, and relative spatial references such as "left" and "right" as the Builder’s position and orientation are not considered.
+The world state/built structure is also fairly minimal, either empty or containing a single shape.
+While useful for elementary testing, it is a bit unclear how well insights from this data translate to the full BAP task.
+In contrast, our synthetic data remains simpler than $D_{\mathrm{mc}}$ while also closely emulating the MCBT, addressing these limitations and preserving many of the key BAP features and challenges, such as diverse instructions, rich spatial relations and floating blocks for example (more in Sections 2.1.1 and 5).
+Although designed/motivated for training more than evaluation, its simplicity allows assessment of basic model competencies, while its emulation of the MCBT enhances its applicability to the overall BAP task, thus striking a better balance from an evaluation standpoint.
+Section 8.2 details additional benefits of evaluating on our synthetic data.
+Additionally, our synthetic data evaluation, combined with fairer F1, enables meaningful performance comparisons between synthetic data and $D_{\mathrm{mc}}$ (an example utility of this is the insights it provides to motivate Curriculum Learning).
+Chaturvedi, Thompson, and Asher (2024) are unable to conduct such comparisons (albeit understandably) due to the limitations of the legacy/original F1 metric.
+At a high level, Chaturvedi, Thompson, and Asher (2024) propose evaluation on their synthetic data as a solution to BAP evaluation issues, whereas we take the complementary approach by addressing them directly to propose the BAP v2 evaluation benchmark.
+Additionally, as described above, we also highlight the importance of evaluation on our synthetic data and contrast it with theirs.
+15.2 Nebula+
+While Chaturvedi, Thompson, and Asher (2024) primarily design their synthetic data for evaluation, they also report results from finetuning Nebula (Llama-3-8B) jointly on it and $D_{\mathrm{mc}}$.
+We refer to this model as Nebula+.
+(Recall that the original Nebula model was trained on $D_{\mathrm{mc}}$ alone.) Nebula+ improves performance on their synthetic data but remains similar to Nebula on $D_{\mathrm{mc}}$ (achieving an F1 score of 39.1% under the legacy F1, compared to Nebula’s 39.2%).
+This is fairly understandable given that their synthetic data, designed mainly for evaluation rather than training, does not capture sufficient task complexity, as discussed above.
+In contrast, as shown in Section 6, our approach of training jointly on our richer synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and $D_{\mathrm{mc}}$ improves performance on both (and then further boosted by Curriculum Learning).
+Similar to Nebula’s evaluation in Section 8.3, we evaluate Nebula+ here.
+Detailed results for the full v2 benchmark are provided in Tables 15 and 16.
+Table 12 presents Nebula and Nebula+ performance (micro F1) on both our synthetic data and the original $D_{\mathrm{mc}}$ test sets (extending the previous Table 9 with an additional row for Nebula+).
+The key trends and conclusions remain consistent with those previously discussed for Nebula—demonstrating the impact and robustness of our v2 evaluation, providing a fairer model assessment, and reinforcing the importance of evaluating on our synthetic data.
+Additionally, Table 12 shows that Nebula+ slightly underperforms compared to Nebula on our synthetic data while maintaining similar performance on $D_{\mathrm{mc}}$.
+This suggests that improvements on their synthetic data do not necessarily translate to better performance on ours (fairly understandable as their synthetic data does not capture sufficient task complexity).
+Model Test Data
+$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
+Nebula 51.5 33.1 30.9 50.1
+Nebula+ 48.8 32.4 30.4 50.1
+Table 12: Nebula and Nebula+’s performance across datasets (v2 evaluation on $D_{\mathrm{mc}}$)
 Dataset Type Color Location Shape Overall
@@
 Table 14: Macro F1 scores for the Nebula model under the v2 benchmark
-[…]
+Dataset Type Color Location Shape Overall
+EB 94.8 94.8 80.6 83.7 80.6
+NEB 78.9 77.7 46.9 58.7 46.3
+Overall 80.7 79.6 50.7 61.6 50.1
+Table 15: Micro F1 scores for the Nebula+ model under the v2 benchmark
+Dataset Type Color Location Shape Overall
+EB 97.5 97.5 83.6 88.9 83.6
+NEB 82.8 82.0 48.6 64.1 48.1
+Overall 84.0 83.3 51.5 66.1 51.0
+Table 16: Macro F1 scores for the Nebula+ model under the v2 benchmark
+Acknowledgements.
+We would like to thank Hetvi Patel, Sana Madhavan, Marc Canby, Risham Sidhu, Siddarth Madala, Rajarshi Haldar, Katya Yegorova, Adam Davies, Ansel Blume, and Sara Aghajanzadeh for their valuable help with the data annotation.
+This work was supported by Contract W911NF-15-1-0461 with the US Defense Advanced Research Projects Agency (DARPA) Communicating with Computers Program and the Army Research Office (ARO).
+Approved for Public Release, Distribution Unlimited.
+The views expressed are those of the authors and do not reflect the official policy or position of the Department of Defense or the U.S.
+Government.
+The work was also supported by the Distributed Research Experiences for Undergraduates (DREU) program, a joint project of the CRA Committee on the Status of Women in Computing Research (CRA-W) and the Coalition to Diversify Computing (CDC), which is funded in part by the NSF Broadening Participation in Computing program (NSF BPC-A #1246649).
+The work utilizes resources Kindratenko et al.
+(2020) supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign.

### v3

@@
 Abstract
-Interactive agents capable of understanding and executing instructions in the physical world have long been a central goal in AI research.
-The Minecraft Collaborative Building Task (MCBT) provides one such setting to work towards this goal (Narayan-Chen, Jayannavar, and Hockenmaier, 2019).
-It is a two-player game in which an Architect (A) instructs a Builder (B) to construct a target structure in a simulated Blocks World Environment.
-We focus on the challenging Builder Action Prediction (BAP) subtask of predicting correct action sequences in a given multimodal game context with limited training data (Jayannavar, Narayan-Chen, and Hockenmaier, 2020).
-We take a closer look at evaluation and data for the BAP task, discovering key challenges and making significant improvements on both fronts to propose BAP v2, an upgraded version of the task.
-This will allow future work to make more efficient and meaningful progress on it.
-It comprises of: (1) an enhanced evaluation benchmark that includes a cleaner test set and fairer, more insightful metrics, and (2) additional synthetic training data generated from novel Minecraft dialogue and target structure simulators emulating the MCBT.
-We show that the synthetic data can be used to train more performant and robust neural models even with relatively simple training methods.
-Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs.
-Although modeling is not the primary focus of this work, we also illustrate the impact of our data and training methodologies on a simple LLM- and transformer-based model, thus validating the robustness of our approach, and setting the stage for more advanced architectures and LLMs going forward.
+Developing interactive agents that can understand language, perceive their surroundings, and act within the physical world is a long-standing goal of AI research.
+The Minecraft Collaborative Building Task (MCBT) (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), a two-player game in which an Architect (A) instructs a Builder (B) to construct a target structure in a simulated 3D Blocks World environment, offers a rich platform to work towards this goal.
+In this work, we focus on the Builder Action Prediction (BAP) subtask: predicting B’s actions in a multimodal game context (Jayannavar, Narayan-Chen, and Hockenmaier, 2020) – a challenging testbed for grounded instruction following, with limited training data.
+We holistically re-examine this task and introduce BAP v2 to address key challenges in evaluation, training data, and modeling.
+Specifically, we define an enhanced evaluation benchmark, featuring a cleaner test set and fairer, more insightful metrics that also reveal spatial reasoning as the primary performance bottleneck.
+To address data scarcity and to teach models basic spatial skills, we generate different types of synthetic MCBT data.
+We observe that current, LLM-based SOTA models trained on the human BAP dialogues fail on these simpler, synthetic BAP ones, but show that training models on this synthetic data improves their performance across the board.
+We also introduce a new SOTA model, Llama-CRAFTS, which leverages richer input representations, and achieves an F1 score of 53.0 on the BAP v2 task and strong performance on the synthetic data.
+While this result marks a notable 6 points improvement over previous work, it also underscores the task’s remaining difficulty, establishing BAP v2 as a fertile ground for future research, and providing a useful measure of the spatial capabilities of current text-only LLMs in such embodied tasks.
 1 Introduction
-There has long been interest in developing interactive agents that can communicate with humans about and operate within the physical world (e.g., Winograd (1971)).
-The goal for these agents is not only to engage in rich natural language dialogue, but also to ground it to physical objects, and execute instructions in the real world.
-We work toward this goal using Minecraft (https://minecraft.net/), a popular multiplayer game where players control avatars to navigate a 3D world and build structures from block-like materials.
-Minecraft, in recent years, has garnered a lot of interest as an AI experimentation platform.
-Our previous work—the Minecraft Collaborative Building Task (MCBT) and the corresponding Minecraft Dialogue Corpus (MDC) Narayan-Chen, Jayannavar, and Hockenmaier (2019)—was among the first to use it to study grounded task-oriented dialogue.
-Jernite et al.
-(2019) is another such example.
-More recent works, such as Mohanty et al.
-(2024), have been directly inspired by the MCBT.
-Additionally, there now exists a well-developed set of platform and tools that support AI experimentation within Minecraft, including Malmo Johnson et al.
-(2016), Craftassist Gray et al.
-(2019), TaskWorldMod Ogawa et al.
-(2020), MC-Saar-Instruct Köhn et al.
-(2020), IGLU GridWorld Zholus et al.
-(2022), and Narayan-Chen, Jayannavar, and Hockenmaier (2019).
-The latter builds on the Malmo platform.
+Developing interactive agents that can understand language, perceive their surroundings, and act within the physical world is a long-standing goal of AI research (e.g., Winograd (1971)).
+We work toward this goal in Minecraft (https://minecraft.net/), a popular 3D game that has become a prominent platform for AI experimentation.
+Our previous work introduced the Minecraft Collaborative Building Task (MCBT) and the corresponding Minecraft Dialogue Corpus (MDC) Narayan-Chen, Jayannavar, and Hockenmaier (2019), one of the first efforts to use this environment to study grounded, task-oriented dialogue.
 Figure 1: In the Minecraft Collaborative Building Task, the Architect (A) has to instruct a Builder (B) to build a target structure.
@@
 (NB: We show B’s actions in the dialogue as a visual aid to the reader.)
-The MCBT situates a dialogue task—designed to be asynchronous, asymmetric, and with minimal language constraints—in a simulated 3D Minecraft environment where participants, with constantly shifting perspectives, must refer to and modify a dynamic world.
-In this task, two players collaborate: an Architect (A) instructs a Builder (B) to construct a target structure using multi-colored blocks (Figure 1).
-The MDC comprises 509 human-human game logs for this task.
-Narayan-Chen, Jayannavar, and Hockenmaier (2019) focus on generating Architect utterances, and Jayannavar, Narayan-Chen, and Hockenmaier (2020) focus on end-to-end neural models for building an automated Builder agent.
-The latter defined the challenging Builder Action Prediction (BAP) task, aimed at predicting the sequence of actions (block placements and/or removals) a human Builder performed at a specific point in a human-human game.
-Despite the task’s inherent challenges and limited training data, we demonstrate that models trained with a suitable amount of game history, enriched world representations, and a sufficient amount of synthetic, diversified data (via data augmentation techniques) yield promising initial results.
-We recap relevant prior work on the Minecraft Builder in Section 2.
-Building on this foundation, we next take a closer look at evaluation and training data for the BAP task, uncovering key challenges and making significant improvements on both fronts to propose BAP v2, an upgraded version of the task.
-This will allow future work to make more efficient and meaningful progress on it.
-Given the inherent complexity of the task, we find that evaluation is nuanced and requires closer examination.
-We identify key challenges that hinder fair and insightful evaluation, then systematically address them.
-To this end, we introduce a new, cleaner v2 test set to replace the original legacy test set for evaluating BAP models (Section 3).
-We also refine the F1 metric, proposing a fairer variant that accounts for nuances in the data.
-The current evaluation, which relies solely on an aggregated and opaque F1 metric, lacks detailed insights into model behavior.
-To enhance insightfulness, we introduce additional metrics that measure specific model capabilities.
-Together, these contributions form the updated BAP v2 evaluation benchmark (Section 4).
-Our work on evaluation highlights spatial reasoning as a key bottleneck in model performance, alongside the challenge of limited training data (even with the aforementioned data augmentation).
-Instead of merely collecting more data (which can be impractical/expensive in realistic scenarios for such complex situated dialogue tasks), we create additional diverse synthetic training data generated from novel Minecraft dialogue and target structure simulators that emulate the MCBT (Section 5).
-This data, though naturally simpler than it, is rich in spatial relations and referential expressions and we show that it can be successfully used to train better models, even with straightforward training methods (Section 6).
-We demonstrate that jointly training on combined synthetic and original BAP (augmented) data already improves performance.
-Further, a straightforward Curriculum Learning method provides an additional boost.
-Together, the synthetic and augmented BAP data form the BAP v2 training set.
-Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs.
-We provide a detailed analysis of model performance of our best performing model relative to the baseline in Section 7.
-Although modeling is not the primary focus of this work, solely to test the robustness of our synthetic data and training methodologies and show their applicability to relatively contemporary model architectures, we experiment with very simple LLM (BERT)- and transformer-based variants of our neural models.
-The simplicity also allows us to assess whether/to what extent simpler LLMs like BERT and vanilla Transformers alone can already effectively handle the BAP task.
-Our results show that the synthetic data and training methodologies improve these models as well, validating the robustness of our approach (Section 6.3).
-Future work can build on these insights to explore more sophisticated LLMs and architectures while leveraging the BAP v2 framework.
-Section 8 provides a detailed discussion on four topics – an additional key challenge in BAP evaluation, the importance of evaluation on the synthetic data, relevant concurrent work, and broader implications of our work that go beyond BAP.
-The latter further highlights the potential impact of our work and directions for future research.
-Related work is covered in Section 9.
-We conclude and discuss future work in Section 10.
-We plan to release code and data involved in this work.
+In the MCBT, two players collaborate to construct a target structure from multi-colored blocks (Figure 1): an Architect (A) provides instructions to a Builder (B).
+The task is situated in a dynamic 3D world that participants must refer to and modify from constantly shifting perspectives.
+The dialogue is asynchronous, asymmetric, and has minimal linguistic constraints.
+The MDC contains 509 human-human game logs from this task.
+While Narayan-Chen, Jayannavar, and Hockenmaier (2019) focused on generating Architect utterances, Jayannavar, Narayan-Chen, and Hockenmaier (2020) introduced end-to-end models for an automated Builder.
+The latter work defined the challenging Builder Action Prediction (BAP) task: predicting the sequence of actions (block placements and/or removals) that a human Builder performed at a given point in a human-human game.
+We review prior work on the Minecraft Builder and related recent works in Section 2.
+The BAP task remains a relevant and challenging testbed for grounded instruction following, presenting a unique combination of difficulties.
+Unlike navigation or object-centric benchmarks such as VLN (Anderson et al., 2018), CVDN (Thomason et al., 2019), ALFRED (Shridhar et al., 2020), DialFRED (Gao et al., 2022), and TEACh (Padmakumar et al., 2022), where agents primarily interact with existing objects, the MCBT requires grounding instructions that refer to objects and structures which do not yet exist, demanding greater abstraction and planning.
+The dialogue is also more complex than in un-embodied benchmarks like MultiWoz (Budzianowski et al., 2018) or other Minecraft datasets like IGLU (Mohanty et al., 2024; Kiseleva et al., 2022b, a); it features free-form, asynchronous communication within a world that is dynamically modified.
+Finally, the task poses a significant planning problem due to its large action space—over 7,600 possible actions compared to just six in typical VLN work—and its reliance on complex, perspective-dependent spatial relations (e.g., “to your left"), which are more difficult to resolve than the absolute cardinal directions used in IGLU.
+This combination of challenges makes BAP a compelling setting for research.
+However, to enable future research to make more efficient and meaningful progress, this paper holistically re-examines the BAP task framework.
+In particular, we identify and address key challenges in evaluation, training data, and modeling, and propose BAP v2, an upgraded version of the task: [^1] All code and data developed for this research will be released at https://github.com/prashant-jayan21/bap-v2
+Evaluation (Sections 3 and 4)
+The complexity of the BAP task makes evaluation nuanced.
+We systematically address challenges that hinder fair and insightful assessment by introducing a cleaner v2 test set (Section 3), propose a fairer F1 metric and introduce new auxiliary metrics to measure specific model capabilities (Section 4).
+Together, these contributions establish the BAP v2 evaluation benchmark and help us move beyond the original, opaque F1 score, providing deeper insights.
+Adopting this new benchmark also reveals that spatial reasoning is the key performance bottleneck in the BAP task.
+Training Data (Section 5)
+To overcome the dual challenges of limited training data for the BAP task—stemming from the small size of the MDC (509 game logs)—and the key performance bottleneck of spatial reasoning, we create a diverse set of synthetic training data.
+This data is generated using novel Minecraft dialogue and target structure simulators that were designed emulate the MCBT.
+While simpler than the human MDC data, the synthetic dialogues are rich in the spatial relations and referring expressions crucial for this task.
+The combination of this new data with the original corpus forms the BAP v2 training set.
+Our approach builds on established work showing the effectiveness of synthetic data for various tasks, from vision-and-language navigation to dialogue systems (Kamath et al., 2023; Ku et al., 2020; Wang et al., 2023; Kang et al., 2023; Das et al., 2017; Kim et al., 2022; Bao et al., 2023; Zhan et al., 2023, 2024; Tang et al., 2024), although synthesizing complete, task-oriented embodied dialogues remains a relatively underexplored area (Padmakumar et al., 2023), especially since the MCBT is significantly more complex than other Blocks World settings (Bisk, Yuret, and Marcu, 2016; Dan, Han, and Roth, 2021).
+Modeling and Analysis (Sections 6 and 7)
+To demonstrate the utility of our synthetic data in addressing the challenge of limited original BAP data, we show that it produces more performant models, for both LLM and non-LLM architectures, even with straightforward training methods (Section 6).
+The data also reveals that both LLM and non-LLM models trained solely on the complex BAP human dialogues lack robustness, counterintuitively underperforming on our simpler synthetic data.
+As the final component of the BAP v2 framework, we extend prior work on LLM-based models for the BAP task (Chaturvedi, Thompson, and Asher, 2024) by introducing richer input text representations that capture essential game context, and we find that these enhanced representations and the synthetic data are both necessary and mutually reinforcing.
+Our best model, Llama-CRAFTS (Context Rich And Fine-Tuned On Synthetic Data), establishes a new state of the art on the BAP task as well as achieves strong performance on the synthetic data.
+But while its F1 score of 53.0 on BAP represents a 6 point improvement over the previous SOTA (Chaturvedi, Thompson, and Asher, 2024) (under the new v2 evaluation), it is nevertheless clear that there is much room for further improvement.
+A detailed analysis of model output (Section 7) illustrates how Llama-CRAFTS improves over the previous SOTA (especially in spatial reasoning), where it falls short, what this implies for current text-only LLMs’ spatial capabilities in such embodied tasks, and indicates both strengths and limitations of our new evaluation metrics.
+While spatial reasoning in language models has been explored in text and across other modalities (Mirzaee and Kordjamshidi, 2022; Shi, Zhang, and Lipani, 2022; Rizvi, Zhu, and Gurevych, 2024; Gopinathan et al., 2024; Li et al., 2024a; Bisk et al., 2018; Du et al., 2024; Kamath, Hessel, and Chang, 2023; Wang et al., 2024), our work addresses a specific gap in the literature: exploring the capabilities of text-only LLMs in a 3D, embodied, collaborative dialogue setting.
+Prior work has shown LLMs struggle with even simple 2D grid navigation and path-planning in text (Yamada et al., 2024; Aghzal, Plaku, and Yao, 2024), and has explored spatial reasoning in games like chess (Feng et al., 2023a; Zhang et al., 2025b) or puzzles (Chollet et al., 2025), but no existing research combines this with the challenges of 3D grids in text, and embodied, collaborative dialogue, thus motivating our focus.
+Part 4: Looking Ahead (Sections 8 and 9)
+Finally, we discuss future research directions (Section 8) We outline the key next challenges for BAP that can be tackled by building on the BAP v2 framework, and explore the broader implications of our work beyond BAP, including way in which ideas from our synthetic data and simulators hold the potential for furthering other related works and tasks (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Mohanty et al., 2024; Bonn et al., 2020; Thompson, Hunter, and Asher, 2024; Bonial et al., 2021; Madge et al., 2025).
+We conclude in Section 9.
 2 Prior Work on the Minecraft Builder
-This section provides a high-level overview of the dataset, task, and baseline model, summarizing our prior work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Jayannavar, Narayan-Chen, and Hockenmaier, 2020).
-We also introduce some necessary formalisms and figures for better exposition.
-((a))
-((b))
-((c))
-((d))
-((e))
-((f))
-((g))
-((h))
-Figure 2: A sample sequence of human-human game states.
+Sections 2.1,2.2,2.3,2.5 and 2.6 provide a high-level overview of the Minecraft Collaborative Building Task and Dialogue corpus, the BAP task and dataset, evaluation metrics, and baseline model, summarizing our prior work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Jayannavar, Narayan-Chen, and Hockenmaier, 2020).
+Section 2.4 formalizes the BAP task.
+Section 2.7 discusses recent related work around more contemporary LLM-based models for BAP (Chaturvedi, Thompson, and Asher, 2024; Kranti, Hakimov, and Schlangen, 2024), and Section 2.8 discusses related work around the task and dataset itself, explaining why the BAP task continues to be a relevant and challenging testbed for grounded instruction following.
+A snippet from the MDC corpus
+Refer to caption Refer to caption Refer to caption Refer to caption
+(a) (b) (c) (d)
+Refer to caption Refer to caption Refer to caption Refer to caption
+(e) (f) (g) (h)
+Figure 2: A sample sequence of human-human game states from the MDC corpus.
 The game starts with an empty board and an initial A instruction (a), which B executes in the first action sequence (b) by placing a single block.
 In (c), B begins to execute the next A instruction given in (b).
-However, A interrupts B in (c), leading to two distinct B action sequences: (b)–(c) (single block placement), and (c)–(h) (multiple placements and removals).
+But A interrupts B in (c), leading to a new B action sequence in (c)–(h) (multiple placements and removals).
 2.1 The Minecraft Collaborative Building Task and Dialogue Corpus
@@
 The task is complete when the Built structure inside the Build Region matches the Target, allowing for translations and rotations in the horizontal plane.
-2.1.1 Features and Challenges
-Since target structures can be fairly complex, Architects typically give step-by-step instructions ("now a yellow 6") for different parts of the target.
-Builders should execute these instructions, but may also ask questions ("Where does the 6 start") which the Architects has to answer ("behind the 7 from your perspective").
-Architects may also need to identify and correct mistakes the Builder may have made ("too much overlap unfortunately").
+2.1.1 Features and Challenges of the MDC
+Since target structures can be fairly complex, Architects typically give step-by-step instructions (“now a yellow 6") for different parts of the target.
+Builders should execute these instructions, but may also ask questions (“Where does the 6 start") which the Architects has to answer (“behind the 7 from your perspective").
+Architects may also need to identify and correct mistakes the Builder may have made (“too much overlap unfortunately").
 A can move around freely, but remains invisible to B and views the structure from behind B when giving instructions.
@@
 Due to the asynchronous nature of the dialogue, the Architect often talks while the Builder is placing blocks (“so it will look like a v"), leading to an apparent interruption of the Builder’s action sequences.
-Figure 2 shows an example from the MDC that highlights some of these challenges.
-Floating blocks
-Blocks do not need to be placed on the ground if their cell is adjacent to a supporting block.
+Finally, Minecraft, blocks do not need to be placed on the ground if their cell is adjacent to a supporting block.
 If that supporting block is later removed, the remaining block (and any structure supported by it) will “float” in place.
 Thus, placing floating blocks needs the placement and subsequent removal of such placeholder supporting blocks.
-Instructions for floating structures vary greatly, ranging from step-by-step instructions involving temporary supporting blocks to single-shot descriptions such as, simply, “place a floating yellow block”.
+Instructions for floating structures vary greatly, ranging from step-by-step instructions involving temporary supporting blocks to single-shot descriptions such as, simply, “place a floating yellow block".
+Figure 2 shows an example from the MDC that highlights some of these features and challenges.
 2.1.2 MCBT subtasks: Architect Utterance Generation and Builder Action Prediction
-We define the Architect Utterance Generation (AUG) Task as the task of generating a suitable Architect utterance at any point in a human-human game at which the human Architect produced an utterance (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), and the Builder Action Prediction (BAP) Task as the corresponding task of generating a suitable Builder action sequence (consisting of block placements/removals) at any point in a human-human game at which the human Builder placed or removed blocks (Jayannavar, Narayan-Chen, and Hockenmaier, 2020).
+Although the MDC should ultimately contribute to the creation of fully interactive agents that can complete an entire game, we have found it helpful, and even necessary, to start with simpler tasks.
+Specifically, we have defined the Architect Utterance Generation (AUG) Task as the task of generating a suitable Architect utterance at any point in a human-human game at which the human Architect produced an utterance (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), and the Builder Action Prediction (BAP) Task as the corresponding task of generating a suitable Builder action sequence (consisting of block placements/removals) at any point in a human-human game at which the human Builder placed or removed blocks (Jayannavar, Narayan-Chen, and Hockenmaier, 2020).
 Although both of these tasks consider only a single turn, and assume the context of a game and dialogue between two human players, they are important first steps towards the creation of agents that can complete an entire game in the MCBT with a human counterpart.
-We note that compared to BAP, Shi, Feng, and Lipani (2022) propose a slightly different building task involving predicting a single action instead of a sequence of actions.
-Thus, their setup is not directly comparable to ours.
+Three BAP task items corresponding to the snippet in Figure 2
+Item 1 corresponding to (a)–(b) in Fig. 2
+Item 2 corresponding to (b)–(c) in Fig. 2
+Item 3 corresponding to (c)–(h) in Fig. 2
+Figure 3: The snippet in Figure 2 corresponds to three BAP task items shown here.
 2.2 The Builder Action Prediction (BAP) Task
@@
 Games start with an empty board, and consist of an alternating sequence of dialogues $D^{t}$ (which may each consist of a single utterance, or a back-and-forth exchange between the two players) and Builder actions $A^{t}$ that result in an updated built structure $S^{t}$.
-Each action sequence $A^{t}$ yields a BAP item $(\mathcal{H}^{t-1},S^{t-1},D^{t},A^{t},S^{t})$ that consists of the game history $\mathcal{H}^{t-1}$ that culminated in the previous built structure $S^{t-1}$, the new dialogue $D^{t}$, and the new action sequence $A^{t}$ that resulted in the new structure $S^{t}$.
+Each action sequence $A^{t}$ yields a BAP item $(\mathcal{H}^{t-1},S^{t-1},B^{t},D^{t},A^{t},S^{t})$ that consists of the game history $\mathcal{H}^{t-1}$ that culminated in the previous built structure $S^{t-1}$, the Builder’s position and orientation, $B^{t}$, the new dialogue $D^{t}$, and the new action sequence $A^{t}$ that resulted in the new structure $S^{t}$.
 Given $\mathcal{H}^{t-1}$, $S^{t-1},$ and $D^{t}$, BAP models $M$ should predict an action sequence $A^{t}_{M}$ that yields $S^{t}$.
-Dataset statistics
-The training, test and development splits of the MDC contain 3709, 1616, and 1331 items respectively, and the average sequence length of an action sequence (in the development set) is 4.3 (with a std.
-deviation of 4.5).
+Figure 3 illustrates how the game snippet shown in Figure 2 corresponds to three different items for the BAP task.
+The game starts with an empty board $S^{0}$, and an initial dialogue between the two players $D^{1}$, and the first BAP item consists of predicting that first block placement $A^{1}$ (and consequent structure $S^{1}$) (Figure 3, Item 1).
+Then, another instruction follows, which the Builder executes by placing another block, resulting in the next structure to be predicted (Figure 3, Item 2), and the third instruction yields a complex action sequence that results in the desired v-shape structure (Figure 3, Item 3).
+How to interpret BAP item figures in this paper
+The leftmost columns in each BAP item figure (e.g.
+Figure 3) display the (truncated) game history ($S^{t-2}$, $D^{t-1}$, $S^{t-1}$, and $D^{t}$), while the rightmost column shows the resulting built structure $S^{t}$.
+Columns are read top to bottom and left to right.
+We use $S_{H}$ instead of $S$ to indicate that the structure was built by a human ($H$), not a model).
+To highlight the effect of the final action sequence, previously placed blocks are shown in lighter colors.
+Structures like $S^{t-1}$ are rotated to align with the B’s POV $B^{t}$ (corresponding to the next turn and dialogue segment $D^{t}$ below it).
+For B’s POV, we only use the yaw angle, and B’s exact position and pitch angle are omitted for clarity.
+Any temporary supporting blocks are not shown as well.
+2.3 The original BAP dataset
+The training, test and development splits of the MDC contain 3709, 1616, and 1331 BAP items respectively, and the average sequence length of an action sequence (in the development set) is 4.3 (with a standard deviation of 4.5).
 Target structures in the test data do not appear in the training or development data.
-Example
-To illustrate this task, Figure 3 shows how the game snippet shown in Figure 2 corresponds to three different items for the BAP task.
-The leftmost columns in each figure display the (truncated) game history ($S^{t-2}$, $D^{t-1}$, $S^{t-1}$, and $D^{t}$), while the rightmost column shows the resulting built structure $S^{t}$.
-Columns are read top to bottom and left to right.
-Note that in this figure (and similar ones in this work), we use $S_{H}$ instead of $S$ for clarity ($H$ indicates that the structure was built by a human, not a model).
-To highlight the effect of the final action sequence, previously placed blocks are shown in lighter colors, and each structure is rotated based on B’s gaze (only the yaw angle) at the end of the dialog snippet below it.
-B’s exact position and pitch angle are omitted for clarity.
-Any temporary supporting blocks are not shown as well.
-The game starts with an empty board $S^{0}$, and an initial dialogue between the two players $D^{1}$, and the first BAP item consists of predicting that first block placement $A^{1}$ (and consequent structure $S^{1}$) (Figure 3(a)).
-Then, another instruction follows, which the Builder executes by placing another block, resulting in the next structure to be predicted (Figure 3(b)), and the third instruction yields a complex action sequence that results in the desired v-shape structure (Figure 3(c)).
-((a)) Item 1
-((b)) Item 2
-((c)) Item 3
-Figure 3: The snippet in Figure 2 corresponds to three BAP task items as shown.
+Since the small size of the training set is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three simple data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B).
+Empirically, we found that increasing the training data to 14,836 (4x) items gave the best performance for the GRU-based baseline model (Section 2.6).
+Henceforth, we will refer to this combination of the augmented training data and the original test and development data as the original BAP data.
 2.4 Formalizing the BAP task
@@
 The Builder can move freely around the environment.
-Since spatial relations in the instructions (“the block in front of you", “to your left") often depend on the Builder’s current location and viewpoint, we record the position and pose/orientation of the Builder agent $\textsc{Pos}_{\textbf{B}}=(\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle,\phi_{\textbf{B}},\gamma_{\textbf{B}})$, where $\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle$ is a (real-valued) location (that may or may not be inside the build region), $\phi_{\textbf{B}}\in[-90,...,+90]$ indicates a pitch (vertical rotation) and $\gamma_{\textbf{B}}\in[-180,...,+180]$ a yaw (horizontal orientation).
-Note that $\textsc{Pos}_{\textbf{B}}$ is a critical component of game context as it dictates how spatial relations are interpreted.
+Since spatial relations in the instructions (“the block in front of you", “to your left") often depend on the Builder’s current location and viewpoint, we record the position and pose/orientation of the Builder agent $\textsc{Pos}_{\textbf{B}}=(\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle,\phi_{\textbf{B}},\gamma_{\textbf{B}})$, where $\langle x_{\textbf{B}},y_{\textbf{B}},z_{\textbf{B}}\rangle$ is a (real-valued) location (that may or may not be inside the build region), $\phi_{\textbf{B}}\in[-90,...,+90]$ indicates a pitch (vertical rotation) and $\gamma_{\textbf{B}}\in[-180,...,+180]$ a yaw (horizontal orientation) angle.
+For brevity, we use $B^{t}$ to refer to $\textsc{Pos}_{\textbf{B}}$ at time $t$ in the formal definition of BAP items.
 The Builder’s actions
-[…]
+We specify Builder actions $a=(t,c,\mathbf{c})$ as triplets consisting of an action type $t\in\mathcal{T}=\{\textsc{Place},\textsc{Remove}\}$, a block color $c\in\textsc{BlockColors}$, and a grid cell $\mathbf{c}=\langle x,y,z\rangle\in\textsc{BuildRegion}$ whose state is changed by aa.
+[^4] Since the Malmo API can be used to move agents into positions in which a particular block is accessible to them, we ignore the Builder’s movement, and only consider its block placement and removal actions.
+If a feasible action $a$ is executed, it changes the built structure $S$ to $S^{\prime}\neq S$: $S~\xrightarrow{a}~S^{\prime}$.
+A placement is feasible if $\mathbf{c}$ is empty and either on the ground or adjacent to a non-empty cell, and results the addition of a $c$-colored block located in $\mathbf{c}$ to $S^{\prime}$.
+A removal action is feasible if $S$ contains a contains a block of color $c$ in $\mathbf{c}$, and results the exclusion of that block from $S^{\prime}$.
+When Builders follow an instruction, they execute an action sequence $A=\langle a^{1},...,a^{k}\rangle$ for $k\geq 1$ and $a^{1},...,a^{k}\in\textsc{ActionSpace}$ that changes $S$ to $S^{\prime}$: $S\xrightarrow{A=\langle a^{1},...a^{k}\rangle}S^{\prime}$.
+ActionSpace is the set of all 7623 possible actions (7 actions available per cell in the $11\times 9\times 11$ build region—placing a block in one of 6 colors or removing a block).
 2.5 Evaluating BAP models
-[…]
+To compare two action sequences, or to measure the change from $S$ to $S^{\prime}$, it is helpful to note that any action $a=(t,c,\mathbf{c})$ is undone by the inverse action $a^{-}=(t^{-},c,\mathbf{c})$ where $t^{-}=\textsc{Remove}$ if $t=\textsc{Place}$, and $t^{-}=\textsc{Place}$ if $t=\textsc{Remove}$.
+Most commonly, this occurs when blocks are placed to serve as a necessary supporting block for a floating structure and removed in the same action sequence.
+These supporting blocks can be ignored since they do not occur in $S^{\prime}$.
+Moreover, builders are free to use any color they wish, and can often choose among a variety of possible locations where to place these blocks.
+Human Builders are also prone to accidentally placing or removing blocks, but typically recognize and correct such mistakes immediately within the same action sequence.
+Finally, many structures do not require their blocks to be placed in a particular order.
+We can therefore transform any (feasible) action sequence $A$ into a set of net actions $A^{\mathrm{net}}$ by first removing from $A$ any actions $a$ and their inverse $a^{-}$ and turning the resulting (infeasible) sequence $A^{\prime}$ into a set $A^{\mathrm{net}}$.
+The distance $\Delta(S,S^{\prime})$ between two structures $S$ and $S^{\prime}$ can then be defined as the number of net actions of any action sequence $A:S\xrightarrow{A}S^{\prime}$ that changes $S$ to $S^{\prime}$: $\Delta(S,S^{\prime})=|A^{\mathrm{net}}|$, and any two action sequences $A_{1}$ and $A_{2}$ are equivalent ($A_{1}\equiv A_{2}$), i.e.
+lead to the same subsequent structure, if they have the same set of net actions ($A_{1}^{\mathrm{net}}=A_{2}^{\mathrm{net}}$).
+We therefore evaluate BAP models against human Builders by comparing their respective net actions, $A_{M}^{\textrm{net}}$ and $A_{H}^{\textrm{net}}$ to compute a (strict) F1 score (we report a micro-averaged F1 score over all action sequences in the test/dev data):
 Definition 2.1.
 Given a BAP item consisting of a (human) reference action sequence $A_{H}=\langle a_{H}^{1},...,a_{H}^{k}\rangle$ that leads from $S$ to $S_{H}$, corresponding to a reference net action set $A_{H}^{\textrm{net}}$, and a (model) predicted action sequence $A_{M}=\langle a_{M}^{1},...,a_{M}^{l}\rangle$ that leads from $S$ to $S_{M}$, corresponding to a predicted net action set $A_{M}^{\textrm{net}}$, strict Precision, Recall and F1 scores assume that a Builder action $a_{M}^{m}=(t,c,\langle x,y,z\rangle)\in A_{M}^{\textrm{net}}$ is correct if and only if there is an equal reference action $a_{H}^{h}=(t,c,\langle x,y,z\rangle)\in A_{H}^{\textrm{net}}$:
-Strict $\textbf{Strict Precision}~{}P_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{M}^{\textrm{net}}|}$
-Strict $\textbf{Strict Recall}~{}R_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{H}^{\textrm{net}}|}$
-[…]
-2.5 Data Augmentation
-Since the small size of the training set (3,709 examples) is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B).
-Empirically, we found that increasing the training data to 14,836 (4x) items gave the best performance for our baseline model (Section 2.6).
-We will refer to this as the original BAP (augmented) data.
+$\textbf{Strict Precision}~P_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{M}^{\textrm{net}}|}$
+$\textbf{Strict Recall}~R_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(A_{M}^{\textrm{net}}\cap A_{H}^{\textrm{net}})|}{|A_{H}^{\textrm{net}}|}$
+$\textbf{Strict F1}~F1_{\mathrm{strict}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{2\cdot P_{\mathrm{strict}}\cdot R_{\mathrm{strict}}}{P_{\mathrm{strict}}+R_{\mathrm{strict}}}$
 2.6 A CNN- and GRU-based baseline model
-Figure 4: Baseline model in this work
 In Jayannavar, Narayan-Chen, and Hockenmaier (2020), we proposed a neural model for the BAP task.
-In this work, we build upon it and refer to it as the baseline model.
-The model (Figure 4) is based on a recurrent encoder-decoder architecture (Sutskever, Vinyals, and Le, 2014; Cho et al., 2014) in which a GRU-based encoder (bottom left box) captures the game context (dialogue and action history) via GloVe embeddings (Pennington, Socher, and Manning, 2014), and a CNN-based encoder (top left box) captures the world state at each time step.
+In this work, we build upon it and refer to it as the GRU-based baseline model.
+The model is based on a recurrent encoder-decoder architecture (Sutskever, Vinyals, and Le, 2014; Cho et al., 2014) in which a GRU-based encoder captures the game context (dialogue and action history) via GloVe embeddings (Pennington, Socher, and Manning, 2014), and a CNN-based encoder captures the world state at each time step.
 The world state encompasses the current built structure, action history and the Builder’s position and orientation.
-The decoder (right box) predicts one action per time step, based on the game history, the world state at that time, and the last action taken.
-It consists of another GRU backbone over action sequences (bottom right), and a multi-class classifier that reads in the output of the GRU backbone as well as the world state encoding produced by the CNN to predict either the next action (block placement or removal) to be taken, or a special STOP token that terminates the action sequence.
-Actions are represented as feature vectors and embedded via feedforward layers before being fed to the decoder GRU.
-The world state representation gets updated and re-encoded after each predicted action.
-We train this model with teacher forcing and cross entropy loss on the original BAP (augmented) data, and use greedy decoding (max.
+We trained this model with teacher forcing and cross entropy loss on the original BAP data, and used greedy decoding (max.
 sequence length of 10 actions) at test time.
 This yields an F1 score of 21.2% on the BAP test set.
+2.7 Related Work: LLMs for BAP
+Related to our work are the recent works of Chaturvedi, Thompson, and Asher (2024) and Kranti, Hakimov, and Schlangen (2024), both exploring more modern LLMs for the BAP task.
+The former finetunes the Llama-3-8B model (Dubey et al., 2024) using QLoRA (Dettmers et al., 2023) on the original BAP data, resulting in the Nebula model, which predicts the next action sequence in a text-to-text fashion using all preceding conversation and action sequences as context.
+The latter employs in-context learning (ICL) with GPT-4.
+Both achieve state-of-the-art results over the GRU-based model.
+Nebula achieves an F1 score of 39.2% and GPT-4 scores 39.0% (under the now legacy BAP evaluation), demonstrating that both approaches yield very similar top-line performance (Nebula being marginally better).
+A slightly more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided later in Section 7.4 given some other relevant aspects of their work.
+2.8 Related Work: Tasks and Datasets
+In this section, we explain why the BAP task continues to be a relevant and challenging testbed for grounded instruction following.
+The MDC dataset is a collaborative embodied task, while BAP focuses on grounded language instruction following.
+General collaborative task-oriented dialogue tends to focus on un-embodied domains.
+Typical of the domain is the MultiWoz dataset (Budzianowski et al., 2018) which explores conversations around tasks such as trip planning, which require both dialogue and information retrieval.
+Even tasks that include a visual modality are not explicitly embodied: take Udagawa and Aizawa (2019)’s OneCommon task, which requires two agents to work together to identify the overlap in their fields of view, yet does not include turning to adjust a perspective, etc.
+The MDC, on the other hand, is closer to some vision-and-language works.
+Vision-and-Language Navigation (VLN) (Anderson et al., 2018), and its dialogue counterpart, Cooperative Vision-and-Dialog Navigation (CVDN) (Thomason et al., 2019), which explore instruction-following in photorealistic navigation settings, are good examples.
+Within this domain, others have combined vision and dialogue navigation with task completion, including CEREALBAR (Suhr et al., 2019), a game task in which an instruction-follower and an instruction-giver cooperate to traverse a map and collect items.
+Others have combined vision and language-based task completion in ways that go beyond navigation.
+Roy et al.
+(2021) consider a human and a robot collaborating to search for an object, while ALFRED (Shridhar et al., 2020) is a benchmark for mapping natural language instructions and egocentric vision to action sequences for household tasks in simulated environments.
+Gao et al.
+(2022) extend this by incorporating dialogue instead of just instructions.
+Many of the challenging aspects of the MDC are shared across these tasks, e.g.
+asking for clarification or help is a central component in VLN works Nguyen and Daumé (2019); Chi et al.
+(2019).
+Central to all of these tasks is that agents and humans must communicate with natural language about a world that exists outside of text.
+Although more recent tasks require real vision to represent this world, their underlying world state space (defined by fixed viewpoints and underlying navigation graphs) is just as highly discretized as a Minecraft grid.
+And, while many works do opt for visual inputs, Jia et al.
+(2024) explore the potential of leaving vision out of task-oriented dialogue, and works in less grounded settings, such as Xu et al.
+(2024), often utilize fully textual agents.
+Despite not requiring vision as most of the counterparts discussed above do, our task poses a possibly more challenging planning problem: its action space is much larger (7623 possible actions vs.
+six actions in the vision-language navigation work).
+By excluding vision, MDC can focus solely on instruction following and the decision-making inherent in that.
+Furthermore, Minecraft dialogues are more complex than the other discussed datasets due to the MDC’s asynchronous communication, longer turn sequences, and dynamic environment changes during construction.
+Unlike navigation or simple cooperative games, where referring expressions target existing objects, construction often involves instructions for objects yet to be built.
+The TEACh benchmark (Padmakumar et al., 2022) addresses the former of these limitations with free-form dialogue, similar to the MDC, but not the latter as the task does not require references to unbuilt objects.
+While it does feature object state changes in its 3000+ human-human dialogues, the instruction giver also receives a predefined plan of steps, limiting the diversity since there is only one way to proceed (compared to the many unique ways to build a final MDC structure).
+Additionally, the MDC allows for greater abstraction in instructions, such as referring to substructures, individual blocks, or shapes, whereas TEACh focuses on well-defined objects in the environment.
+This combination of dialogue structure, references to unbuilt objects, plan diversity, and the instruction abstraction make the MDC and BAP tasks both difficult and compelling, as well as providing areas of exploration not addressed by other similar tasks.
+Minecraft Environment
+Minecraft has garnered a lot of interest as an AI experimentation platform in recent years.
+Our previous work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) was among the first to use it to study grounded task-oriented dialogue, and introduced the MCBT and MDC.
+Facebook’s CraftAssist (Gray et al., 2019; Jernite et al., 2019; Szlam et al., 2019) is another such example, enabling two-way human-bot interactions where a human architect instructs an automated builder to build complex structures.
+Their data includes synthetically generated and crowdsourced instructions paired with logical tree structures consisting of action primitives, unlike the MDC, which features human-human dialogues with greater ambiguity, variety, and noisier Builder actions.
+MineDojo (Fan et al., 2022) focuses on creating versatile agents for diverse tasks via an internet-scale knowledge base, contrasting with tasks like IGLU (Mohanty et al., 2024; Kiseleva et al., 2022b, a) and MCBT, which prioritize grounded natural language dialogue and clarification for interactive agents.
+Among Minecraft-related efforts, IGLU is most closely related to the MDC/MCBT and was directly inspired by it.
+Unlike the full dialogue structure of the MDC, the IGLU dataset contains mostly single turns (i.e.
+instruction-action pairs without a broader game context or history).
+A small multi-turn subset is available (127 dialogues), but has strict turn taking in place of MDC’s free-form asynchronous dialogue.
+It further simplifies the task by using cardinal directions (e.g “North", “South") which are absolute and more consistent that relative references (e.g.
+“left", “right").
+Some other works also build on or are inspired by the MCBT/MDC: Bonn et al.
+(2020), Thompson, Hunter, and Asher (2024), Bonial et al.
+(2021), and Madge et al.
+(2025) provide additional linguistic annotations for the MDC, while Shi, Feng, and Lipani (2022) explore when the Builder should ask for clarification or execute an instruction, although their building task consists of predicting single actions instead of action sequences, and is therefore not directly comparable to ours.
 3 Revisiting BAP Evaluation: A Cleaner Test Set
-Since the MDC consists of real human-human game logs collected in a low-stakes, free-flowing setting with minimal constraints on the participants’ language and actions, rather than standalone instructions, this strategy does not always yield BAP items that are suitable for evaluation.
+Suitable BAP test items (with correct structures)
+Clear context with unique interpretation (non-empty board)
+Clear context with unique interpretation (empty board)
+Clear context with multiple interpretations (empty board)
+Figure 4: BAP items with clear contexts and correct structures are suitable for evaluation purposes. If the board is empty, there may be multiple valid interpretations.
+Unsuitable BAP test items: Unclear contexts
+Figure 5: BAP items with unclear contexts are unsuitable for evaluation purposes
+Unsuitable BAP test items: Incorrect structures
+Figure 6: BAP items with incorrect structures are unsuitable for evaluation purposes.
+Fixable BAP items with incorrect structures
+The structure of item $i$ can be replaced by the structure of item $i+1$
+Item $i$
+Item $i+1$
+The structures of item $i$ and $i+1$ can be replaced by the structure of item $i+2$
+Item $i$
+Item $i+1$
+Item $i+2$
+Figure 7: A correct structure following one or more BAP items with incorrect built structures can be used to fix the incorrect BAP items.
+Since the MDC consists of real human-human game logs collected in a low-stakes, free-flowing setting with minimal constraints on the participants’ language and actions, rather than standalone instructions, not all action sequences correspond to BAP items that are suitable for evaluation purposes.
 First, Builders might move blocks even when no new instructions were given, or before it is completely clear what they should be building next (e.g.
@@
 In the next section, we revisit the strict F1 metric, and define a fairer variant that accounts for the fact that some empty-board test items have multiple valid interpretations.
-3.1 Annotating the test set
-We manually annotate items in the test set so that we can identify cases that have to be distinguished for evaluation purposes.
-Annotators were asked to annotate all items in the same dialog in sequence.
-For each non-empty board (NEB) item, annotators were first asked whether the context is clear and has a unique interpretation , i.e.
-clearly specifies what structure to be built next, and there is only a single way of placing that structure on the board (as is typically the case when the board is not empty), or whether the context is unclear, i.e.
-it is unclear what the Builder should do next, either because more information is required, or no new instruction was given.
-For empty board (EB) items, annotators could also indicate that the context is clear and has multiple interpretations , i.e.
-it is clear what structure should be built, even though there are multiple possible placements or orientations within the build region, as is often the case when the board is empty.
-If the context is clear (with unique or multiple interpretations), a second question asked whether the built structure is correct or incorrect.
-Details about the annotation setup are provided in the appendix (Section 14).
-We now illustrate the resulting categories of BAP items with a few examples.
-Clear context with unique interpretation and correct structure (Figure 5)
-It is clear what structure B should build next, only one such structure can be built, and the human Builder actually built that structure (Figure 5(a)).
-This can also arise when the board is empty (Figure 5(b)), since B is first made to stand at the bottom right corner of the board, and A clearly specifies the locations of the two blocks to be placed, given B’s new orientation.
-Clear context with unique interpretation and incorrect structure (Figure 6)
-Here, it is clear what structure B should build next (and only one such structure can be built), but B’s actions did not result in that structure, perhaps because B got interrupted by A, or because B made a genuine mistake.
-In Figure 6(a), A wants B to place two red blocks directly below the middle block of the top of the arch, but B only places one of those blocks.
-In Figure 6(b), A wants B to place a red block diagonally above (“on top and in front of") the topmost green block, but B places a red block in the wrong location.
-In the next BAP item from that dialog (Figure 6(c)), A tries to correct B’s mistake, but B now places two red blocks, only one of which should be there in the final structure.
-Clear context with multiple interpretations and correct structure (Figure 7)
-When the board is empty, it is often clear what structure B should build next, but the placement and orientation of that structure may not have been uniquely specified.
-In this case, the context is clear, although it can be (correctly) interpreted in multiple ways.
-Unclear context (Figure 8)
-Finally, some action sequences occur in contexts in which it is unclear what the Builder should do next.
-In Figure 8(a), it is unclear what exact location A implies by “the side of the red".
-In Figure 8(b), A only signals they had made an error, but it is unclear what the error is or what its recency is and A does not provide any information that would allow B to correct the error.
-((a)) Item 1
-((b)) Item 2
-Figure 5: Clear context with unique interpretation and correct structure
-((a)) Item 1
-((b)) Item 2
-((c)) Item 3
-Figure 6: Clear context with unique interpretation and incorrect structure
-Figure 7: Empty board: Clear context with multiple interpretations and correct structure
-((a)) Item 1
-((b)) Item 2
-Figure 8: Unclear context
-3.2 Creating a clean test set
-To obtain a clean BAP test set, we remove all items with unclear contexts, as the the correctness of the built structure cannot (and should not) be assessed for the BAP task.
-We also automatically fix/replace incorrect structures that can be fixed in such manner, and remove the items with structures that can’t be fixed.
-This results in a new, BAP v2 test set that contains only items with clear contexts and correct structures, and thus suitable for fairer BAP evaluation.
-Fixing incorrect structures by looking ahead to the next correct structure
-If the Builder made a mistake or their action sequence was interrupted by an utterance, the built structure of a BAP item may be incorrect.
-E.g.
-in Figure 9(a) B places only one instead of two blocks.
-However, the next BAP item (Figure 9(b)) has a correct structure, and its structure is the one that should have been built in 9(a) because 9(b) completes the truncated action sequence.
-We can therefore replace the structure in 9(a) with the one in 9(b).
-Figure 10 shows a sequence of three BAP items: in 10(a) B misinterprets A and makes an error, in 10(b) B realizes the mistake post A’s correction and proceeds to correct it, but A interrupts them, and in 10(c), the mistake is finally rectified post the interruption.
-Items 10(a) and 10(b) with incorrect structures should therefore be amended to result in the correct structure built in 10(c).
-Manual analysis shows that sequences of one or more incorrect items (with clear contexts) that are immediately followed by a correct item (with clear context) can be fixed by replacing their built structures with the built structure of that subsequent correct item.
-Those items with incorrect structures that conform to this pattern are thus fixed automatically, and the remaining that can’t are excluded from the final test set.
-((a)) Item $i$ has an incorrect built structure
-((b)) Item $i+1$ has a correct built structure
-Figure 9: An item with an incorrect built structure (9(a)) that can be fixed by the correct built structure in the following item (9(b))
-((a)) Item $i$ has an incorrect built structure
-((b)) Item $i+1$ has an incorrect built structure
-((c)) Item $i+2$ has the correct built structure for items $i$, $i+1$ and $i+2$
-Figure 10: A sequence of items with incorrect built structures (10(a), 10(b)) that can be fixed by the following item’s correct built structure (10(c))
-Statistics of the clean BAP v2 test set
-The BAP v2 test set consists of 1,155 BAP items, all of which have a clear context and a correct built structure.
+3.1 Identifying suitable BAP test items
+To distinguish the cases mentioned above, we manually annotate all BAP items in the test set (for details, see Appendix Section 12).
+For each non-empty board (NEB) item, annotators were asked whether the context is clear and has a unique interpretation (i.e.
+it is clear what structure to be built next, and where to place it on the board), or whether the context is unclear (i.e.
+it is unclear what to build next, e.g.
+because more information is required, or no new instruction was given).
+For empty board (EB) items, annotators could also indicate that the context is clear and has multiple interpretations, (i.e.
+it is clear what structure should be built, even though there are multiple possible placements or orientations of that structure on the empty board.) If the context is clear, a second question asked whether the built structure is correct or incorrect.
+Suitable BAP test items
+Items whose contexts are clear and whose structures are correct can clearly be used for evaluation purposes (Figure 4).
+If the board is non-empty, this requires contexts to have a unique interpretation, but if the board is empty, Builders can chose any of typically many correct interpretations.
+Unsuitable BAP test items
+Items are unsuitable for evaluation if the context is unclear (e.g.
+because the Architect gave unclear instructions, Figure 5) or if the structure is incorrect (e.g.
+the Builder made mistakes, Figure 6).
+Fixable BAP test items
+Manual analysis shows that sequences of one or more incorrect built structures that are immediately followed by a correct structure can be fixed by replacing the incorrect structures with the built structure of that subsequent correct item (Figure 7).
+Such cases occur typically because the Builder’s (correct) action sequence was interrupted by an utterance, or because the Architect issued a quick correction that the Builder immediately executed.
+3.2 Statistics of the clean BAP v2 test set
+To obtain our final evaluation set, we start with the original BAP test set of 1,616 BAP items, but correct the structures of 149 items that we were able to identify and fix automatically, and remove 461 items that are unsuitable for evaluation purposes (418 with unclear contexts and 43 with incorrect structures that could not be fixed).
+This results in a new, BAP v2 test set that consists of 1,155 BAP items, all of which have a clear context and a correct built structure.
 1,071 of these items have a unique correct interpretation, and 84 of the empty board (EB) items have multiple correct interpretations.
-This set was obtained from the original BAP test set of 1,616 BAP items by removing 461 items that should not be used for evaluation (418 with unclear contexts and 43 with incorrect structures that could not be fixed), and by fixing the structures of 149 items whose original structures were incorrect.
 4 Revisiting BAP Evaluation: Metrics
@@
 4.1 Fairer F1
-The clean v2 test set consists only of BAP items with a clear context and correct structures, but 84 of the EB items (7.3% of the v2 test set) do not have a unique interpretation, since their structure can be placed and oriented in many different ways (see Figure 7 above).
+The clean v2 test set consists only of BAP items with a clear context and correct structures.
+84 of the EB items (7.3% of the v2 test set) do not have a unique interpretation, since their structure can be placed and oriented in many different ways (Section 3.1).
 The v2 test set can therefore be split up into items with a unique correct interpretation ($\mathcal{U}$) and items with multiple correct interpretations ($\mathcal{M}$).
@@
 That is, a fairer F1 metric computes a strict F1 score for the predicted action sequences $A_{M}^{\mathrm{net}}$ for any item in $\mathcal{U}$, because the reference sequences for these items yield the only correct structures for these items, but for any item in $\mathcal{M}$, it computes a strict F1 score for the optimally aligned predicted action sequences $\widetilde{A}_{M}^{\mathrm{net}}$ because this item has multiple correct interpretations that only differ from each other in location and orientation.
-Thus, in the example in Figure 7, any predicted orange 3x3 U-shaped structure placed on the ground will receive a fairer F1 score of 1.0.
+Thus, in the example in Figure 4 (bottom), any predicted orange 3x3 U-shaped structure placed on the ground will receive a fairer F1 score of 1.0.
 4.2 Auxiliary metrics
@@
 […]
-[…]
+$\textbf{Location equivalence:}~(t,c,\langle x,y,z\rangle)\equiv_{\textrm{Loc}}(t^{\prime},c^{\prime},\langle x^{\prime},y^{\prime},z^{\prime}\rangle)$ $\iff\langle x,y,z\rangle=\langle x^{\prime},y^{\prime},z^{\prime}\rangle$
 (Note that the condition for Location equivalence is effectively the same as $t=t^{\prime}~\textrm{and}~\langle x,y,z\rangle=\langle x^{\prime},y^{\prime},z^{\prime}\rangle$).
 Accordingly, we can transform both net action sets $A_{M}^{\textrm{net}}$ and $A_{H}^{\textrm{net}}$ into Type, Color or Location multi-sets:
-$M^{\textrm{Type}}$ =$=\{t\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
-$M^{\textrm{Color}}$ =$=\{(t,c)\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
-$M^{\textrm{Loc}}$ =$=\{\langle x,y,z\rangle\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
-which yields Type, Color and Location (strict) precision, recall and F1 scores by comparing the sizes of the corresponding multisets and their intersections (we only show Type F1 below, as Color and Location F1 are similarly defined):
-Strict Type $\textbf{Strict Type Precision}~{}P_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{M}^{\textrm{Type}}|}$
-Strict Type $\textbf{Strict Type Recall}~{}R_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{H}^{\textrm{Type}}|}$
-[…]
-Intuitively, the Type F1 metric quantifies how good a model is at understanding what type of actions to perform – placements, or removals, or both, and how many of each type.
-The Color F1 metric narrows down Type F1 a bit more to factor for colors as well.
-It quantifies how good a model is at understanding the colors of blocks that need to be placed or removed.
-The Location F1 metric quantifies how good a model is at understanding the specific locations of blocks that need to be placed or removed (without regard to color), i.e., spatial reasoning.
-Also, as mentioned earlier, these metrics are analogous to the strict F1 metric.
-They can then be used to define their fairer counterparts in the same way the strict F1 was used to define the fairer F1.
-Henceforth, we will use Type, Color and Location F1 to refer to their fairer counterparts.
+$M^{\textrm{Type}}$ $=\{t\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
+$M^{\textrm{Color}}$ $=\{(t,c)\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
+$M^{\textrm{Loc}}$ $=\{\langle x,y,z\rangle\mid(t,c,\langle x,y,z\rangle)\in A^{\textrm{net}}\}$
+which yields (strict) Type, Color and Location precision, recall and F1 scores by comparing the sizes of the corresponding multisets and their intersections (we only show Type F1 below, as Color and Location F1 are similarly defined):
+$\textbf{Strict Type Precision}~P_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{M}^{\textrm{Type}}|}$
+$\textbf{Strict Type Recall}~R_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{|(M_{M}^{\textrm{Type}}\cap M_{H}^{\textrm{Type}})|}{|M_{H}^{\textrm{Type}}|}$
+$\textbf{Strict Type F1}~F1_{\textrm{Type}}(A_{M}^{\textrm{net}},A_{H}^{\textrm{net}})$ $=\frac{2\cdot P_{\textrm{Type}}\cdot R_{\textrm{Type}}}{P_{\textrm{Type}}+R_{\textrm{Type}}}$
+Fairer counterparts to these strict metrics are defined analogously to the overall metric, and we will use the fair variants in the remainder of this paper.
+Intuitively, the Type metric captures how good a model is at understanding what type of actions to perform – placements, or removals, or both, and how many of each type.
+The Color metric is stricter than he Type metric, in that it quantifies how good a model is at understanding the colors of blocks that need to be placed or removed.
+The Location metric quantifies how good a model is at understanding the specific locations of blocks that need to be placed or removed (without regard to color), i.e., it helps to evaluate spatial reasoning.
 Shape F1
-To evaluate how good a Builder model is at understanding the "shape" in which the required blocks need to be placed or removed, we define a metric, Shape F1, that is agnostic to the absolute $\langle x,y,z\rangle$ locations of the required blocks, but only cares that they be placed/removed in the right locations relative to each other.
-Essentially, it compares net actions modulo their exact placement and orientation.
-Therefore, similar to the methodology used for fairer F1 (Section 4.1), we optimally align the model’s net actions $A_{M}^{\mathrm{net}}$ against the human reference net actions $A_{H}^{\mathrm{net}}$ (instead of the corresponding resulting built structures themselves like was done for fairer F1), yielding aligned net actions $\hat{A}_{M}^{\mathrm{net}}$ that can now be compared against $A_{H}^{\mathrm{net}}$ by our original strict F1 metric.
+To evaluate how good a Builder model is at understanding the structure in which the required blocks need to be placed or removed, we define a metric, Shape F1, that is agnostic to the absolute $\langle x,y,z\rangle$ locations of the required blocks, but only cares that they be placed/removed in the right locations relative to each other.
+Essentially, this metric compares net actions modulo their exact placement and orientation.
+Similar to the methodology used for fairer F1 (Section 4.1), where we align the resulting built structures, we optimally align the model’s net actions $A_{M}^{\mathrm{net}}$ against the human reference net actions $A_{H}^{\mathrm{net}}$, yielding aligned net actions $\hat{A}_{M}^{\mathrm{net}}$ that can now be compared against $A_{H}^{\mathrm{net}}$ by our original strict F1 metric.
 This is done for all items in the v2 test set (and not just the ones with multiple interpretations like was done for fairer F1).
-We will see examples of Shape F1 and the other auxiliary metrics later in Sections 7.2 and 8.1.
-4.3 Micro vs Macro F1
-Current BAP evaluation reports the micro-averaged F1 (micro F1) over all action sequences in the test data (Section 2.4), reflecting a model’s quality in predicting individual actions.
-However, since the BAP task requires sequence prediction, it’s equally important to assess the model’s quality in predicting entire action sequences using the macro-averaged F1 (macro F1).
-Intuitively, this is also particularly relevant for Shape F1, which measures the model’s performance on a broader sequence-level property (the "shape" of the net actions).
-Therefore, hereafter, we will report both micro and macro F1.
+We will see examples of Shape F1 and the other auxiliary metrics in Sections 7.2.
 4.3 BAP v2 evaluation benchmark
-The above metrics enable a fairer, robust, and more insightful evaluation on the clean v2 test set.
-We use the fairer F1 metric to report both micro and macro F1 scores in Tables 1 and 2 respectively.
-For each, we include both auxiliary and overall F1 metrics, and report them on the overall/full test set and separately for the EB and NEB subsets as well.
-This forms the BAP v2 evaluation benchmark.
-Our approach also enables more accurate evaluation within the EB and NEB categories, offering a clearer understanding of model performance across these two distinct scenarios.
-This division is important for the following reasons:
-  • EB examples are generally the easiest for the BAP task, making EB performance a basic competency or sanity test, and it also reflects a model’s maximum potential quality.
-  • EB performance is also an estimate of the ability of a model to begin a game accurately – an elementary yet important trait to measure.
-Moreover, in an interactive/online setting, from a user (human Architect) experience perspective and considering the potential for subsequent cascading errors, mistakes made in EB scenarios can potentially be costlier than those in NEB, as EB marks the start of a game (even though EB examples in the data are far fewer than NEB).
-Quantifying EB performance will help provide a sense for how a model can potentially fare in an interactive setting wrt this aspect.
-We begin with the micro scores in Table 1 for direct comparison with our prior work.
-Recall that the baseline model achieves an F1 score of 21.1% under the strict F1 metric.
-This increases to 27.3% under the fairer F1 metric (the overall F1 on the overall dataset), a notable difference of 6.2 percentage points and a 29.4% improvement.
-This highlights the importance of the high quality cleaner test data and the fairer F1 metric in providing a fairer reflection of model capability.
-On the overall dataset, the model achieves reasonably good F1 scores of 65.2% for Type and 62.4% for Color, but only 27.7% for Location F1, which is also very close to the overall F1 of 27.3%.
-This suggests that spatial reasoning (Location F1) is the key bottleneck for high performance on this task.
-As expected, the model performs better on EB than NEB across all metrics.
-Recall that the number of EB items are fewer than NEB items though (and the metrics on the overall dataset will therefore be more influenced by NEB).
-Next, we turn to the macro scores in Table 2.
-While all values increase compared to the micro scores, the overall trends remain the same.
-The overall F1 on the overall test set increases from 27.3% (micro) to 28.1% (macro).
-The model achieves a Shape F1 of 36.3% (micro) and 41.1% (macro) on the overall test set, with the macro score providing a more intuitive measure in this context, as discussed in Section 4.3.
-Hereafter, we will use the v2 evaluation to report model performance.
+The above metrics enable a fairer, robust, and more insightful evaluation.
+We validate this on the GRU-based baseline model for the BAP task (Section 2.6).
+Table 1 reports the full battery of scores included in the BAP v2 evaluation benchmark, i.e.
+overall and auxiliary F1 scores on the overall test set as well as for the empty and non-empty board (EB and NEB) subsets.
+We note that the overall F1 score has risen from 21.1 reported in Jayannavar, Narayan-Chen, and Hockenmaier (2020) to 27.3 under our newer, fairer evaluation, or 53.2 for EB and 24.2 on the (more frequent) NEB items.
+The distinction between EB vs.
+NEB is important since EB examples are generally easier, making EB performance a basic competency or sanity test, and a reflection of a model’s maximum potential quality.
+EB performance is also an indication of a model’s ability of a model to begin a game accurately – an elementary yet important trait to measure, especially for interactive/online settings where errors might cascade throughout a game.
+The auxiliary metrics provide further insight: in particular, we see that the model achieves reasonably good F1 scores of 65.2% for Type and 62.4% for Color, but only 27.7% for Location F1, which is also very close to the overall F1 of 27.3% (similar results hold both for the EB and NEB setting).
+This suggests that spatial reasoning is the key bottleneck for high performance on this task.
 Dataset Type Color Location Shape Overall
@@
 Overall 65.2 62.4 27.7 36.3 27.3
-Table 1: Micro F1 scores for the baseline model under the v2 benchmark
-Dataset Type Color Location Shape Overall
-EB 80.9 79.8 61.5 64.9 61.2
-NEB 73.1 69.9 25.7 39.0 25.2
-Overall 73.8 70.7 28.5 41.1 28.1
-Table 2: Macro F1 scores for the baseline model under the v2 benchmark
-5 Revisiting BAP Data: Generating Synthetic Data
-One key challenge of the BAP task is the limited training data, despite the data augmentation (Section 2.5).
-We aim to avoid simply collecting more data like the MDC, due to scalability issues and the desire to stay aligned with many realistic scenarios, where data collection for complex situated dialogue tasks, like the MCBT, can be impractical/expensive, warranting exploration of alternative methods.
-Additionally, we observe that the baseline model struggles primarily with spatial reasoning, as reflected in its Location F1 score.
-These reasons motivate the need for additional synthetic training data that, while naturally simpler than the MDC, remains rich in spatial relations and referential expressions to help train more robust models.
+Table 1: F1 scores for the GRU baseline model on the v2 benchmark
+5 Revisiting BAP Training Data: Generating Synthetic Data
+Synthetic Target Structures
+Random Shape-Based
+Refer to caption Refer to caption Refer to caption
+Figure 8: Synthetic target structures
+Synthetic BAP items
+(A) Block-based dialogues for random targets ($D_{\mathrm{r}}$)
+(B) Block-based dialogues for shape-based targets ($D_{\mathrm{sb}}$)
+(C) Shape-based dialogues for shape-based targets ($D_{\mathrm{ss}}$)
+Figure 9: We generate three different types of synthetic BAP items, resulting in datasets $D_{\mathrm{r}}$, $D_{\mathrm{sb}}$ and $D_{\mathrm{ss}}$
+Our process to generate synthetic BAP items
+The synthetic BAP item generated by the process above
+Figure 10: Our process to generate synthetic BAP items (top) and a BAP item generated by that process (bottom)
+Feature MDC ($D_{\mathrm{mc}}$) $D_{\mathrm{r}}$ $D_{\mathrm{sb}}$ $D_{\mathrm{ss}}$
+Target Structure Human-designed, complex and meaningful (e.g., furniture, animals).
+Dynamically and randomly generated; typically disordered.
+Predefined composites of 6 elementary shapes.
+Predefined composites of 3 simpler shapes.
+Architect Planning Complex, strategic & adaptive human planning with hierarchical decomposition of target structures based on goals of efficiency, naturalness in instruction giving, etc.
+Minimalistic; plans single random block placement/removal per turn.
+Mid-level; plans one shape at a time, decomposed into block-level instructions.
+High-level; plans construction of one full shape instance per turn.
+Instruction Abstraction Multi-level: from high-level concepts (“flower") to low-level commands.
+Low-level; single block references only.
+Low-level; but more abstract than $D_{\mathrm{r}}$; decomposes shapes into block-level actions.
+Mid-level; refers to entire shapes (“row", “plane").
+Semantic Naming Frequent use of names for commonplace objects (“chicken"), shapes & user-defined concepts (“r-windows").
+N/A (block-level instructions only).
+N/A (block-level instructions only).
+Limited; template-based names for shapes (“pillar", “line").
+Spatial Relations Highly diverse, complex, and nuanced (“parallel", “symmetrical", “hanging").
+Simple 1D/2D relations anchored to arbitrary reference blocks.
+Simple 1D/2D relations anchored to the “last placed block".
+All 1D/2D/3D relations to specify shape’s starting point and Builder-anchored relations (“going to the left of you") to specify orientation.
+Floating Blocks Present; instructions range from single-shot commands to step-by-step guidance.
+Present; single-shot commands only.
+Present; single-shot commands only.
+Present; single-shot commands only.
+Context Clarity Variable; includes unclear/ambiguous contexts (Section 3). Always clear & unambiguous. Always clear & unambiguous. Always clear & unambiguous.
+Dialogue Synchronicity Asynchronous; mutual interruptions by A and B are possible.
+Synchronous; discrete, structured turns.
+Synchronous; discrete, structured turns.
+Synchronous; discrete, structured turns.
+Structural Correctness Contains correct & incorrect structures (due to mistakes/interruptions; Section 3). Always correct. Always correct. Always correct.
+Dialogue Act Diversity High; wide mix (instructions, confirmations, corrections, social chat, etc.).
+Low; dominated by instructions & short clarification exchanges.
+Low; dominated by instructions & short clarification exchanges.
+Low; dominated by instructions and short clarification exchanges.
+Builder Initiative High; B actively verifies, suggests, gives updates and extrapolates.
+Low; limited to simple clarification questions.
+Low; limited to simple clarification questions.
+Low; limited to simple clarification questions.
+Clarification Exchanges Varied and natural questions about ambiguities ( “is it flat?").
+Simulated; on missing block color or location.
+Simulated; on missing block color or location.
+Simulated; on missing color, location, size, or orientation.
+Referring Expressions Highly varied; includes implicit arguments and ellipsis (“two more in the same direction").
+Simple; disambiguate blocks by color and relative position (“leftmost red block").
+Simple; refers to the last placed block as spatial anchor.
+Simple; refers to last placed block of previous shape as spatial anchor.
+Builder’s Low-Level Planning Required & involved; scales with instruction abstraction.
+Minimal; executes single-block instructions.
+Minimal; executes block-level instructions.
+Required but simple; uses heuristics (e.g., column-by-column, row-by-row, or zigzag patterns) for shape instructions.
+Table 2: Detailed qualitative comparison of MDC and synthetic datasets. For all rows, MDC features are a superset of the ones for the synthetic datasets.
+Despite the data augmentation (Section 2.3), the BAP task suffers from relatively limited training data, but collecting a significantly larger amount of human data would have been prohibitively expensive and impractical.
+More importantly, while the MDC dialogues are far more linguistically rich and varied than our synthetic data, the baseline models struggle even with basic spatial understanding and reasoning (as reflected in the Shape and Location F1 scores).
+In order to achieve comprehensive coverage of various spatial relations, we therefore opted to create synthetic training data that would allow us to train our models in these basic skills.
+Specifically, we distinguish between two different types of target structures ––random and shape-based (Figure 8) –, and three different types of dialogues ––block-based for random or shape-based targets, or shape-based for shape-based targets–– (Figure 9), described in detail below.
+We design novel Minecraft target structure and dialogue simulators that emulate the MCBT, and use them to generate a set of three synthetic dialogue datasets with corresponding target structures.
+Each dataset is modeled after the MDC and, while understandably simpler, preserves essential elements of realistic human behavior observed in the MDC as much as possible (including, e.g.
+clarification exchanges, Architect A ’s and Builder B ’s planning perspectives, dialogue flow, etc.).
+While we specifically emphasize spatial relations (as it is a core motivation) and representing examples of different complexities, we also vary the level of instruction abstraction, the complexity of referring expressions, the complexity of target structures, etc.
+across these datasets.
+Each is thus unique, contributing to the diversity and richness of the overall set.
 Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs.
-(We, in fact, also illustrate the impact of this data on training a simple LLM and transformer-based model in Section 6.3.)
-We design novel Minecraft dialogue and target structure simulators that emulate the MCBT, and use them to generate a set of three synthetic dialog datasets with corresponding target structures .
-Each dataset is modeled after the MDC and is understandably simpler, while preserving essential elements of realistic human behavior observed in the MDC as much as possible, such as the language, clarification exchanges, Architect A ’s and Builder B ’s planning perspectives, dialog flow, etc.
-Across these datasets, we vary the level of instruction abstraction, the complexity of spatial relations and referential expressions, the complexity of target structures, etc.
-Special emphasis is laid on spatial relations as that is a core part of our motivation here.
-Each dataset is thus unique, contributing to the overall diversity and richness of the overall set.
-Note that although our primary focus is on BAP, the simulators and data introduced here are applicable to other MCBT subtasks as well (e.g.
+And although our primary focus is on BAP, the simulators and data introduced here are applicable to other MCBT subtasks as well (e.g.
 the AUG subtask; Section 2.1.2), supporting advancements across the broader MCBT ecosystem.
 We discuss this in more detail in Section 8.2.
-5.1 General dialog simulation framework
-5.1.1 Desiderata
-Similar to the target structures and game logs in the MDC, the following components of synthetic dialogues need to be generated:
-  • The target structure
-  • The dialogue/game – a chronological sequence that interleaves the following elements:
-    • A dialogue between A and B, consisting of utterances that include instructions, potential clarification exchanges, etc.
-    • B’s position and orientation at appropriate points in the game, i.e., when an utterance or action is recorded (Grounded elements of the dialog, e.g.
-spatial relations, are always wrt this frame of reference.)
-    • B’s actions, which reflect the evolving structure being built
-5.1.2 High-level Algorithm
-As described in Section 2.1, the MCBT task is asynchronous and loosely structured.
-The MDC dialogs are real human-human game logs collected in a low-stakes, free-flowing setting with minimal constraints on participants’ language and actions.
-To enable tractable simulation and facilitate simpler synthetic data generation, we make some simplifications and structure the task as follows.
-Our dialog simulators generate synthetic dialogs that meet the desiderata, following a general framework.
-At a high level, it is an iterative procedure that continues until the target structure is built.
-The target structure may be provided as input or generated dynamically.
-In the former case, a separate simulator generates target structures.
-Each iteration produces one dialog segment—A’s instruction followed by B’s actions in response, with optional clarification exchanges and confirmations—leading to the construction of a part of the target structure.
-We are essentially simulating A and B, including their latent planning processes.
-Each iteration proceeds as shown below.
-(We use Figure 11 as an example.
-It illustrates a BAP item $(\mathcal{H}^{t-1},S^{t-1},D^{t},A^{t},S^{t})$ from one of our synthetic datasets (recall notation from Section 2.2).
-In this iteration, the dialog and actions following $S^{t-1}$ are simulated to produce the next structure, $S^{t}$ – specifically, simulating $D^{t}$, $A^{t}$, and $S^{t}$.)
+5.1 Our dialogue simulation framework
+As described in Section 2.1, the MCBT is characterized by asynchronous, loosely structured dialogue between human players with minimal constraints on their language and actions.
+To make the simulation of this complex interaction tractable and to facilitate a simplified data generation process, we introduce a simpler and more structured framework.
+Like the original MDC data, for each game, a target structure and a complete log that interleaves the dialogue between A and B (consisting of instructions, potential clarification exchanges), B’s actions (which reflect the evolving structure being built), and B’s changing position and orientation (so that spatial references in the utterances can be correctly interpreted).
+Our simulators generate grounded synthetic dialogues that meet these desiderata by iteratively simulating the MCBT until the target structure is built.
+Each iteration of the algorithm produces a BAP item, $(\mathcal{H}^{t-1},S^{t-1},B^{t},D^{t},A^{t},S^{t})$ (as defined in Section 2.2), by simulating the dialogue $D^{t}$ and actions $A^{t}$ following $S^{t-1}$ to produce the next structure, $S^{t}$.
+Each item consists of the specification of B’s current position $B^{t}$, a non-empty dialogue segment $D^{t}$A’s instruction, optional clarification exchanges with B, and confirmations– followed by B’s actions $A^{t}$ that alter the built structure from $S^{t-1}$ to $S^{t}$, thus constructing part of the target structure.
+$\mathcal{H}^{t-1}$ and $S^{t-1}$ represent the current game context that is input to this iteration of the algorithm.
+The target structure for the dialogue is handled in one of two ways: it is either provided as a predefined input to the algorithm (generated by a separate target structure simulator) or generated dynamically.
+If predefined, each iteration involves selecting the next blocks to place from this target.
+If generated dynamically, each iteration involves choosing blocks randomly to place or remove, and the final constructed configuration is designated as the target structure for the dialogue post-hoc.
+Each iteration proceeds according to the steps detailed below, which is also visualized in the running example in Figure 10 (bottom).
   1. 1.
-    Architect Planning: First, A’s plan for the next steps is simulated, including which blocks to place or remove and their order if multiple blocks are involved.
-A reference block is chosen when necessary, enabling A to describe the new block(s) with appropriate spatial relations wrt the reference block.
-E.g., in Figure 11, A decides to place the red floating block, using the previously placed blue block as the reference.
+    Architect Planning (Step 1): The simulation begins by sampling the Architect’s plan for the next construction phase.
+This includes determining which blocks to place or remove and in what order.
+A reference block is chosen, when necessary, enabling A to describe the location of the new block(s) relative to the reference block in the subsequent dialogue $D^{t}$.
+In our running example, A decides to place a red floating block, using the previously placed blue block as a reference (diagonally from it at $(x_{1},y_{1},z_{1})$.
   2. 2.
-    Builder Position and Orientation: A "valid" position and orientation for B is sampled, ensuring a frame of reference for the utterances in the dialog segment.
-B is positioned at a suitable distance from the action area and oriented to face it while adhering to world constraints.
-E.g., in Figure 11, the structure $S^{t-1}$ is rotated according to B’s gaze (yaw angle only).
-[^4] B’s position and pitch angle is omitted in our figures for clarity as noted in Section 2.2
+    Adjusting the POV to a new Builder Position and Orientation (Step 2): As noted in Section 2.1.1, A and B can move around freely, but A views the structure from behind B when giving instructions.
+As a result, A instructions (including spatial relations, etc.) are grounded in B’s perspective.
+To emulate this, this step simulates B’s position and orientation $B^{t}$ (see Section 2.4 for a formal definition).
+A valid $B^{t}$ is sampled for B, establishing the frame of reference for the upcoming dialogue, $D^{t}$.
+$B^{t}$ is considered valid if B is at a suitable distance from the action area and oriented to face it, while respecting certain world constraints.
+E.g., Steps 2A and 2B show $B^{t}$ being chosen.
+In Step 2C, the grid is also rotated to align with B’s point-of-view and to view $S^{t-1}$ according to that POV – just for the reader’s visual clarity.
+Note that technically,
+    Step 5.1.
+    1 and
+    Step 5.2.
+    2 are independent and can be performed in any order.
   3. 3.
-    Architect Instruction: An instruction from A is synthesized, incorporating necessary block details (e.g.
-color), spatial relations, and referential expressions, with template-based lexical and syntactic variations.
-E.g., in Figure 11, A instructs, "put a floating block …".
+    Generating the Main Dialogue (Step 3):
+    1. (a)
+      Dialogue Style (Step 3A): The simulation first determines the dialogue style for the turn.
+It randomly chooses between a straightforward, complete instruction and, with much lower probability, an instruction that intentionally omits information to elicit a clarification question from the Builder.
+    2. (b)
+      Dialogue Generation (Step 3B): Next, the dialogue $D^{t}$ is synthesized using templates that incorporate lexical and syntactic variations.
+If a direct instruction was chosen, the utterance includes all necessary details (e.g., color, spatial relations (5.2.3), referential expressions).
+If a clarification was chosen, a key detail is omitted, and a corresponding question-answer exchange between B and A is generated.
+We show both possibilities for $D^{t}$ (omitting the block’s color for the clarification case).
   4. 4.
-    Optional Clarification Exchange: Some information may be omitted from the instruction with a small probability, allowing B to ask a clarification question.
-An appropriate exchange is then generated, with B asking and A providing the missing information, with template-based lexical and syntactic variations.
-E.g., in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
+    Generating the Builder Actions (Step 4): The Builder’s actions ($A^{t}$) to execute the instruction are then simulated.
+This includes low-level planning to decide the precise order of actions and to account for temporary supporting blocks required for construction.
+These actions update the built structure, producing the new state, $S^{t}$.
+In our example, B places a red floating block to produce structure $S^{t}$, temporarily adding and removing a support block to do so.
+[^5] As noted in Section 2.2, temporary supporting blocks are not shown in our figures.
   5. 5.
-    Builder Actions and Low-Level Planning: B’s actions to add/remove blocks are then simulated, including any necessary low-level planning for deciding order of actions and accounting for temporary supporting blocks.
-This updates the built structure.
-E.g., in Figure 11, B places the red floating block to produce structure $S^{t}$, temporarily adding/removing support for the floating block.
-[^5] Temporary supporting blocks are not shown in our figures as noted in Section 2.2.
-  6. 6.
-    Optional Confirmation from Builder: Optionally, a confirmatory utterance from B, like "done," may be synthesized with a small probability.
-Figure 11: A BAP item from one of our synthetic datasets
-We implement this general framework in three distinct ways, resulting in three different simulators.
-Each simulator uniquely instantiates steps 1, 3, and 4, along with independent target structure generation, while steps 2, 5, and 6 remain largely consistent across simulators.
-For brevity, we provide only a summary of each simulator in the following sections and provide more details in the appendix (Section 11).
-5.1.3 Spatial Relations
-Given the special emphasis on spatial relations in our dataset design, we briefly outline them here. We focus on the following types of spatial relations:
-  1. 1.
+    Generating an Optional Confirmation from Builder (Step 5): Finally, the turn may (with a low probability) conclude with a confirmatory utterance from B, such as “done!".
+    The complete BAP item produced by this five-step process is shown in Figure 10 (bottom), which follows the standard visualization format used throughout this paper, so that structures like $S^{t-1}$ are rotated to align with the Builder’s POV $B^{t}$ (corresponding to the next turn and dialogue segment $D^{t}$ below it), and not with $B^{t-1}$, whereas Steps 2A-2C in the top of Figure 10 explicitly show the transition from the previous perspective ($B^{t-1}$) to the new one ($B^{t}$).
+    We developed three distinct simulators based on this general framework. While
+    Step 5.3.
+    2,
+    Step 5.4.
+    4, and
+    Step 5.5.
+    5 are implemented largely consistently across all three, each simulator features a unique implementation of Architect Planning (
+    Step 5.6.
+    1) and Generating the Main Dialogue (
+    Step 5.7.
+    3), as well as its own target structure generation.
+For brevity, we provide only a summary of each simulator in the following sections and provide more details for each in the appendix (Section 10).
+    5.2 Generating Grounded Aspects of the Synthetic Dialogs
+    We now briefly elucidate some key details of
+    Step 5.8.
+    2 and the spatial relations part of
+    Step 5.9.
+    3B, common across all three simulators.
+    5.2.1 Sampling the Builder’s Position and Orientation
+    Step 5.10.
+    2
+    For each turn in the simulation, we sample a new position and orientation for the Builder $\text{Pos}_{\textbf{B}}=(\text{loc}_{\textbf{B}},\phi_{\textbf{B}},\gamma_{\textbf{B}})$, where $\text{loc}_{\textbf{B}}=\langle x_{B},y_{B},z_{B}\rangle$ is the Builder’s location, $\phi_{\textbf{B}}$ is its pitch (vertical rotation), and $\gamma_{\textbf{B}}$ is its yaw (horizontal rotation) (Section 2.4).
+This sampling is grounded by the current state of the world, specifically the location of the reference block, $r_{p}$, which will be used as a spatial anchor in the Architect’s instruction.
+    The process begins by identifying a set of all possible locations from which the Builder has an unobstructed view of the reference block.
+We then sample the Builder’s position, $\text{loc}_{\textbf{B}}$, uniformly from this feasible set.
+With the position established, we calculate the optimal yaw and pitch angles, $(\gamma_{\textbf{B}}^{*},\phi_{\textbf{B}}^{*})$, that would orient the Builder to look directly at the reference block from its new position:
+    $\gamma_{\textbf{B}}^{*}$ $=\arctan\left(-\frac{r_{x}-x_{B}}{r_{z}-z_{B}}\right)$
+    $\phi_{\textbf{B}}^{*}$ $=\arctan\left(\frac{y_{B}-r_{y}}{\sqrt{(x_{B}-r_{x})^{2}+(z_{B}-r_{z})^{2}}}\right)$
+    To introduce natural variation and avoid perfectly robotic behavior, the Builder’s final orientation is sampled from a normal distribution centered on these optimal angles:
+    $\gamma_{\textbf{B}}$ $\sim\mathcal{N}(\gamma_{\textbf{B}}^{*},h/2)$
+    $\phi_{\textbf{B}}$ $\sim\mathcal{N}(\phi_{\textbf{B}}^{*},v/2)$
+    where $h$ and $v$ are the Builder’s horizontal and vertical Field of View (FOV), respectively.
+We assume a “normal” game setting and 16:9 ratio of FOV, i.e., $h=102.4^{\circ}$ and $v=70^{\circ}$.
+[^6] https://minecraft.fandom.com/wiki/Options This ensures the Builder is generally facing the area of interest without being perfectly fixated on it.
+    5.2.2 Grounding Spatial Relations in B’s Perspective (
+    Step 5.12.
+    3B)
+    Computing perspective coordinates
+    Once the Builder’s POV for the turn is established, we can generate spatial descriptions that are naturally grounded in this perspective.
+This is achieved by converting absolute world coordinates into ‘perspective’ coordinates (Jayannavar, Narayan-Chen, and Hockenmaier, 2020), which remap the 3D space relative to the Builder’s position and orientation.
+Any absolute location $\text{loc}=\langle x_{c},y_{c},z_{c}\rangle$ is transformed into perspective coordinates $\text{loc}^{\prime}=\langle x_{c}^{\prime},y_{c}^{\prime},z_{c}^{\prime}\rangle$ by first translating the origin to the Builder’s location $\text{loc}_{\textbf{B}}$, and then applying rotations to account for the Builder’s yaw ($\gamma_{\textbf{B}}$) and pitch ($\phi_{\textbf{B}}$): [^7] The rotation matrices for pitch ($P$) and yaw ($Y$) are defined as: $P\!=\!\left(\begin{smallmatrix}1&0&0\\ 0&\cos\phi_{\textbf{B}}&\sin\phi_{\textbf{B}}\\ 0&-\sin\phi_{\textbf{B}}&\cos\phi_{\textbf{B}}\end{smallmatrix}\right)$ and $Y\!=\!\left(\begin{smallmatrix}\cos\gamma_{\textbf{B}}&0&-\sin\gamma_{\textbf{B}}\\ 0&1&0\\ \sin\gamma_{\textbf{B}}&0&\cos\gamma_{\textbf{B}}\end{smallmatrix}\right)$.
+    $\langle x_{c}^{\prime},y_{c}^{\prime},z_{c}^{\prime}\rangle=P\cdot Y\cdot\langle x_{c}-x_{B},y_{c}-y_{B},z_{c}-z_{B}\rangle$
+    Computing spatial relations
+    We leverage these perspective coordinates to generate phrases describing the position of the next block to be placed, $p$, relative to the reference block, $r$.
+We first convert both their absolute positions, $p_{p}$ and $r_{p}$, into perspective coordinates, yielding $p_{p}^{\prime}$ and $r_{p}^{\prime}$.
+The choice of directional words (e.g., "left/right" vs.
+"front/behind") is determined by comparing the difference between these new coordinates along the axes.
+    For example, to determine the primary horizontal relation, we compare the magnitude of the difference along the Builder’s perspective x-axis (left/right) with the difference along the z-axis (front/behind).
+If $|p_{x}^{\prime}-r_{x}^{\prime}|>|p_{z}^{\prime}-r_{z}^{\prime}|$, the most salient direction is left or right.
+The specific term is chosen based on the sign of the difference: if $p_{x}^{\prime}>r_{x}^{\prime}$, the instruction would be "to the left of" the reference block.
+Conversely, if the difference along the z-axis is more significant, the instruction would use "in front of" or "behind".
+This method provides a deterministic way to generate simple and intuitive spatial language that is directly tied to what the Builder would be seeing from its simulated vantage point.
+    5.2.3 Types of Spatial Relations
+    We generate the following types of spatial relations:
+    1. (a)
       Relative to a Reference Block: We can describe the location of a target block in relation to another block, referred to as the reference block.
@@
 This leads to the following spatial relations (all of them are considered with respect to B’s POV/frame of reference):
-    1. (a)
+      1. i.
         1D: The target block lies in one of the rows, differing from the reference block in only one dimension.
-This results in 6 basic spatial relations: left, right, top, bottom, front, and behind (e.g., "to the right of …", "behind …").
-    2. (b)
+This results in 6 basic spatial relations: left, right, top, bottom, front, and behind (e.g., “to the right of …", “behind …").
+      2. ii.
         2D: The target block lies in one of the quadrants, differing in exactly two dimensions.
 This results in 12 possible combinations of the 1D relations, such as top+right, bottom+left, etc.
-(e.g., "to the top right of …").
-    3. (c)
+(e.g., “to the top right of …").
+      3. iii.
         3D: The target block lies in one of the octants, differing in all three dimensions.
-This results in 8 possible combinations of the 1D relations, such as top+right+front (e.g., "one to the left, one above, and one block behind …").
+This results in 8 possible combinations of the 1D relations, such as top+right+front (e.g., “one to the left, one above, and one block behind …").
       These 1D/2D/3D relations apply when one needs to specify spatial relations between blocks or substructures/shapes.
-  2. 2.
-    Using B as a Spatial Anchor: Spatial relations can also be described using B as a spatial anchor directly (e.g., "going up and to the left of you").
+    2. (b)
+      Using B as a Spatial Anchor: Spatial relations can also be described using B as a spatial anchor directly (e.g., “going up and to the left of you").
 These only apply when one needs to specify the direction in which to build a certain substructure/shape, e.g., a diagonal.
-We will see concrete examples of these in the following sections.
-5.2 Random target structures and dialogs
-Dialogs
-The dialogs correspond to random target structures generated dynamically during the dialog simulation (Section 5.1.2) which runs for a (bounded) random number of iterations.
-Blocks are placed/removed randomly, and one at a time in each iteration, with limited planning on A ’s part for block placement.
-As a result, the target structure is disordered, and A ’s instructions are fairly low-level, addressing only one block at a time.
-Most of them use spatial relations relative to arbitrary blocks identified by color and location (e.g., "to the right of the red" or "to the right of the bottom-most red block").
-Figure 12 shows an example BAP item from this dataset.
-A identifies the orange block as a reference block and uses a 1D spatial relation "behind" to specify the target yellow block.
-We provide more details about the simulator in the appendix (Section 11).
-Figure 12: A BAP item from the random structure-based dialogs
-Target structures
-Target structures are generated dynamically during the dialog simulation and correspond to the final built structure at the simulation’s end. Some global constraints are enforced:
-  1. 1.
-    The structure must have at least one block on the ground.
-  2. 2.
-    The entire structure must be "connected," i.e., each block shares at least one face or edge with at least one other block.
-Figure 13 provides an example.
-Figure 13: A random target structure
-5.3 Shape-based target structures and dialogs
+    We will see concrete examples as we describe each dataset simulator in turn.
+    5.3 Dataset $D_{\mathrm{r}}$: Dialogues for random target structures
+    In this dataset, target structures are generated dynamically as a direct outcome of the dialogue simulation, corresponding to the final configuration of blocks at the end of the game.
+Dialogues are generated by running the simulation framework (Section 5) for a bounded, random number of turns.
+The key characteristics of this simulator emerge from its specific implementation of the Architect’s planning and dialogue generation steps.
+The Architect Planning (
+    Step 5.14.
+    1) process is intentionally minimalistic: at each turn, the Architect decides to place or remove a single block at a random valid location adjacent to the existing structure.
+This lack of sophisticated, high-level planning results in the dynamically generated, disordered target structures mentioned above, and directly influences the Generating the Main Dialogue (
+    Step 5.15.
+    3).
+Because the Architect considers only one block at a time, the resulting instructions are inherently low-level.
+They typically use 1D or 2D spatial relations that are anchored to arbitrary reference blocks, which are identified by their color or location (e.g., “behind" or “to the right of the red" or “to the right of the bottom-most red block", or, as in Figure 9 (top), “put a yellow block behind the orange one".
+Further details on this simulator are available in Appendix 10.
+    5.4 Datasets $D_{\mathrm{sb}}$ and $D_{\mathrm{ss}}$: Dialogues for shape-based target structures
     Target structures in the MDC are not random; they are far more ordered (e.g., a dining table) and often composed of meaningful shapes/concepts (e.g., rows, towers, diagonals, planes, etc.).
-Architect instructions describe these structures at various levels of abstraction, from high-level references to the whole structure ("flower" or "bell tower"), mid-level references to sub-shapes ("row" or "plane"), down to low-level instructions referencing individual blocks.
+Architect instructions describe these structures at various levels of abstraction, from high-level references to the whole structure (“flower" or “bell tower"), mid-level references to sub-shapes (“row" or “plane"), down to low-level instructions referencing individual blocks.
 This observation motivates the need for similarly structured synthetic data as well.
-We first describe a simulator that can generate these structures (Section 5.3.1), and then outline two types of dialogs simulated based on these structures (Sections 5.3.2 and 5.3.3).
+We first introduce a simulator that can generate such ordered and shape-based target structures (Section 5.4.1), and then outline two types of distinct dialogue simulators based on these structures (Sections 5.4.2 and 5.4.3).
     5.4.1 Shape-based structures
     We define six unique elementary shapes: rows, diagonals, T-shapes, L-shapes, U-shapes, and planes.
-These shapes were chosen for their simplicity and occurrence in MDC dialogs and structures.
+These shapes were chosen for their simplicity and occurrence in the MDC.
 (Formal definitions of the shapes are provided in the appendix (Section 10.5)).
-We design a simulator that generates target structures by combining multiple randomly sampled elementary shapes into a single composite structure.
-Sampling is performed with replacement, allowing multiple instances of the same shape type within a structure.
-For each shape instance, properties such as size, color, orientation, and exact 3D location are randomly sampled within specific constraints (e.g., minimum and maximum size bounds).
-Some global constraints are also enforced (similar to those for random structures in Section 5.2, with an added constraint for shapes):
-  1. 1.
-    The structure (i.e. at least one shape) must have at least one block on the ground.
-  2. 2.
-    The entire structure must be "connected", i.e., each block shares at least one face or edge with at least one other block, and each shape instance shares at least one block face or edge with at least one other shape instance.
-The simulator is flexible, allowing parameters such as shape types, number of shapes per structure, etc. to be customized. Figure 14 illustrates two examples of such target structures.
-Figure 14: Shaped-based target structures
-5.3.2 Blocks-based dialogs for shape-based structures
-Target structures
-Each target structure consists of three shape instances, sampled from the full set of the six elementary shapes.
-Dialogs
-The dialog simulation (Section 5.1.2) takes the target structure as input and continues iteratively until the target structure is built.
-Construction proceeds one shape at a time.
-A ’s planning is involved and follows a heuristic that selects one or more blocks to place in each iteration until all shapes are completed.
-Consequently, A ’s instructions reference a single block or multiple ones collectively, with each shape decomposed into multiple instructions.
-Most of A ’s instructions include spatial relations relative to the last block placed by B (e.g., "now, place two red blocks behind the last block").
-Figure 15 illustrates an example BAP item from this dataset.
+We then implemented a flexible target structure simulator that programmatically combines these elementary shapes into complex, composite structures.
+The simulator operates by randomly sampling multiple shape instances, allowing for repetition.
+For each instance, it further randomizes key properties such as size, color, orientation, and 3D location within specified constraints (e.g., minimum and maximum size bounds).
+The simulator is flexible, allowing parameters such as shape types, number of shapes per structure, etc.
+to be customized.
+    5.4.2 Dataset $D_{\mathrm{sb}}$: Blocks-based dialogues for shape-based targets
+    Each target structure consists of three shape instances sampled from the full set of six elementary shapes.
+Dialogues are generated by applying the simulation framework to the predefined shape-based structures until construction is complete.
+The simulation breaks down the construction of complex shapes into a sequence of simpler, block-level instructions.
+The Architect Planning (
+    Step 5.16.
+    1) is more sophisticated here.
+The Architect formulates a plan to build the target one shape instance at a time.
+Within each shape, it uses a heuristic to select one or more nearby blocks to place in each turn, ensuring a contiguous and logical build order.
+Consequently, the Generating the Main Dialogue (
+    Step 5.17.
+    3) produces low-level instructions that decompose each shape into multiple steps, referencing either a single block or a small group of blocks.
+A key linguistic feature of this dataset is that spatial relations are primarily anchored to the last block placed by B (e.g., “now, place two red blocks behind the last block").
+Figure 9 (middle) provides examples.
 The first blue square has been built, and the second (yellow) shape is under construction.
-A is instructing B to continue building the latter by placing three more blocks, using a 1D spatial relation "on top of" relative to the last block.
-(In the previous step, A used a 2D spatial relation, "diagonally in front of and to the right of".) We provide more details about the simulator in the appendix (Section 11).
-Figure 15: A BAP item from the blocks-based dialogs for shape-based structures
-5.3.3 Shape-based dialogs for shape-based structures
-Target structures
-For tractability in dialog simulation and to prevent overly complex utterances, each target structure is limited to two shape instances, sampled from only three elementary shapes: rows, diagonals, and planes.
-Dialogs
-The dialog simulation (Section 5.1.2) takes the target structure as input and proceeds shape by shape, with one complete shape placed per iteration until the target is built.
-A ’s planning employs simple heuristics to determine the order of shapes and select a target block within each shape for defining spatial relations, which also serves as B ’s starting point for building the shape.
-A ’s instructions, therefore, describe entire shapes rather than individual blocks, resulting in higher-level, abstractive instructions.
-Since each target structure consists of two shapes, each placed in one go, a 1D/2D/3D spatial relation is only used to specify the position of the initial block of the second shape relative to the last block placed in the first shape.
-Additionally, A can also use spatial relations that use B as a spatial anchor (Section 5.1.3), for building either shape.
-B’s low-level planning also uses heuristics to determine the order of blocks within a shape, mimicking human Builders.
-Multiple orderings are possible, with some being more optimal or natural than others.
-E.g., 2D shapes like planes can be built either by column or by row.
+A instructs B to continue building the latter by adding blocks relative to the last one placed.
+A uses a 1D spatial relation “on top of" and the last block as the reference block.
+(In the previous step, A used a 2D spatial relation, “diagonally in front of and to the right of".) Further details on this simulator are available in Appendix 10.
+    5.4.3 Dataset $D_{\mathrm{ss}}$: Shape-based dialogues for shape-based targets
+    This simulator generates dialogues at a higher level of abstraction than the block-based simulator described above, but, to ensure the dialogue simulation remains tractable and to avoid overly complex utterances, each target structure here is limited to two shape instances, sampled from only three elementary types: rows, diagonals, and planes.
+In Architect Planning (
+    Step 5.18.
+    1), the Architect plans to construct an entire shape instance within a single turn.
+The plan involves simple heuristics to select the order of the two shape instances and to identify a starting block for the second shape relative to the first.
+This high-level plan leads to more abstract, mid-level instructions during Generating the Main Dialogue (
+    Step 5.19.
+    3).
+Instead of referring to individual blocks, utterances describe whole shapes (e.g., “add a five block long red line…").
+Spatial relations are used in two distinct ways: 1D, 2D, or 3D relations are used to specify the starting point of the second shape relative to the first, while relations that use B as a spatial anchor (Section 5.2.3) are used to define a shape’s orientation (e.g., “…going to the left of you").
+This approach also simulates more complex Generating the Builder Actions (
+    Step 5.20.
+    4).
+Given a high-level instruction to build a shape, the Builder must use its own heuristics to determine the actual sequence of block placements, mimicking how a human might build.
+E.g., 2D shapes like planes can be built either column-by-column or row-by-row.
 Additionally, B may also adopt a zigzag pattern, alternating the starting side for each row or column.
-Figure 16 illustrates a BAP item from this dataset.
-The first shape (a yellow column) has been built, and A is instructing B to build the second shape (a red row) in one go.
-A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row, along with the 3D spatial relation "one to the left, two underneath, and one block in front of" relative to the last block (the topmost yellow one), to indicate the starting position for the row.
-We provide more details about the simulator in the appendix (Section 11).
-Figure 16: A BAP item from the shape-based dialogs for shape-based structures
-6 Training with Synthetic Data
-Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with straightforward training methods, thereby addressing the challenge of insufficient training data.
-We first show that just joint training on the aggregation of the synthetic data and the original BAP (augmented) data already yields a better model (Section 6.1).
-We then go a step further and show that a simple Curriculum Learning method can further boost performance (Section 6.2).
-The synthetic data together with the original BAP (augmented) data now constitutes the BAP v2 training dataset, completing the BAP v2 framework.
-We also illustrate the impact of the synthetic data and training methodologies on training a simple LLM (BERT) and transformer-based model (Section 6.3).
-A note on experimental design choices
-For the GRU-based models in Sections 6.1 and 6.2, we maintain the same architecture and hyperparameters as the baseline model (Section 2.6), along with the same amount of data augmentation in the BAP training data as in our prior work (Section 2.5).
-Fine-tuning these aspects for different training methods is beyond the scope of our objectives here and left for future work to further optimize performance.
-Similar choices made for the BERT-based models are discussed later in Section 6.3.
+Figure 9 (bottom) shows an example.
+The first shape (a yellow column) has been built, and A instructs B to build the second shape (an entire red line) in one go, using both types of spatial relations.
+A uses the spatial relation “going to the left of you" (using B as a spatial anchor) to indicate the direction for the row, along with the 3D spatial relation “one to the left, two underneath, and one block in front of" relative to the last block (the topmost yellow one), to indicate the starting position for the row.
+Further details on this simulator are available in Appendix 10.
+    5.5 A Qualitative Comparison of the Datasets and MDC
+[…]
+    5.6 Related Work: Synthetic Data
+    Synthetic data has proven beneficial for various tasks, including those mentioned in Section 2.8.
+VLN agents, constrained by limited human instruction data and diversity in training environments, often struggle with complex language grounding and spatial language understanding.
+Kamath et al.
+(2023) address this by using synthetic data, achieving SOTA performance on the RxR dataset (Ku et al., 2020), while Wang et al.
+(2023) are similar in spirit and achieve SOTA on the aforementioned CVDN and other benchmarks.
+Similarly, Kang et al.
+(2023) demonstrate strong performance gains in low-data regimes for Visual Dialogue Das et al.
+(2017) by leveraging synthetic data.
+Synthetic data has also been effective for dialogue systems, particularly those focused on text-only dialogues (Kim et al., 2022; Bao et al., 2023; Zhan et al., 2023).
+Zhan et al.
+(2024) is an example of extending this to multi-modal dialogue agents, demonstrating the effectiveness of synthetic visual descriptions in enhancing agents’ grounding capabilities.
+We have also seen synthetic data used to improve VLM spatial reasoning.
+Tang et al.
+(2024) do so by constructing simple images and fine-tuning over simple tasks such as 2D object localization, distance estimation, navigations, etc to improve spatial reasoning capabilities in 3D images.
+    Synthesizing Task-Oriented Embodied Dialogues
+    However, fewer efforts have addressed synthesizing complete task-oriented embodied dialogues (involving both utterances and environment actions).
+Padmakumar et al.
+(2023) were the first to design a framework to do so.
+They extend agenda-based dialogue simulation (Schatzmann and Young, 2009) to a multimodal embodied agent framework, and demonstrate the impact of the synthetic dialogues on the TEACh task (see Section 2.8 for a comparison of TEACh to MDC).
+Our work serves to provide another useful example in this relatively underexplored general domain, further advancing investigation of synthetic data generation for such complex tasks.
+Dan, Han, and Roth (2021) is another closely related work that demonstrates the potential of synthetic data in the Blocks World domain.
+They use simpler synthetic data for the benchmarks of Bisk, Yuret, and Marcu (2016), which require understanding single-shot instructions that transform one world state to another using simulated 3D blocks.
+Blocks are viewed from a fixed bird’s eye perspective, initialized randomly in the initial world state, and uniquely identifiable.
+The varying Builder perspective and lack of easily identifiable referents, along with the need to understand utterances in a dialogue context, make the BAP task a much more challenging problem.
+[^8] As noted in Section 2.1.1, unlike traditional Blocks World, Minecraft allows blocks to float (requiring non-monotonic action sequences where placement is followed by removal), or attach to any side of an existing block.
+Our work on synthetic data thus serves as a demonstration of the potential for richer, embodied task-oriented dialogue tasks that extend beyond simpler Blocks World settings.
+    LLM-generated Synthetic Data
+    While other domains have adopted LLM-generated synthetic data, embodied task-oriented dialogue has generally not followed suit.
+LLMs are able to generate diverse dialogues (Samarinas et al., 2024) or dialogue counterparts to an existing modality (Zhang et al., 2025a), but they struggle to generate the real-world environment and track state-changes within it, even on the simpler un-embodied task-oriented dialogue settings (Li et al., 2024b; Feng et al., 2023b; Kulkarni et al., 2024; Hudeček and Dusek, 2023).
+Generating dialogue from scratch (as Du et al.
+(2025) do) requires complex systems of planning and synthesis that are extremely domain-specific and must be evaluated by humans to verify coherence, flow, and accuracy.
+Our work, therefore, utilizes a programmatic simulation framework to overcome such limitations.
+This method guarantees that all generated data is accurately grounded in the game’s world state via highly nuanced geometric computations (Section 5.2), providing a reliable and systematic way to generate high-fidelity training data for complex, embodied tasks like BAP.
+    6 Revisiting BAP Modeling: Training with Synthetic Data
+    Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with straightforward training methods, thereby addressing the challenge of insufficient training data, and thus validating one of the motivations behind our synthetic data.
+We demonstrate this for both non-LLM- and LLM-based models, highlighting the robustness of our approach.
+The synthetic data together with the original BAP training data now constitutes the BAP v2 training data in the BAP v2 framework.
     Key Terminology
-Hereafter, we refer to the random structure-based dialogs (Section 5.2) as $D_{\mathrm{r}}$, blocks-based dialogs for shape-based structures (Section 5.3.2) as $D_{\mathrm{bs}}$, and shape-based dialogs for shape-based structures (Section 5.3.3) as $D_{\mathrm{ss}}$.
+    Recall that we refer to the synthetic datasets as $D_{\mathrm{r}}$, $D_{\mathrm{sb}}$ and $D_{\mathrm{ss}}$.
 The Minecraft-based original BAP data is denoted as $D_{\mathrm{mc}}$.
-Each dataset includes train, test, and val splits, with the $D_{\mathrm{mc}}$ train set also incorporating augmented data from Section 2.5.
-The baseline model (Section 2.6), trained on $D_{\mathrm{mc}}$ alone, is denoted as $M_{\mathrm{mc}}$.
-6.1 Joint training on aggregated data
-The most straightforward approach to use the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) is to just combine them all together along with the original $D_{\mathrm{mc}}$ data and retrain the baseline model on this aggregated dataset.
+Each dataset includes train, test, and val splits.
+The GRU-based baseline model (Section 2.6), trained on $D_{\mathrm{mc}}$ alone, is denoted as $M_{\mathrm{mc}}$.
+    6.1 The GRU-based models
+    The most straightforward approach to use the synthetic datasets ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) is to just combine them all together along with the original $D_{\mathrm{mc}}$ data and retrain the GRU-based baseline model on this aggregated dataset.
     6.1.1 Experimental Setup
     Our setup closely follows the one used for the baseline model $M_{\mathrm{mc}}$ when it was trained on $D_{\mathrm{mc}}$ alone (Section 2.6).
-We retain the same exact model architecture and hyperparameters as in $M_{\mathrm{mc}}$.
 Training employs teacher forcing with cross-entropy loss, and decoding is performed via greedy decoding, with a maximum sequence length of 20 actions.
 Further details are provided in the appendix (Section 11.2).
-We use the fairer F1 metric to report micro F1 scores on the BAP v2 test set for $D_{\mathrm{mc}}$ (Section 4.4) and the test sets of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$.
-[^6] For fairer F1 on $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $\mathcal{M}$ consists of all EB items, unlike $D_{\mathrm{mc}}$, where only a subset is considered (Section 4.1).
-This is because, by design, all EB items in $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ have multiple correct interpretations.
-Results on the full v2 evaluation benchmark are reported for the baseline and best performing models in this work later in Section 7.1.
+We use the fairer F1 metric to report micro F1 scores on the BAP v2 test set for $D_{\mathrm{mc}}$ (Section 4.3) and the test sets of $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$.
+[^9] For fairer F1 on $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $\mathcal{M}$ consists of all EB items, unlike $D_{\mathrm{mc}}$, where only a subset is considered (Section 4.1).
+This is because, by design, all EB items in $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ have multiple correct interpretations.
     How much of each synthetic dataset to use during train time?
@@
     Dataset Train Val Test
-$D_{\mathrm{bs}}$ 9890 1186 1181
+    $D_{\mathrm{sb}}$ 9890 1186 1181
     $D_{\mathrm{ss}}$ 11868 1000 1000
@@
     6.1.2 Experiments and Results
-Table 4 presents the results.
-We denote the model trained on the aggregated dataset as $M_{\mathrm{agg}}$, and include comparisons to the baseline model ($M_{\mathrm{mc}}$– trained on $D_{\mathrm{mc}}$), and models trained on the three synthetic datasets separately ($M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$– trained on $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ respectively).
-F1 performance is reported (as percentages) on the respective test sets of all four datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$).
-$M_{\mathrm{agg}}$ achieves an F1 score of 30.3% on $D_{\mathrm{mc}}$, up from the baseline model $M_{\mathrm{mc}}$ ’s 27.3%, a notable 3-percentage point improvement.
-Additionally, $M_{\mathrm{agg}}$ is highly performant across the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}$ by large margins on these.
-For a sense of what high performance means on these datasets, we reference the diagonal entries in rows for models $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, representing performance of these models on the same test data distribution as train: 83.9% for $D_{\mathrm{bs}}$, 75.4% for $D_{\mathrm{ss}}$, and 64.4% for $D_{\mathrm{r}}$.
-In light of these numbers, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
+    Table 4 report overall F1 performance on the respective test sets of all four datasets ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$) for all variants of our GRU models.
+We denote the model trained on the aggregated dataset as $M_{\mathrm{agg}}$, and include comparisons to the baseline model ($M_{\mathrm{mc}}$, trained on $D_{\mathrm{mc}}$), and models trained on the three synthetic datasets separately ($M_{\mathrm{sb}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, trained on $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ respectively).
+    $M_{\mathrm{agg}}$ achieves an F1 score of 30.3 on $D_{\mathrm{mc}}$, up from the baseline model $M_{\mathrm{mc}}$ ’s 27.3, a notable 3- point improvement.
+Additionally, $M_{\mathrm{agg}}$ is highly performant across the synthetic datasets ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}$ by large margins on these (83.4 vs.
+35.2 on $D_{\mathrm{sb}}$, 68.8 vs.
+19.1 on $D_{\mathrm{ss}}$, and 63.6 vs.
+12.4 on $D_{\mathrm{r}}$).
+It is also informative to compare both the $M_{\mathrm{mc}}$ and $M_{\mathrm{agg}}$ numbers with the corresponding diagonal entries in the $M_{\mathrm{sb}}$, $M_{\mathrm{ss}}$ and $M_{\mathrm{r}}$ rows, which report the performance of these models when trained only on these datasets: 83.9 for $D_{\mathrm{sb}}$, 75.4 for $D_{\mathrm{ss}}$, and 64.4 for $D_{\mathrm{r}}$ with the baseline $M_{\mathrm{mc}}$ numbers.
+$M_{\mathrm{mc}}$ clearly underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
 In contrast, $M_{\mathrm{agg}}$ is more robust, and retains much of the diagonal-entry performance on the synthetic datasets while also significantly improving on $D_{\mathrm{mc}}$.
-It is also better on the synthetic datasets than on $D_{\mathrm{mc}}$ by large margins (as it should be), unlike $M_{\mathrm{mc}}$.
 Lastly, the models trained solely on synthetic datasets perform poorly on $D_{\mathrm{mc}}$, as does $M_{\mathrm{mc}}$ on synthetic data, indicating that joint training on the aggregated dataset is essential for robust performance across both real and synthetic data.
     Model Train Data Test Data
-$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
+    $D_{\mathrm{sb}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
     $M_{\mathrm{mc}}$ $D_{\mathrm{mc}}$ 35.2 19.1 12.4 27.3
-$M_{\mathrm{bs}}$ $D_{\mathrm{bs}}$ 83.9 12.9 44.9 6.3
+    $M_{\mathrm{sb}}$ $D_{\mathrm{sb}}$ 83.9 12.9 44.9 6.3
     $M_{\mathrm{ss}}$ $D_{\mathrm{ss}}$ 11.3 75.4 7.9 9.1
     $M_{\mathrm{r}}$ $D_{\mathrm{r}}$ 43.8 10.1 64.4 7.7
-$M_{\mathrm{agg}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 83.4 68.8 63.6 30.3
-Table 4: Models trained on datasets separately ($M_{\mathrm{mc}}$, $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$) and together ($M_{\mathrm{agg}}$)
-6.2 Curriculum Learning (CL)
-Another key observation from Table 4 is that performance of the $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, and $M_{\mathrm{mc}}$ models decreases in the order $M_{\mathrm{bs}}$ > $M_{\mathrm{ss}}$ > $M_{\mathrm{r}}$ > $M_{\mathrm{mc}}$, suggesting that the complexity/difficulty of examples increases in the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$, with the most complex $D_{\mathrm{mc}}$ examples naturally appearing at the end.
-This observation raises the possibility of using this ordering to improve over the naive joint training approach (where data is randomly ordered) when training on the aggregated dataset, so as to achieve even better performance on the complex $D_{\mathrm{mc}}$.
-Curriculum Learning (CL) is a natural next step to explore in this direction.
-CL is a training strategy that sequences training data from simpler to more complex examples, similar to human learning.
-Introduced by Bengio et al.
-(2009), CL begins with easier examples, gradually increasing in difficulty.
-This progression allows the model to first learn basic patterns and representations, building a foundation before tackling more challenging cases, which can enhance generalization and improve performance on complex tasks.
-6.2.1 CL Mechanics
-Difficulty Metrics
-A key prerequisite in CL is establishing difficulty metric(s) that quantify how difficult a training example is for a model to learn, allowing the data to be sorted accordingly during training.
-In this work, we explore three straightforward, natural, and intuitive difficulty metrics, along with their combinations (e.g., sorting first by one metric and then sorting (breaking ties) by another).
-The first metric is the dataset $D$ an example belongs to ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$).
-As mentioned before, the datasets can be ordered as $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$ in terms of difficulty.
-Additionally, we consider the lengths of the action (output) and input sequences: the action (output) sequence length $L_{A}$ reflects the number of actions in the gold action sequence, while the input sequence length $L_{I}$ measures the number of tokens in the game history fed to the model.
-These are also intuitive as longer sequences are harder to generate as output and process as input.
-CL Algorithm
-We make use of a simple CL algorithm, specifically using the Curriculum Arrangement algorithm within the CL framework proposed by Xu et al.
-(2020).
-Details of the method are provided in the appendix (Section 12.1).
+    $M_{\mathrm{agg}}$ $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 83.4 68.8 63.6 30.3
+    Table 4: Models trained on datasets separately ($M_{\mathrm{mc}}$, $M_{\mathrm{sb}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$) and together ($M_{\mathrm{agg}}$)
+    6.2 LLMs (Llama)
+    To demonstrate the robustness of our synthetic data and its applicability to contemporary Large Language Models (LLMs), we conduct experiments with Llama (Dubey et al., 2024).
+Specifically, we consider the Nebula model recently proposed in Chaturvedi, Thompson, and Asher (2024) that established a new SOTA on the BAP task and aim to further improve upon it.
+We refer to it as our LLM-based baseline model.
+It is a Llama-3-8B model that takes as context all conversation and action sequences up to the action sequence $A$ to predict $A$ in a text-to-text fashion.
+It is finetuned on the Minecraft data $D_{\mathrm{mc}}$ using QLoRA (Dettmers et al., 2023).
     6.2.1 Experiments and Results
-Overall experimental setup remains similar to the one used for joint training (Section 6.1.1).
-We experiment with each of the three difficulty metrics individually and the four possible combinations of them, making a total of seven metrics – $D$, $L_{A}$, $L_{I}$, $(D,L_{A})$, $(D,L_{I})$, $(L_{I},L_{A})$, and $(D,L_{I},L_{A})$ (ordering within a tuple denotes the sorting order).
-We also experiment with the CL algorithm’s parameter, $N$, trying six values – 5, 10, 15, 20, 25, and 30.
-(This controls the number of buckets/difficulty levels – see appendix (Section 12.1).) This results in a grid of $7\times 6=42$ experiments.
-Table 5 presents the best model, $M_{\mathrm{agg+cl}}$, selected based on $D_{\mathrm{mc}}$ performance.
-For comparison, we also include the baseline $M_{\mathrm{mc}}$, and the aforementioned $M_{\mathrm{agg}}$ model (Table 4).
-On $D_{\mathrm{mc}}$, $M_{\mathrm{agg+cl}}$ achieves an F1 of 33.0%, a notable 2.7-point improvement over $M_{\mathrm{agg}}$, and an even more significant 5.7-point improvement over the baseline $M_{\mathrm{mc}}$.
-It also continues to be highly performant on the synthetic data similar to $M_{\mathrm{agg}}$.
-Thus, Curriculum Learning on the aggregated data provides a meaningful performance boost on the BAP task over naive joint training.
-Hereafter, we refer to $M_{\mathrm{agg+cl}}$ as the best model.
-Model Train Data Test Data
-$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
-$M_{\mathrm{mc}}$ $D_{\mathrm{mc}}$ 35.2 19.1 12.4 27.3
-$M_{\mathrm{agg}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 83.4 68.8 63.6 30.3
-$M_{\mathrm{agg+cl}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 84.0 68.8 60.3 33.0
-Table 5: Curriculum learning ($M_{\mathrm{agg+cl}}$) compared to models $M_{\mathrm{mc}}$ and $M_{\mathrm{agg}}$
-6.3 Transformers and LLMs (BERT)
-As noted in Section 1, modeling is not the primary focus of this work.
-However, solely to test the robustness of our synthetic data and training methodologies and show their applicability to relatively contemporary model architectures, we experiment with a very simple alternative version of the baseline model (Section 2.6) that incorporates a Transformer architecture (Vaswani et al., 2017) and an LLM (BERT; Devlin et al.
-(2019)).
-Specifically, we replace the GRU-based encoder (and GloVe embeddings) with BERT and substitute the GRU decoder with a Transformer, keeping the rest of the architecture unchanged (Figure 4).
-(A projection layer is added atop BERT for dimensionality reduction.) The simplicity also allows us to assess whether/to what extent simpler LLMs like BERT and vanilla Transformers alone can already effectively handle the BAP task.
-Future work can build on the insights we obtain to explore more sophisticated LLMs and architectures while leveraging the BAP v2 framework.
-Model details
-To control the representation dimensionality and complexity of the decoder transformer, we add a feedforward projection layer atop BERT, compressing its output (as otherwise the dimensionality of the decoder transformer would be forced to be the same as the BERT encoder and make it overly complex than need be).
-The BERT encoder and transformer decoder are connected in standard encoder-decoder fashion but via the projection layer.
-BERT’s token embeddings are compressed by the projection layer, and the transformer decoder then attends to these representations at each decoding time-step, and the representation dimensionality of the decoder matches the output size of the projection layer.
-This also matches the embedding dimensionality for action vectors fed into the decoder.
-Positional embeddings are also used for the decoder input.
-The output of the decoder’s last layer conditions the CNN-based action predictor (replacing the GRU hidden state from the baseline model).
-6.3.1 Experiments and Results
-We conduct similar experiments with this model like with the GRU-based model, i.e., training on $D_{\mathrm{mc}}$ only, joint training on the aggregated dataset, and CL.
-The experimental setup and the CL experiments are the same as for the GRU scenario (Sections 6.1.1 and 6.2.2).
-Further details specific to the BERT scenario are provided in the appendix (Section 12.4).
-We tested various sizes of BERT models (Turc et al., 2019), and keeping BERT frozen versus finetuning it.
-We found that frozen $\text{BERT}_{\text{BASE}}$ (110.1M parameters) performed the best.
-In the transformer decoder, five layers yielded optimal results, and a projection layer output size of 256 proved most effective.
-Other hyperparameters are provided in the appendix (Section 12.4).
-For tractability of experimentation, we only experimented with the transformer’s and projection layer’s hyperparameters, keeping other architectural hyperparameters identical to the GRU-based models (e.g.
-those of the CNNs).
-Results are shown in Table 6.
-(Note: We report only the best-performing BERT-based model, which was from the frozen scenario as mentioned earlier.
-However, similar to the results for the frozen case discussed below, synthetic data, joint training, and CL also improved model performance in the finetuning scenario, validating our approach.)
+    We conduct experiments analogous to those performed with the GRU-based models, along with additional experiments on the input textual representation provided to the LLM.
+The results are detailed in Table 6.
+All models involve finetuning Llama-3-8B using QLoRA.
+    LLM-based baseline:
+    First, we reproduce the Nebula LLM-based baseline by fine-tuning Llama-3-8B on $D_{\mathrm{mc}}$ but this time evaluating it on $D_{\mathrm{mc}}$ using the v2 evaluation as well as on the synthetic data.
+As shown in the first row of Table 6, this model ($M_{\mathrm{mc}}^{\prime}$) achieves an F1 score of 47.0 on $D_{\mathrm{mc}}$, significantly better than its GRU-based counterpart $M_{\mathrm{mc}}$ (27.3).
+However, similar to $M_{\mathrm{mc}}$, it underperforms on the easier synthetic data ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), despite being trained on the far more complex $D_{\mathrm{mc}}$, albeit its performance is better than that of $M_{\mathrm{mc}}$.
+Thus, this counterintuitive behavior of the GRU-based model still persists even with the far more advanced LLM-based model.
+    Joint training on aggregated data
+    Next, we perform joint training on the aggregated dataset ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$) to produce model $M_{\mathrm{agg}}^{\prime}$.
+This yields substantial improvements on $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, as expected, but only a marginal gain on $D_{\mathrm{mc}}$ (47.8 compared to 47.0.
+Furthermore, its performance on the synthetic datasets still lags behind its GRU-based counterpart, $M_{\mathrm{agg}}$.
+    Input Representation: NN
+    Predict the action sequence (AS) for the Minecraft excerpt:
+    <Architect> place one in the center of one of the lines near the edge
+    place yellow 5 1 0
+    <Architect> good; repeat to make a V
+    place yellow 3 1 -1
+    place yellow 3 1 1
+    <Architect> still diagonal from the previous squares
+    <Builder> extended?
+    <Architect> not really; the first step is orange, but it is in line with the yellow blocks
+    pick 3 2 1
+    pick 3 1 1
+    pick 2 2 2
+    <Architect> now one orange on top of teh second orange
+    place orange 1 2 2
+    <Architect> now delete the second orange
+    pick 1 1 2
+    <Architect> good repeat to make a third step
+    ### AS:
+    Input Representation: $N+\textsc{Pos}_{\textbf{B}}$ (appends builder’s position and orientation to the input above)
+    Builder’s current position is 0.2 1.0 -0.4 and orientation (yaw) is -6.5 degrees
+    ### AS:
+    Input Representation: $N+\textsc{Pos}_{\textbf{B}}+S$ (appends current built structure to the input above)
+    Current built structure is:
+    yellow 5 1 0
+    yellow 4 1 1
+    yellow 4 1 -1
+    yellow 3 1 2
+    yellow 3 1 -2
+    orange 2 1 2
+    orange 1 2 2
+    ### AS:
+    Output (Common to all inputs)
+    place orange 1 3 2
+    place orange 0 3 2
+    pick 1 3 2
+    Table 5: The three different input representations ($N$, $N+\textsc{Pos}_{\textbf{B}}$, $N+\textsc{Pos}_{\textbf{B}}+S$) and the (common) ground truth output.
+Ellipses (…) denote truncation for readability.
+$N$ is from Chaturvedi, Thompson, and Asher (2024), and $N+\textsc{Pos}_{\textbf{B}}$, $N+\textsc{Pos}_{\textbf{B}}+S$ are ours.
+    Richer input representations
+    We hypothesize that the lower-than-expected results of $M_{\mathrm{agg}}^{\prime}$ stems from the input representation, which lacks a critical component of the game context available to the GRU models – the Builder’s position and orientation, $\textsc{Pos}_{\textbf{B}}$, which is essential for correctly interpreting spatial relations (Section 2.4).
+The GRU models incorporate $\textsc{Pos}_{\textbf{B}}$ (Jayannavar, Narayan-Chen, and Hockenmaier, 2020), making the comparison with the LLM models uneven.
+We therefore enrich the input text by appending information about $\textsc{Pos}_{\textbf{B}}$, as illustrated in Table 5 (representation $N+\textsc{Pos}_{\textbf{B}}$).
+We denote the original input representation used in Nebula as $N$ and show that at the very top of the table.
+(The ground truth output representation is shown at the very bottom.)
+    Training on the aggregated data with this richer input yields model $M_{\mathrm{agg+posb}}^{\prime}$, which shows significant improvement on $D_{\mathrm{mc}}$ and even more on the synthetic data, in comparison to $M_{\mathrm{agg}}^{\prime}$.
+It achieves F1 scores of 93.4 on $D_{\mathrm{sb}}$, 84.9 on $D_{\mathrm{ss}}$, and 82.3 on $D_{\mathrm{r}}$, while also boosting the $D_{\mathrm{mc}}$ score to 52.5, and now also overtakes the performance of the GRU models on all datasets, validating out hypothesis that suitable input representations are essential for this task.
+    However, to verify that this gain is not from the representation alone, we trained a model ($M_{\mathrm{mc+posb}}^{\prime}$) on just $D_{\mathrm{mc}}$ but with the new representation.
+This results in an F1 score of 47.6 on $D_{\mathrm{mc}}$, a marginal improvement over the 47.0 from its counterpart $M_{\mathrm{mc}}^{\prime}$ that used the original Nebula representation.
+This confirms that both the enriched input representation and the synthetic data are necessary to achieve significant performance gains on both synthetic and real Minecraft data; each on their own are not enough.
+The explicit spatial context in the input allows the model to better leverage the patterns in the synthetic data and transfer that knowledge to the $D_{\mathrm{mc}}$ domain.
+    We also note that while our textual representation of $\textsc{Pos}_{\textbf{B}}$ is fairly straightforward, its effect on performance is very substantial.
+This input only explicitly provides the Builder’s frame of reference, but for the model to correctly ground spatial instructions, it must "infer" a more complex, situated understanding of the 3D world state from this frame.
+Our results demonstrate that the LLM seems to become, to a significant extent, capable of this complex inference, just through the textual signal in the data and standard training (teacher forcing and cross entropy loss).
+Crucially, however, this ability due to $\textsc{Pos}_{\textbf{B}}$ only emerges when the model is trained on enough signal from the rich spatial examples in our synthetic data in conjunction with the original $D_{\mathrm{mc}}$.
+    A further boost
+    Finally, we investigate whether an explicit representation of the current world state/built structure $S$ (like in the GRU-based models) helps further.
+While the action history contained in the input text already "implicitly" defines $S$, we append an explicit version of it to the input (representation $N+\textsc{Pos}_{\textbf{B}}+S$ in Table 5).
+Joint training on the aggregated data with this input produces model $M_{\mathrm{agg+posb+s}}^{\prime}$ which further boosts performance on $D_{\mathrm{mc}}$ by a bit to 53.0 (up from the 52.5 of $M_{\mathrm{agg+posb}}^{\prime}$), while performance on $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ continues to remain strong and comparable to $M_{\mathrm{agg+posb}}^{\prime}$.
+    This result suggests that for the longer and more complex dialogues in $D_{\mathrm{mc}}$, where tracking and inferring the world state/built structure $S$ from just the textual action history is more challenging, an explicit representation of $S$ provides a beneficial signal.
+On the shorter and simpler dialogues in the synthetic datasets, however, this inference is not that challenging, and therefore explicit state information may not be as necessary.
+This likely explains why adding in this representation only improves performance on $D_{\mathrm{mc}}$.
+Moreover, given the already strong performance on the synthetic data after including $\textsc{Pos}_{\textbf{B}}$, further gains from additional inputs such as $S$ may anyway be subject to diminishing returns on the synthetic data and the transfer of knowledge to the $D_{\mathrm{mc}}$ domain (unlike for $M_{\mathrm{agg+posb}}^{\prime}$ as discussed above).
+    We designate $M_{\mathrm{agg+posb+s}}^{\prime}$ as our best model.
+As it makes use of richer game context and synthetic data (with each one enabling the other), we name it Llama-CRAFTS (Context Rich And Fine-Tuned On Synthetic Data).
+This model establishes a new state of the art (SOTA) on the BAP task ($D_{\mathrm{mc}}$), achieving an F1 score of 53.0.
+This represents a 6 point improvement over the 47.0 F1 score of the previous SOTA, the Nebula (the LLM baseline).
+    6.3 Broader Takeaways
     Robustness of our approach
-Overall trends align with those observed for the GRU-based models in Table 5.
-The synthetic data, joint training on the aggregated dataset, and CL, all also improve the BERT-based model’s performance on both synthetic and real data, highlighting the robustness of our approach.
-$M_{\mathrm{agg}}^{\prime}$ achieves an F1 score of 29.4% on $D_{\mathrm{mc}}$, up from $M_{\mathrm{mc}}^{\prime}$ ’s 23.0%, a notable 6.4-percentage point improvement.
-It is also highly performant across the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}^{\prime}$ by large margins on these.
-On $D_{\mathrm{mc}}$, CL again yields the best-performing model, as with the GRU-based models.
-$M_{\mathrm{agg+cl}}^{\prime}$ achieves an F1 of 32.3%, a notable 2.9-point improvement over $M_{\mathrm{agg}}^{\prime}$, and an even more significant 9.3-point improvement over $M_{\mathrm{mc}}^{\prime}$.
-It also achieves significant gains in performance over $M_{\mathrm{agg}}^{\prime}$ on the synthetic data as well.
-Comparison to GRU-based models
-The lower performance of the BERT-based models than their respective GRU-based counterparts’ is more pronounced for performance on the synthetic data than the Minecraft $D_{\mathrm{mc}}$ data.
-$M_{\mathrm{mc}}^{\prime}$ achieves an F1 of 23.0% on $D_{\mathrm{mc}}$, 4.3 points lower than $M_{\mathrm{mc}}$’s 27.3%.
-This gap between the GRU-based and BERT-based counterparts reduces significantly as we move on to look at $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg+cl}}^{\prime}$.
-It goes down to 0.9 and 0.7 points respectively.
-$M_{\mathrm{agg+cl}}^{\prime}$ is therefore not too far below $M_{\mathrm{agg+cl}}$.
-The gap in synthetic data performance between $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg}}$ is fairly significant.
-But CL reduces this quite a bit – $M_{\mathrm{agg+cl}}^{\prime}$ is very close to $M_{\mathrm{agg+cl}}$ on $D_{\mathrm{ss}}$, although still behind on the other two.
-Two key factors contribute to the relatively lower performance of the BERT-based models compared to their GRU counterparts:
-  • As mentioned earlier, the sizes of all four datasets were optimized for the GRU-based models.
-These might need separate tuning for the BERT-based models (and will probably need to be scaled up also as transformers are more complex and data-hungry than GRUs).
-  • For tractability of experimentation, we only experimented with the transformer’s and projection layer’s hyperparameters, keeping other architectural hyperparameters identical to the GRU-based models (e.g.
-those of the CNNs).
-These choices were intentional, as our goal here (as stated in the beginning of this section) was primarily to demonstrate the robustness of our synthetic data and training approach, and not to optimize dataset sizes or architectural hyperparameters for maximizing performance.
-We leave such fine-tuning for future work.
-Model Train Data Test Data
-$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
-$M_{\mathrm{mc}}^{\prime}$ $D_{\mathrm{mc}}$ 29.1 17.8 12.0 23.0
-$M_{\mathrm{agg}}^{\prime}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 70.2 55.0 36.8 29.4
-$M_{\mathrm{agg+cl}}^{\prime}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 75.9 67.4 45.3 32.3
-Table 6: BERT- and transformer-based models
+    By providing the LLM-based models with an input representation that is more comparable to the one used by the GRU-based models, we show that our synthetic data significantly improves performance of both non-LLM- and more modern LLM-based models on both the synthetic and original Minecraft data.
+This highlights the robustness of our data and simple training methodology across different model architectures.
+    Motivating future work
+    With the fairer v2 evaluation now possible, we can more accurately gauge current performance on the BAP task and the remaining gap for improvement.
+While Llama-CRAFTS’s 53.0 F1 score marks improved performance, it also underscores that even with SOTA LLMs, significant room for improvement remains, thus reinforcing the task’s challenging nature and its importance for future research.
+    Importance of evaluation on the synthetic data
+    Going forward, evaluation on the synthetic data will also be important for making more meaningful progress on the BAP task.
+First, the more structured and simpler nature of the synthetic data allows us to assess more basic model competencies than is possible on the more complex $D_{\mathrm{mc}}$.
+Second, as noted above, both GRU and LLM baselines $M_{\mathrm{mc}}$ and $M_{\mathrm{mc}}^{\prime}$ underperform significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$; they also perform lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
+This counterintuitive behavior shows that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust.
+Evaluating on the synthetic data will help build models that do not show such behavior (like we have shown by using the data effectively for training), and complements evaluation on $D_{\mathrm{mc}}$.
+In fact, future models should aim to surpass previous performance on $D_{\mathrm{mc}}$ while also achieving strong results on synthetic data, ideally improving on both.
+    Model Input Train Data Test Data
+    $D_{\mathrm{sb}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
+    $M_{\mathrm{mc}}^{\prime}$ (Nebula) NN $D_{\mathrm{mc}}$ 49.5 31.9 33.3 47.0
+    $M_{\mathrm{agg}}^{\prime}$ NN $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 75.5 54.9 43.8 47.8
+    $M_{\mathrm{mc+posb}}^{\prime}$ $N+\textsc{Pos}_{\textbf{B}}$ $D_{\mathrm{mc}}$ 49.7 32.2 32.0 47.6
+    $M_{\mathrm{agg+posb}}^{\prime}$ $N+\textsc{Pos}_{\textbf{B}}$ $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 93.4 84.9 82.3 52.5
+    $M_{\mathrm{agg+posb+s}}^{\prime}$ (Llama-CRAFTS) $N+\textsc{Pos}_{\textbf{B}}+S$ $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 92.3 83.7 82.2 53.0
+    Table 6: Llama performance with different datasets and input representations.
+    6.4 Related Work: Role of specialized datasets and smaller finetuned models
+    These results highlight a key contemporary question in applied NLP: the relative merits of fine-tuning smaller, specialized models versus leveraging large, general-purpose models via ICL.
+The performance parity between Nebula and GPT-4 on this task (Section 2.7) suggests that a smaller, fine-tuned specialist model can be as effective as a much larger generalist one.
+This finding coupled with the fact that we are able to improve the former by a substantial 12.8% via leveraging synthetic data, motivates a deeper exploration of the fine-tuning paradigm and the role of specialized datasets, especially concerning the most effective pathways to further improve performance.
+    While generalist models offer impressive few-shot capabilities, avenues for their improvement often center on prompt engineering, increased test-time computation, or scaling to even larger (and often proprietary) models, which may not always be practical.
+In contrast, the fine-tuning approach offers more vectors for enhancement, e.g., synthetic data (like we show), better training regimes, etc.
+in addition to the ones for ICL.
+Furthermore, smaller specialized and open models present practical advantages in terms of computational efficiency and accessibility for both academic research and deployment in resource-constrained settings.
+For instance, in real-world counterparts of BAP such as instruction-following robots, stringent low-latency requirements often render the extensive test-time computation demanded by larger models infeasible.
+This pragmatic consideration is supported by emerging theoretical work suggesting that fine-tuning is not merely a matter of convenience.
+Studies indicate that fine-tuned models can achieve generalization capabilities comparable to ICL (Mosbach et al., 2023; Zhao et al., 2025; Ponce and Etchegoyhen, 2025) and may learn representations that are more efficiently embedded in lower-dimensional manifolds, potentially better capturing real-world data/features (Janapati and Ji, 2025).
+    Finally, the development of specialized datasets need not be an endeavor exclusive to fine-tuning.
+The synthetic data we have generated, rich in specific linguistic and spatial phenomena, could also potentially enhance ICL performance.
+For instance, it could be used to select more effective few-shot exemplars, drawing inspiration from works like Kulkarni et al.
+(2024), or be integrated into curriculum learning frameworks for ICL, as explored by Liu et al.
+(2024), given that the synthetic data is simpler than the original BAP data.
+This suggests that creating high-quality, task-specific data is a valuable pursuit that can advance multiple modeling paradigms.
     7 Model Analysis
-We now analyze the performance and behavior of the best model $M_{\mathrm{agg+cl}}$ (Section 6.2.2) relative to the baseline $M_{\mathrm{mc}}$ (Section 2.6).
+    We now analyze the performance of the best LLM-based model Llama-CRAFTS ($M_{\mathrm{agg+posb+s}}^{\prime}$) relative to the LLM-based baseline Nebula ($M_{\mathrm{mc}}^{\prime}$).
 In Section 7.1, we do so using the BAP v2 evaluation benchmark introduced in Section 4.3.
@@
     Dataset Model Type Color Location Shape Overall
-EB Baseline ($M_{\mathrm{mc}}$) 76.4 75.2 53.5 56.0 53.2
-Best ($M_{\mathrm{agg+cl}}$) 81.4 80.0 61.8 65.5 61.6
-NEB Baseline ($M_{\mathrm{mc}}$) 63.8 60.8 24.6 33.8 24.2
-Best ($M_{\mathrm{agg+cl}}$) 66.0 63.2 30.2 39.4 29.6
-Overall Baseline ($M_{\mathrm{mc}}$) 65.2 62.4 27.7 36.3 27.3
-Best ($M_{\mathrm{agg+cl}}$) 67.7 65.0 33.6 42.2 33.0
-Table 7: Micro F1 scores for the baseline and best models under the v2 benchmark
-Dataset Model Type Color Location Shape Overall
-EB Baseline ($M_{\mathrm{mc}}$) 80.9 79.8 61.5 64.9 61.2
-Best ($M_{\mathrm{agg+cl}}$) 83.1 81.7 69.0 72.6 68.6
-NEB Baseline ($M_{\mathrm{mc}}$) 73.1 69.9 25.7 39.0 25.2
-Best ($M_{\mathrm{agg+cl}}$) 73.7 71.2 32.1 46.5 31.5
-Overall Baseline ($M_{\mathrm{mc}}$) 73.8 70.7 28.5 41.1 28.1
-Best ($M_{\mathrm{agg+cl}}$) 74.5 72.0 35.0 48.6 34.5
-Table 8: Macro F1 scores for the baseline and best models under the v2 benchmark
-Tables 7 and 8 present all metrics obtained using the v2 benchmark, augmenting Tables 1 and 2 (Section 4.4) with results for the best model.
-Overall, the best model outperforms the baseline across all metrics.
-Key takeaways are discussed below.
-We begin with the micro scores in Table 7.
-On the overall dataset, the best model outperforms the baseline on all three F1 metrics: Type, Color, and Location.
-As noted in Section 4.4, Location F1 is the primary bottleneck to achieving high performance in this task.
-The best model achieves an overall F1 of 33%, a notable 5.7 percentage-point improvement over the baseline’s 27.3%.
-This increase in overall F1 is primarily driven by the improvement in Location F1, which similarly rises from 27.7% to 33.6%.
-These results validate the motivation and impact of our synthetic data.
-The best model outperforms the baseline in both EB and NEB scenarios.
-However, as discussed in Section 4.4 regarding the importance of performance on EB, there remains room for improvement: the best model reaches only 61.6%, in spite of the increase from the baseline’s 53.2%.
-This insight can inform further error analysis and targeted improvements in future models for this task.
-Turning to the macro scores in Table 8, we observe an increase across all metrics compared to the micro scores, but the same trends persist.
-The best model achieves an overall F1 of 34.5% on the overall test set, a notable 6.4 percentage-point improvement over the baseline’s 28.1%.
-For Shape F1 on the overall test set, the best model scores 42.2% (micro) and 48.6% (macro), up from the baseline’s 36.3% and 41.1%, marking significant gains of 5.9 and 7.5 percentage points, respectively.
+    EB Nebula 92.6 92.0 74.7 80.9 74.2
+    Llama-CRAFTS 94.7 94.7 80.3 85.8 80.3
+    NEB Nebula 78.8 77.6 44.3 57.1 43.5
+    Llama-CRAFTS 80.8 79.5 50.4 59.4 49.4
+    Overall Nebula 80.4 79.2 47.7 59.8 47.0
+    Llama-CRAFTS 82.4 81.2 53.8 62.4 53.0
+    Table 7: F1 scores for the LLM-based baseline Nebula and Llama-CRAFTS
+    As can be seen from Table 7, Llama-CRAFTS outperforms Nebula across all metrics on the v2 benchmark, reaching an overall F1 of 53.0, a notable 6.0 points improvement over Nebula ’s 47.0.
+This increase is primarily driven by an improvement in Location F1, which similarly rises from 47.7 to 53.8.
+However, the BAP task remains far from being solved.
+Although the empty board (EB) can be seen as the most basic variant of the BAP task (Section 4.3), the best model’s 80.3 (albeit a substantial increase from the baseline’s 74.2) is a clear indication that there remains much room for improvement.
     7.2 Qualitative evaluation
-We present examples of output from both the baseline and best models, demonstrating the substantial improvement of the latter.
-For each example, we also discuss the overall F1 metric and relevant auxiliary metrics.
-A complete table of scores for all metrics across examples is provided in the appendix (Table 10).
-Figure 17: Example 1
-Example 1
-As shown in Figure 17, copying substructures is a common strategy in the MDC.
-Here, A instructs B to build an arrowhead, asking B to "mirror" the completed half on the opposite side.
-The baseline model only captures the "on the top" aspect of this instruction, while the best model fully understands and completes the arrowhead.
-This example also involves placing a floating block.
-The best model accomplishes this too, by placing and then removing a supporting block (not shown in the figure).
-It achieves an overall F1 score of 1.0, compared to the baseline model’s 0.
-However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need for a single purple block to be placed (in terms of net actions).
-Figure 18: Example 2
-Example 2
-As shown in Figure 18, A gives a complex instruction requiring blocks to meet several highly specific criteria, including a criterion involving the builder’s position and orientation.
-The baseline model only grasps the "two other orange blocks" aspect but poorly interprets their placement.
-In contrast, the best model fully understands the instruction.
-Similar to the previous example, this example also involves placing floating blocks, which the best model accomplishes successfully.
-It achieves an overall F1 score of 1.0, compared to the baseline model’s 0.
-However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need for two orange blocks to be placed (in terms of net actions).
-Figure 19: Example 3
-Example 3
-As shown in Figure 19, this example mirrors the previous one but focuses on block removals instead of placements.
-The initial structure, $S_{H}^{t-1}$, consists of three steps extending from a base square.
-A instructs B to remove a very specific block, using terms like "center" and "second highest step" to describe its location.
-The baseline model incorrectly selects the third highest step instead of the second, while the best model interprets the instruction accurately.
-The latter achieves an overall F1 score of 1.0, compared to the former’s 0.
-However, both models achieve perfect type and color F1 scores, as the baseline model correctly identifies the need to remove one green block.
-8 Discussion
-We now discuss four topics in detail: an additional key challenge in BAP evaluation and its implications for future work (Section 8.1), the importance of evaluation on the synthetic data (Section 8.2), relevant concurrent work (Section 8.3), and broader implications of our work that go beyond BAP (Section 8.4).
-8.1 An additional key challenge for BAP evaluation
-Through manual inspection of random samples during the qualitative analysis of model outputs (Section 7.2), we identified an additional challenge for fair evaluation in the BAP task, separate from the issue addressed by the fairer F1 metric in Section 4.1.
-Specifically, and rather unsurprisingly, the F1 metric lacks the sensitivity to distinguish nuanced differences in error quality.
-It is possible for two models to make different types of errors, with one type being qualitatively "better" than the other – yet the F1 metric may not capture this distinction effectively.
-This limitation can hinder fair evaluation and reduce our ability to distinguish between models based on their F1 scores.
-Below, we illustrate this challenge through qualitative examples.
-Across all examples, the best model $M_{agg+cl}$ consistently performs qualitatively better than the baseline $M_{mc}$ and is arguably closer to the human reference.
-However, the F1 metric lacks the sensitivity to reflect this substantial quality gap.
-In some cases, the auxiliary metrics proposed in Section 4.2 offer some help.
-A full table of scores for all metrics across examples is provided in the appendix (Table 11).
-It’s worth noting that such examples were frequent enough in our random inspection and the ones discussed here are fairly representative of the problem, not cherry-picked to favor the best model.
-This analysis underscores the inherent challenges in automated evaluation for the BAP task, highlighting the need for further research in this area.
-The examples also illustrate the difficulty of creating a clear, unambiguous rubric for grading model output on an absolute scale, making human evaluation equally complex.
-(In fact, we initially considered human evaluation but ultimately dismissed it due to these challenges.)
-Figure 20: Example A
+    Scenario 1: One model is correct, the other makes mistakes
     Example A
-As shown in Figure 20, A instructs B to place a yellow block meeting multiple highly specific criteria, including one involving the builder’s position and orientation.
-The best model largely understands these criteria and places the correct block, though it also adds an extraneous one on each side.
-In contrast, the baseline model performs poorly, producing seemingly random placements.
-The F1 metric rates the baseline model as 0 and the best model as 0.5, indicating an improvement.
-However, this score does not fully capture the qualitative gap between the two models, especially in spatial reasoning.
-While both models produce two false positives, the errors made by the best model demonstrate a clearer understanding of the task, an aspect that the F1 metric fails to reflect adequately.
-Figure 21: Example B
     Example B
-As shown in Figure 21, A and B are building a 2D giraffe, and A instructs B to place three orange blocks diagonally to form the neck, opposite the tail.
-The baseline model places three orange blocks, but arranges them horizontally and closer to the tail.
-While the best model recognizes the need for a diagonal structure, it places it in the wrong location and orientation.
-Both models receive an F1 score of 0, with identical scores of 1, 1, and 0 for type, color, and location F1, respectively.
-However, shape F1 distinguishes the models more effectively: the baseline achieves a shape F1 of 0.33, while the best model achieves a perfect 1.
-Figure 22: Example C
     Example C
-As shown in Figure 22, A instructs the addition of two rows of orange blocks directly below each hanging row on either side of the arch.
-The baseline model places two orange blocks on top, while the best model attaches blocks to each hanging row—still incorrect but a noticeably closer interpretation of A’s intent.
-Here, too, the F1 metric scores both models as 0. Unlike the previous example, however, none of the auxiliary metrics distinguish between the models, as both receive identical scores.
-Figure 23: Example D
+    Figure 11: The overall F1 score is sufficient to distinguish between correct and incorrect model outputs (Scenario 1)
+    Scenario 2: The auxiliary metrics provide further insight
     Example D
-As shown in Figure 23, A initially directs B to place a red block on the left, but then corrects this to the analogous position on the right, requiring removal of the original block and placement of a new one.
-The baseline model correctly removes the original red block but fails to place a new one, instead removing the top green block randomly.
-The best model performs only the correct removal and places a new red block to the right, though not in the exact position specified—nonetheless, it is clearly a closer interpretation of A’s intent.
-In this case, the F1 metric assigns both models a non-zero score of 0.5, as both have 1 true positive, 1 false positive, and 1 false negative.
-However, the type and color F1 metrics penalize the baseline model, scoring it 0.5 in each, while the best model achieves a perfect score of 1.
-8.2 Evaluation on synthetic data
-Going forward, evaluation on the synthetic data will also be important for making more meaningful progress on the BAP task.
-First, the more structured and simpler nature of the synthetic data allows us to assess basic model competencies.
-Second, as noted in Section 6.1.2, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
-Evaluating on the synthetic data will help build models that do not show such counterintuitive behavior (Section 8.3 provides further evidence supporting this argument), and complements evaluation on $D_{\mathrm{mc}}$.
-In fact, future models should aim to surpass previous performance on $D_{\mathrm{mc}}$ while also achieving strong results on synthetic data, ideally improving on both.
-8.3 Concurrent work
-Concurrent to our work are the recently published (at EMNLP, Nov 12-16, 2024) works of Chaturvedi, Thompson, and Asher (2024) and Kranti, Hakimov, and Schlangen (2024), both exploring SOTA LLMs for the BAP task.
-The former finetunes the Llama-3-8B model (Dubey et al., 2024) on the original BAP data, resulting in the Nebula model, which predicts the next action sequence in a text-to-text fashion using all preceding conversation and action sequences as context.
-The latter employs in-context learning with GPT-4.
-Nebula achieves an F1 score of 39.2% and GPT-4 scores 39.0% (under the now legacy BAP evaluation).
-Nebula is marginally better, although both are very close and mark improved performance on the task.
-As mentioned in Section 1, modeling is not the primary focus of our work.
-Instead, we focus on the BAP task as a whole, addressing key challenges in data and evaluation to propose BAP v2, an upgraded version of the task itself.
-We also illustrate the impact of this data on training a simple LLM and transformer-based model.
-This foundation enables future work to make more efficient and meaningful progress on the task, and sets the stage for more sophisticated models/LLMs.
-These works are therefore complementary to ours.
-A more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in the appendix (Section 15), given some other relevant aspects of their work.
-Here, we evaluate Nebula using the v2 benchmark to further demonstrate the impact and robustness of our evaluation while providing a fairer assessment of its performance.
-We observe a 27.8% gain in performance (compared to the legacy evaluation), similar to the gain for our baseline $M_{\mathrm{mc}}$.
-We also discuss how the v2 evaluation now enables more accurate benchmarking of SOTA LLMs on the BAP task and highlights the significant room for improvement, further motivating the task itself.
-Additionally, it helps us confirm that spatial reasoning remains the key bottleneck for high performance, motivating future efforts leveraging our data and methods, which have shown effectiveness in addressing this challenge.
-We also evaluate Nebula on our synthetic data and find that, like the baseline $M_{\mathrm{mc}}$ (and the BERT-based $M_{\mathrm{mc}}^{\prime}$), it underperforms significantly despite being trained on the far more complex $D_{\mathrm{mc}}$.
-This suggests that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust, reinforcing our argument of Section 8.2 about the importance of the synthetic data evaluation.
-Future work should explore finetuning models such as Llama-3-8B (Nebula) and newer LLMs using our synthetic data and training methodologies like CL.
-BAP v2 evaluation
-Similar to the 29.4% improvement we observed for our baseline model $M_{\mathrm{mc}}$ under the v2 evaluation on $D_{\mathrm{mc}}$, Nebula’s performance also increases from 39.2% to 50.1%, reflecting a 27.8% gain.
-This further underscores the impact and robustness of our v2 evaluation, offering a fairer reflection of Nebula’s performance.
-With this fairer evaluation now possible, we can more accurately gauge current performance on the BAP task and the remaining gap for improvement.
-While the 50.1% score marks improved performance, it also underscores that even with SOTA LLMs, significant room for improvement remains, thus reinforcing the task’s challenging nature and its importance for future research directions.
-Detailed tables for the full v2 benchmark are provided in the appendix (Tables 13 and 14).
-The observed trends generally align with those seen in our models such as the model’s relative strengths and weaknesses.
-They also shed some light on key areas for future improvement.
-For example, spatial reasoning (Location F1) is the key bottleneck for high performance on this task as noted in Section 4.4, and this continues to remain the case.
-Therefore, as demonstrated with $M_{\mathrm{mc}}$, our synthetic data and training methodologies can potentially help address this challenge, motivating future work.
-Evaluation on the synthetic data
-Table 9 shows Nebula’s performance (micro F1) on both the synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and the original $D_{\mathrm{mc}}$ test sets.
-Evaluating Nebula on the synthetic data reveals behavior consistent with our baseline model $M_{\mathrm{mc}}$ (and the BERT-based $M_{\mathrm{mc}}^{\prime}$), as reported in Section 6.1.2 and further discussed in Section 8.2 – it underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$.
-This shows that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust, and reinforces our argument from Section 8.2 about the importance of the synthetic data evaluation: it will help build models that don’t show such counter-intuitive behavior.
-Model Test Data
-$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
-Nebula 51.5 33.1 30.9 50.1
-Table 9: Nebula’s performance across datasets (and using the v2 evaluation on $D_{\mathrm{mc}}$)
-8.4 Broader implications of our work
-While our focus in this work is on the Builder Action Prediction (BAP) subtask within the Minecraft Collaborative Building Task (MCBT) setting, the contributions have broader applications and takeaways as well.
-Below, we discuss their relevance for other subtasks within MCBT, for research building directly on the MCBT, and for similar domains and tasks in the field.
-This highlights the potential impact of our work and directions for future research.
+    Example E
+    Figure 12: Auxiliary metrics provide further insight than the overall F1 score (Scenario 2)
+    Scenario 3: Mistakes that warrant a more semantic evaluation
+    Example F
+    Example G
+    Figure 13: Examples that warrant a more semantic evaluation (Scenario 3)
+    We now present three different scenarios, i.e.
+types of examples of output from Llama-CRAFTS and Nebula that, on the one hand, illustrate different ways in which Llama-CRAFTS arguably displays a better spatial understanding, and, on the other hand, allow us to discuss how this is reflected (or not) in our various metrics.
+    Scenario 1: The overall F1 score is sufficient when one model is correct and the other makes mistakes
+    Figure 11 shows three examples that Llama-CRAFTS gets completely right (receiving perfect scores across all metrics), and Nebula does not.
+In Example A, A instructs B to remove the four corners of the orange ring.
+Nebula removes many extraneous blocks but only three corners (achieving an overall F1 of 0.38).
+In Example B, A asks B to add another red block one tile below and to the left of the last block B had placed.
+Nebula only correctly identifies that a single red block should be placed below the last one, but goes too far down (2 tiles), and misses the second spatial relation (left) altogether, resulting in a perfect type and color F1 scores of 1.0.
+and an overall F1 of 0.0.
+In Example C, A asks B to “copy" the red ring in blue.
+Instructing Builders to copy substructures is a natural and efficient strategy that the human Architects in the MDC employ frequently.
+Nebula achieves perfect type and color scores, but misplaces one of five blocks, resulting in location and shape scores of 0.8.
+In these clear cases, the strict F1 metric is obviously sufficient to capture the qualitative difference between the outputs.
+    Scenario 2: Both models make mistakes, but the auxiliary metrics provide further insight
+    Figure 12 shows examples where both models make mistakes (although Llama-CRAFTS’s mistakes are arguably less egregious), and illustrates how our metrics capture different aspects of this task.
+In Example D, B is asked to make “a red 3 x 3 square" that is “vertical" (and presumably includes the four red blocks already placed).
+Both models add eight red blocks to create squares that leave out the bottom three red blocks and only include the red block above.
+Since Llama-CRAFTS’s square is actually vertical, it achieves a score of 0.77 on all metrics, whereas Nebula built a horizontal square, and hence achieves overall, shape and location scores of 0.31 (and type and color scores of 0.77).
+That is, although the overall score correctly identifies Llama-CRAFTS as the better model, it is the auxiliary metrics that provide additional insight and assign credit to Nebula.
+In Example E, B has just finished the blue-purple-orange (from right to left) line, and A references earlier instructions when saying that “the other one: [should be like the one B already made but with] green red yellow [instead of blue purple orange] …".
+Based on the earlier dialogue (also not shown in the figure), this second line should include the green block already on the board.
+[^10] This example illustrates a shortcoming of our figures: although BAP models have access to the entire preceding context, the reader does not.
+Here, both models receive a 0.0 overall F1 score, Nebula because it produces no net actions (receiving 0.0 scores on all metrics), and Llama-CRAFTS because none of its blocks appear in the intended places.
+However, because Llama-CRAFTS actually places a diagonal line of the correct color (albeit on the other side of the previous line and with the pattern starting from the left instead of the right), it receives type, color, and shape scores of 0.88.
+In this case, the auxiliary metrics are necessary to identify that Llama-CRAFTS’s output is actually quite similar to the desired structure.
+    Scenario 3: Shortcomings of our metrics
+    Figure 13 illustrates shortcomings of our automatic evaluation procedures, and suggests the need for a more complex, semantic evaluation.
+In Example F, B is asked to create a (diagonal) line of four blocks inside the ring-shaped structure, “from top left to bottom right".
+Nebula places three disconnected blocks inside the structure (achieving overall and location scores of 0.0, type and color scores of 0.86, and a shape score of 0.57), while Llama-CRAFTS places a (horizontal) line of four blocks across the inside top edge, as well as an additional block in the (outside) bottom right corner (achieving overall, location and shape scores of 0.22, and type and color scores of 0.89).
+The fact that Nebula’s shape score is higher than Llama-CRAFTS’s points to a limitation of our shape metric (two of Nebula’s blocks can be mapped to the intended diagonal) This example also suggests a need for a more semantic evaluation: none of our metrics capture the fact that, arguably, Llama-CRAFTS showed at least a partial understanding of the “line" concept (and, in a way, even the “bottom right"), but Nebula did not.
+In Example G, A asks B to place two blocks on the “"other side"" of “"what you just made"" (i.e.
+the line of four blocks on the board).
+Such implicit references to spatial layouts and to existing structures are very common in the MDC dialogues.
+But we see another shortcoming of our metrics: because both models place a row of two green blocks on the ground (and because our shape metric allows for rotations on the horizontal plane), they achieve location, shape and overall F1 scores of 0.5, and type and color scores of 1.0, even though Llama-CRAFTS arguably understood “other side" and “90 degrees" better than Nebula.
+    7.3 Takeaways
+    Our qualitative analysis is consistent with the quantitative results in Table 7: Both models are adept at identifying simpler aspects of the task, such as action type and color, reflected in their high Type and Color F1 scores.
+Llama-CRAFTS also demonstrates a superior understanding of required block shapes (e.g., diagonals, corners), which aligns with its higher Shape F1 score.
+The primary differentiator, however, is spatial reasoning.
+Llama-CRAFTS’s improvement in this area, demonstrated by its higher Location F1 score, validates the impact of our synthetic data and its richer input representations.
+    Error Analysis
+    However, our error analysis illustrates that Llama-CRAFTS still makes significant errors in precise spatial understanding, indicating that much work remains to be done to solve this task.
+    Evaluation Challenges and the Role of Auxiliary Metrics
+    The qualitative analysis reveals a more nuanced challenge for fair evaluation: the overall F1 metric lacks the sensitivity to distinguish differences in error quality.
+As our qualitative analysis reveals, it is possible for two models to make different types of errors, with one being qualitatively "better" or closer to the correct solution, yet the F1 metric may not capture this distinction.
+This demonstrates that our proposed auxiliary metrics can play a crucial role, either by identifying the types of mistakes a model makes or by identifying cases where the model displayed at least a partial understanding of the instruction.
+At the same time, our analysis underscores the inherent challenges in automated evaluation for the BAP task.
+    The examples in Scenario 3 also illustrate the difficulty of creating a clear, unambiguous rubric for grading model output on an absolute scale, making human evaluation equally complex.
+This complicates not only absolute scoring but also comparative pairwise evaluations (i.e., "Is model A better than model B?").
+While our qualitative analysis includes examples where a clear distinction can be made, many cases are far more ambiguous.
+In fact, we initially considered human evaluation but ultimately did not pursue it due to these challenges.
+Also, given that the primary goal of this work is to provide a holistic revision of the entire BAP framework—spanning evaluation, data, and modeling—we concluded that a deep investigation into human evaluation methodologies was beyond the current scope and needed thorough investigation owing to the aforementioned challenges.
+This highlights the need for further research into BAP evaluation by building on our v2 benchmark.
+    Implications for LLMs
+    Finally, the persistence of these subtle spatial errors, as observed above in Section 7.2, suggests limitations in how current text-only LLMs handle fine-grained 3D spatial grounding in embodied tasks, specifically in a 3D, embodied, collaborative dialogue setting (see Section 7.5 below for more discussion in the broader context of related work), thus motivating future work.
+    7.4 Related Work: BAP evaluation
+    Chaturvedi, Thompson, and Asher (2024) also identify some of the issues with BAP evaluation on $D_{\mathrm{mc}}$ that we do, and hence propose evaluation on synthetic data.
+They design simpler, non-dialogue-based data with corresponding evaluation metrics.
+The data is single-turn and evaluates a model’s ability to follow an instruction in two scenarios: when the board is empty and the instruction is to build a single shape, or when a single shape is already present and the instruction involves placing or removing one block.
+    While this data is intentionally simple to test elementary Builder capabilities, it overly simplifies the task and therefore deviates significantly from the target MCBT/BAP.
+It lacks a dialogue component, game history, and relative spatial references such as "left" and "right" as the Builder’s position and orientation are not considered.
+The world state/built structure is also fairly minimal, either empty or containing a single shape.
+While useful for elementary testing, it is a bit unclear how well insights from this data translate to the full BAP task.
+In contrast, our synthetic data remains simpler than $D_{\mathrm{mc}}$ while also closely emulating the MCBT, addressing these limitations and preserving many of the key BAP features and challenges, such as diverse instructions, rich spatial relations and floating blocks for example (more in Sections 2.1.1 and 5).
+Although designed/motivated for training more than evaluation, its simplicity allows assessment of basic model competencies, while its emulation of the MCBT enhances its applicability to the overall BAP task, thus striking a better balance from an evaluation standpoint.
+Section 6.3 details additional benefits of evaluating on our synthetic data.
+    Additionally, our synthetic data evaluation, combined with fairer F1, enables meaningful performance comparisons between synthetic data and $D_{\mathrm{mc}}$.
+Chaturvedi, Thompson, and Asher (2024) are unable to conduct such comparisons (albeit understandably) due to the limitations of the legacy/original F1 metric.
+    At a high level, Chaturvedi, Thompson, and Asher (2024) propose evaluation on their synthetic data as a solution to BAP evaluation issues, whereas we take the complementary approach by addressing them directly to propose the BAP v2 evaluation benchmark.
+Additionally, as described above, we also highlight the importance of evaluation on our synthetic data and contrast it with theirs.
+    7.5 Related Work: Spatial Reasoning
+    Early efforts to understand spatial reasoning in language models focused on extracting spatial relations from text and observing what conclusions about a real world a model trained only on rich language data could learn (Mirzaee and Kordjamshidi, 2022).
+Spatial reasoning was often explored in simple text-based games (Shi, Zhang, and Lipani, 2022) or with carefully constructed reasoning tasks (Rizvi, Zhu, and Gurevych, 2024).
+    Spatial reasoning has also been explored in conjunction with other modalities.
+In Vision-and-Language Navigation many explore it over scene-graphs (Gopinathan et al., 2024; Li et al., 2024a).
+Others explore spatial reasoning with text and image inputs.
+Bisk et al.
+(2018) explore what vision and language activations models learn from block tasks represented in 2D images: mapping out how models interpreted spatial relations but also verbs like "mirror" or "twist".
+Du et al.
+(2024) also explore how VLMs can reason over space in photographs, but in reference to objects within those objects (i.e.
+an object being on the right of the photo but to the left of a person in the photo).
+However, Kamath, Hessel, and Chang (2023) found that VLMs spatial reasoning can lag well behind human performance on VQA-style tasks and Wang et al.
+(2024) find that VLMs often do not make adequate use of the images that they are given, especially when provided with rich text representations.
+    Unfortunately, there is not much parallel work exploring the spatial capabilities of LLMs on text-only inputs, particularly for embodied tasks.
+Yamada et al.
+(2024) explore some basic grid navigation abilities of prompted LLMs, which struggle to traverse simple grids represented in text.
+Aghzal, Plaku, and Yao (2024) consider the problem from the perspective of text-only grid path-planning, showing that models can effectively link coordinates and directions, but struggle to generalize to larger grids with more obstacles.
+Others have explored LLMs in playing chess (Feng et al., 2023a; Zhang et al., 2025b) or sliding puzzle games (Chollet et al., 2025) to some degree of success, but we found no works combining 3D text-only grids and embodied collaborative dialogue.
+Together, these gaps in the literature motivate the MDC setup: by excluding the visual modality used in other tasks, it allows us to explore the baseline spatial capabilities of text-only LLMs on embodied tasks in a small controlled setting that is still generalizable to concrete end-goals.
+    8 Future Work
+    We now discuss the implications of our work and outline directions for future research.
+This discussion is organized into two parts: the path forward for BAP (Section 8.1), and the broader impact of our work beyond BAP (Section 8.2).
+    8.1 The path forward for BAP
+    The introduction of the BAP v2 framework, while a significant step forward, also illuminates several key directions for future research.
+The work detailed in this paper suggests that the next advancements on the BAP task can be organized around three core pillars: evaluation, data, and modeling.
+    Advancing Evaluation
+    Section 7.3 discusses the persisting problems around both automated and human evaluation for BAP, in spite of the progress made in BAP v2.
+This highlights the need for further research into evaluation by building on our v2 benchmark.
+    Enhancing Synthetic Data Generation
+    We note the broader challenges in using LLMs to generate coherent, state-aware synthetic data for embodied dialogue tasks (Section 5.6).
+A promising research direction is to explore methods for overcoming these limitations, potentially leveraging LLMs to create even more diverse, complex, and naturalistic training dialogues that still maintain grounding in the 3D world state.
+    Improving Model Capabilities
+    With a new SOTA F1 score of 53.0, there is still considerable room for improvement. Key areas for future modeling work include
+    • Spatial Reasoning: Our analysis (Section 7) confirmed that highly accurate and fine-grained 3D spatial reasoning remains a primary bottleneck, even for top-performing models.
+Future models must develop more robust mechanisms for interpreting complex, perspective-dependent spatial language.
+    • Advanced Training Regimes: More sophisticated training strategies could unlock further gains.
+For instance, exploring curriculum learning, where models are first trained on simpler synthetic data before moving to the more complex human-generated dialogues, could prove effective.
+    • Architectural Exploration: The relative success of a smaller fine-tuned specialist model motivates deeper exploration of this paradigm.
+Future efforts can build on Llama-CRAFTS with more advanced architectures or fine-tuning techniques.
+Concurrently, the synthetic data could also be leveraged to improve the ICL performance of larger, general-purpose models (Section 6.4).
+    8.2 Broader implications of our work beyond BAP
+    While our focus in this work is on the BAP subtask within the MCBT setting, our contributions have broader applications and takeaways as well.
+Below, we discuss their relevance for other subtasks within MCBT, for research building directly on the MCBT, and for similar domains and tasks in the field; thus highlighting the potential applications of our work and directions for future research.
     Within MCBT
@@
 Additionally, other subtasks, such as deciding when the Builder should act versus speak or generating clarification questions during dialogue, could also benefit from our synthetic datasets.
-Our simulators and data were carefully designed to emulate the nuances of the MCBT itself (Section 5.1), making them directly usable or adaptable to these subtasks.
+Our simulators and data were carefully designed to emulate the nuances of the MCBT itself (Section 5), making them directly usable or adaptable to these subtasks.
 The simulators’ parameterized nature enables generating tailored data distributions to this end, supporting further advancements across the broader MCBT ecosystem.
@@
 Several works directly build upon or are inspired by the MCBT/MDC, including IGLU (Mohanty et al., 2024), Bonn et al.
-(2020), Thompson, Hunter, and Asher (2024), and Bonial et al.
+(2020), Thompson, Hunter, and Asher (2024), Bonial et al.
+(2021), and Madge et al.
 (2025).
@@
     For Other Similar Domains and Tasks
-Our findings also offer broader takeaways for research on tasks involving grounded, multimodal task-oriented dialogue systems operating in dynamic, spatially complex environments.
-First, in many realistic scenarios, data collection for complex situated dialogue tasks is often impractical or expensive.
-Our work illustrates that synthetic data generation can be a viable and scalable alternative to simply collecting more data.
-Second, we demonstrate that rich synthetic data can be generated for such tasks in game-based or simulated environments like Minecraft by carefully emulating task and game-specific nuances.
-This approach produces full embodied task oriented dialogues combining utterances, environment actions, and other game state information, offering a viable strategy for similar game-based/simulated and grounded dialog tasks –– an underexplored area as highlighted in Section 9.2.
-Finally, techniques like curriculum learning demonstrate potential and can be explored further in similar domains/tasks in combination with synthetic data.
-9 Related Work
-9.1 Task and Dataset
-We discuss related work to illustrate that the MDC and BAP remain a challenging and relevant testbed for grounded language instruction following.
-Vision-and-Language Navigation (VLN) (Anderson et al., 2018), and its dialog counterpart, Cooperative Vision-and-Dialog Navigation (CVDN) (Thomason et al., 2019), explore instruction following in photorealistic navigation settings.
-However, Minecraft dialogues are more complex due to their asynchronous communication, longer turn sequences, and dynamic environment changes during construction.
-Unlike navigation, where referring expressions target existing objects, construction often involves instructions for objects yet to be built.
-And although more recent navigation tasks require real vision, their underlying world state space (as defined by fixed viewpoints and the underlying navigation graph) is just as highly discretized.
-Our task does not require vision, but poses an arguably more challenging planning problem, since its action space is much larger (7623 possible actions vs.
-six actions in the vision-language navigation work).
-Tasks involving vision and language-based task completion beyond navigation have also been explored.
-ALFRED (Shridhar et al., 2020) is a benchmark for mapping natural language instructions and egocentric vision to action sequences for household tasks in simulated environments.
-However, similar to VLN and unlike the MDC, it lacks dialogue and relies solely on planner-based demonstrations instead of human ones.
-Efforts to combine vision and dialogue navigation with task completion include CEREALBAR (Suhr et al., 2019), which operates in a game-based setting.
-Yet, similar to CVDN, it uses turn-based dialogue, contrasting with the free-form, asynchronous communication of the MDC.
-This limitation is addressed by the more recent benchmark TEACh (Padmakumar et al., 2022), which includes free-form dialogue, similar to the MDC, and features over 3000 human-human interactive dialogues for completing household tasks in simulation and with object state changes.
-However, unlike the MDC, the instruction giver in TEACh is provided with a predefined plan of steps, while the Architect in the MDC independently devises a unique plan for constructing the target structure.
-This results in greater creativity and diversity in the MDC, with multiple possible plans for the same structure.
-Although the TEACh dataset is larger (by number of #dialogs), its dialogues and utterances are shorter.
-Additionally, the MDC allows for greater abstraction in instructions, such as referring to substructures, individual blocks, or shapes, whereas TEACh focuses on well-defined objects in the environment.
-Similar to VLN and CVDN, referring expressions in TEACh refer to objects that exist in the world, but construction instructions like those in the MDC frequently refer to objects that need to be built.
-Minecraft has garnered a lot of interest as an AI experimentation platform in recent years.
-Our previous work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) was among the first to use it to study grounded task-oriented dialogue, and introduced the MCBT and MDC.
-Facebook’s CraftAssist (Gray et al., 2019; Jernite et al., 2019; Szlam et al., 2019) is another such example, enabling two-way human-bot interactions where a human architect instructs an automated builder to build complex structures.
-Their data includes synthetically generated and crowdsourced instructions paired with logical tree structures consisting of action primitives, unlike the MDC, which features human-human dialogues with greater ambiguity, variety, and noisier Builder actions.
-MineDojo (Fan et al., 2022) focuses on creating versatile agents for diverse tasks via an internet-scale knowledge base, contrasting with tasks like IGLU (Mohanty et al., 2024; Kiseleva et al., 2022b, a) and MCBT, which prioritize grounded natural language dialogue and clarification for interactive agents.
-Among Minecraft-related efforts, IGLU is most closely related to the MDC/MCBT and was directly inspired by it.
-The IGLU dataset, unlike the MDC, is single turn only, and not dialogue based.
-It comprises 8136 instruction-action pairs, including 1056 ambiguous instructions with corresponding clarification questions.
-The task setting is similar to the MCBT but incorporates some simplifications which make it relatively less challenging.
-E.g., cardinal directions (e.g., North, South) are displayed in the voxel world, which simplifies instruction giving/following through the use of absolute rather than relative references (e.g., left, right).
-IGLU also introduces a very small multi-turn dataset (127 dialogues) based on this same task setting.
-It features strict turn-taking and clearly demarcated instructions and clarification questions.
-This contrasts with the MDC’s free-form, asynchronous dialogues, which are less structured and whose utterances have more diversity.
-Some other works also build on or are inspired by the MCBT/MDC.
-For instance, Bonn et al.
-(2020), Thompson, Hunter, and Asher (2024), and Bonial et al.
-(2021) provide additional linguistic annotations for the MDC.
-9.2 Synthetic Training Data
-Synthetic data has proven beneficial for various tasks, including those mentioned in Section 9.1.
-VLN agents, constrained by limited human instruction data and diversity in training environments, often struggle with complex language grounding and spatial language understanding.
-Kamath et al.
-(2023) address this by using synthetic data, achieving SOTA performance on the RxR dataset (Ku et al., 2020), while Wang et al.
-(2023) are similar in spirit and achieve SOTA on the aforementioned CVDN and other benchmarks.
-Similarly, Kang et al.
-(2023) demonstrate strong performance gains in low-data regimes for Visual Dialog Das et al.
-(2017) by leveraging synthetic data.
-Synthetic data has also been effective for dialogue systems, particularly those focused on text-only dialogues (Kim et al., 2022; Bao et al., 2023; Zhan et al., 2023).
-Zhan et al.
-(2024) is an example of extending this to multi-modal dialogue agents, demonstrating the effectiveness of synthetic visual descriptions in enhancing agents’ grounding capabilities.
-However, fewer efforts have addressed synthesizing task-oriented embodied dialogues (involving both utterances and environment actions).
-Padmakumar et al.
-(2023) were the first to design a framework to do so.
-They extend agenda-based dialogue simulation (Schatzmann and Young, 2009) to a multimodal embodied agent framework, and demonstrate the impact of the synthetic dialogues on the TEACh task (see Section 9.1 for a comparison of TEACh to MDC).
-Our work serves to provide another useful example in this general domain, further advancing synthetic data generation for such complex tasks.
-Dan, Han, and Roth (2021) is another closely related work that demonstrates the potential of synthetic data in the Blocks World domain.
-They use simpler synthetic data for the benchmarks of Bisk, Yuret, and Marcu (2016), which require understanding single-shot instructions that transform one world state to another using simulated 3D blocks.
-Blocks are viewed from a fixed bird’s eye perspective, initialized randomly in the initial world state, and uniquely identifiable.
-The varying Builder perspective and lack of easily identifiable referents, along with the need to understand utterances in a dialogue context, make the BAP task a much more challenging problem.
-[^7] Also, unlike traditional Blocks World, Minecraft allows blocks to float (requiring non-monotonic action sequences where placement is followed by removal), or attach to any side of an existing block.
-Our work on synthetic data thus serves as a demonstration of the potential for richer, embodied task-oriented dialogue tasks that extend beyond simpler Blocks World settings.
-10 Conclusion and Future Work
-In this work, we revisited the challenging Builder Action Prediction (BAP) subtask within the Minecraft Collaborative Building Task (MCBT) setting, addressing key shortcomings in evaluation and training data.
-We introduced BAP v2, an upgraded version of the task, comprising two core contributions: (1) an enhanced evaluation benchmark that includes a cleaner test set and fairer, more insightful metrics, and (2) diverse synthetic training datasets generated from novel Minecraft dialogue and target structure simulators.
-These contributions together provide a robust foundation for future research on the BAP task.
-Additionally, our synthetic data demonstrates the potential to train more performant and robust models, even with relatively simple training methods and architectures.
-Although modeling was not the primary focus of this work, we also illustrated the impact of the synthetic data and training methodologies on training a simple LLM and transformer-based model, thus validating the robustness of our approach, and setting the stage for more sophisticated models/LLMs going forward.
-Future work can therefore build on our work to explore more advanced LLMs and architectures while leveraging the BAP v2 framework, alongside training methodologies such as Curriculum Learning (CL).
-Further improvements could stem from experimenting with more sophisticated CL approaches or entirely new training paradigms to get even more out of the synthetic data.
-Additionally, as highlighted in Section 8.1, BAP evaluation remains a nuanced challenge, warranting deeper investigation.
-Finally, Section 8.4 outlined the broader implications of our work, while also highlighting potential directions for future research.
+    Our findings also offer a key takeaway for research on embodied, multimodal, task-oriented dialogue in dynamic and spatially complex settings.
+We demonstrate a methodology for generating rich synthetic data within game-based or simulated environments by carefully emulating task- and game-specific nuances.
+This approach yields complete embodied task-oriented dialogues that integrate utterances, environmental actions, and other game state information, providing a robust strategy for such complex dialogue settings—an area that remains underexplored, as highlighted in Section 5.6.
+This enables future work in the domain.
+    9 Conclusion
+    In this work, we addressed the challenges of the Builder Action Prediction (BAP) task, a demanding testbed for grounded instruction following within the Minecraft Collaborative Building Task.
+To foster more efficient and meaningful progress, we introduced BAP v2, a holistically re-examined and upgraded task framework.
+BAP v2 is built on three core contributions.
+First, we established an enhanced evaluation benchmark, featuring a cleaner test set and fairer, more insightful metrics.
+This new benchmark not only provides a more accurate measure of performance but also revealed that spatial reasoning is the primary performance bottleneck.
+Second, to address both data scarcity and the spatial reasoning challenge, we developed novel dialogue simulators to generate diverse synthetic training data rich in spatial language.
+Third, we leveraged this new data to train stronger models.
+Our new SOTA model, Llama-CRAFTS, utilizes rich, context-aware input representations and the synthetic data to achieve an F1 score of 53.0 —a significant 6 point improvement over previous work.
+    While this result marks a substantial advance, it also underscores the task’s remaining difficulty, even for SOTA LLMs.
+By systematically improving the evaluation, data, and modeling paradigms, the BAP v2 framework provides a robust and reliable foundation for future research.
+It paves the way for developing more sophisticated models with advanced training regimes, creating richer synthetic data, and exploring new evaluation techniques to push the boundaries of what is possible for interactive agents in complex, collaborative, and embodied environments.
     10 Appendix: Generating Synthetic Data
-Sections 11.111.2, and 11.3 provide additional details (beyond those covered in Section 5) on how each of the three dialog simulators implements the general simulation framework (Section 5.1.2) – specifically, steps 1 (Architect Planning), 3 (Architect Instruction), and 4 (Optional Clarification Exchange).
-For details on the elementary shapes used for generating shape-based target structures (Section 5.3.1) see Section 11.4.
+    Sections 10.210.3, and 10.4 provide additional details (beyond those covered in Section 5) on how each of the three dialogue simulators implements the general simulation framework (Section 5) – specifically,
+    Step 10.1.
+    1 (Architect Planning),
+    Step 10.2.
+    3 (Architect Instruction), and
+    Step 10.3.
+    4 (Optional Clarification Exchange). Section 10.5 provides details on the elementary shapes used to define shape-based target structures (Section 5.4.1).
     Note
-All distances are measured using Manhattan Distance.
-For Steps 3 and 4, we outline only the key elements of our template-based generation framework.
+    All distances are measured using Manhattan Distance. For
+    Step 10.4.
+    3 and
+    Step 10.5.
+    4, we outline only the key elements of our template-based generation framework.
 Further details will be included in our eventual code and data release.
 Thus, we omit detailed lexical and syntactic variations of utterances, providing only a few representative examples.
-11.1 Random target structures and dialogs
-Architect Planning
+    10.1 Target structures
+    Some global constraints on target structures are enforced during their generation:
+    1. (a)
+      The structure must have at least one block on the ground.
+    2. (b)
+      The entire structure must be "connected," i.e., each block shares at least one face or edge with at least one other block.
+    10.2 Random target structures and dialogues ($D_{\mathrm{r}}$)
+    Step 10.6.
+    1: Architect Planning
     Next Action:
-The first four actions in the dialog are placements.
+    The first four actions in the dialogue are placements.
 Subsequent actions are sampled with a 90% probability for placement and 10% for removal.
@@
 Candidate positions are defined as follows.
-If the board is empty (start of dialog), candidate positions are limited to the ground.
-But if at least one block exists, a candidate position must be within the "connected neighborhood" of an existing block.
-The "connected neighborhood" includes all locations within a Manhattan distance of 2 that share a face or edge with an existing block (but not a sole corner).
+If the board is empty (start of dialogue), candidate positions are limited to the ground.
+But if at least one block exists, a candidate position must be within the ”connected neighborhood” of an existing block.
+The ”connected neighborhood” includes all locations within a Manhattan distance of 2 that share a face or edge with an existing block (but not a sole corner).
 This forms a 3x3x3 cube around the block, excluding the 8 diagonal corners but including the 6 adjacent and 12 2D diagonal locations.
-Reference Block:
-If needed, the algorithm selects a reference block as the closest existing block to the chosen next block.
-However, if the block from the last action that took place has the same distance, it takes precedence.
-A reference block can either be an existing block in the built structure (the last added one or a randomly selected one) or the last block that was removed from it.
-There are two special cases where a reference block is not required or cannot be assigned.
-These are discussed later in Section Special Cases with No Reference Block.
-Architect Instruction
+    Step 10.8.
+    3: Generating the Main Dialogue
     Next Action
-The instruction specifies the color of the block, and when necessary, includes an optional adjective such as "floating" if temporary supporting blocks are required.
-For example, in Figure 11, "blue" and "floating" are used, while in Figure 12, "purple" and "yellow" are used.
+    The instruction specifies the color of the block (e.g. “blue”), and when necessary, includes an optional adjective such as “floating” if temporary supporting blocks are required.
     Reference Block
     A phrase is needed to refer to the reference block, determined by the following priority rules (from highest to lowest):
-  1. 1.
-    If the reference block is the last placed/removed block, the phrase used is "the last block you added/removed" etc.
+    1. (a)
+      If the reference block is the last placed/removed block, the phrase used is “the last block you added/removed” etc.
 If the last action was a placement, the algorithm may omit the reference block in the instruction (ellipsis) with a small probability.
-  2. 2.
-    If the reference block’s color is unique in the built structure, it is referred to as "the [COLOR] block," where [COLOR] represents the block’s color.
-  3. 3.
-    If the color is not unique, the block is disambiguated using its spatial position relative to other blocks of the same color, allowing references such as "the leftmost [COLOR] block," etc.
-For example, in Figure 11, the reference block is described as "the last block" and "the last block you placed." In Figure 12, it is referred to as "the closest yellow block" and "the orange one."
+    2. (b)
+      If the reference block’s color is unique in the built structure, it is referred to as “the [COLOR] block,” where [COLOR] represents the block’s color.
+    3. (c)
+      If the color is not unique, the block is disambiguated using its spatial position relative to other blocks of the same color, allowing references such as “the closest/leftmost [COLOR] block,” etc.
     Optional Clarification Exchange
     The elements of the instruction that may be omitted include the location (i.e., the spatial relation + reference block) or the color of the next action.
-B’s clarification questions, such as "Where?" or "What color?" are then generated accordingly, followed by A’s response.
-For example, in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
+B’s clarification questions, such as “Where?” or “What color?” are then generated accordingly, followed by A’s response.
     Special Cases with No Reference Block
@@
     If the next action is a removal and the block to be removed is the last placed block, the following adjustments are made while keeping everything else unchanged: 1) No reference block is needed.
-2) A’s instruction will be a simple command such as "remove that block".
+2) A’s instruction will be a simple command such as “remove that block”.
 3) No clarification exchanges are generated, as there is nothing to omit in the instruction regarding the location or color of the block to be removed.
-Start of Dialog
-At the start of the dialog (empty board), the following modifications are made: 1.
-The first action is the of a block on the ground (randomly sampled from target structure blocks on the ground).
-2.
-Since there are no prior blocks, A’s instruction will be a simple directive, such as "First start by placing an orange block on the ground".
-11.2 Blocks-based dialogs for shape-based structures
-Architect Planning
-Recall that a target structure is provided as input to the dialog simulation.
+    Start of Dialogue
+    At the start of the dialogue (empty board), the following modifications are made: (a) The first action is the of a block on the ground (randomly sampled from target structure blocks on the ground).
+(b) Since there are no prior blocks, A’s instruction will be a simple directive, such as “First start by placing an orange block on the ground”.
+    10.3 Blocks-based dialogues for shape-based structures ($D_{\mathrm{sb}}$)
+    Step 10.10.
+    1: Architect Planning
+    Recall that a target structure is provided as input to the dialogue simulation.
 In this iteration of the simulation, blocks to be placed next are selected from those that remain unplaced, given the current built structure and the target.
@@
     Multiple Blocks (if any)
-Next, we check whether multiple blocks (i.e., apart from the next block) can be placed "greedily" using the same reference block and spatial relation.
+    Next, we check whether multiple blocks (i.e., apart from the next block) can be placed ”greedily” using the same reference block and spatial relation.
 For instance, if the next block is a yellow block placed to the left of the reference block, and another yellow block exists further to the left, A may plan to place both simultaneously for efficiency.
@@
 It then selects the longest contiguous sequence of such blocks connected to the next block.
-Architect Instruction
-The approach is similar to the one in Section 11.1 but additionally accounts for the number of blocks when multiple blocks are placed.
-For example, in Figure 16, the instruction specifies "three yellow blocks."
+    Step 10.12.
+    3: Generating the Main Dialogue
+    The approach is similar to the one in Section 10.2 but additionally accounts for the number of blocks when multiple blocks are placed (e.g. to specify “three yellow blocks”).
     Optional Clarification Exchange
@@
     Special Cases with No Reference Block
-There is only one special case, which is the same as the "Start of Dialog" one described in Section Special Cases with No Reference Block.
-11.3 Shape-based dialogs for shape-based structures
-Architect Planning
-Recall that a target structure consisting of two shape instances is provided as input to the dialog simulation.
+    There is only one special case, which is the same as the ”Start of Dialogue” one described in Section Special Cases with No Reference Block.
+    10.4 Shape-based dialogues for shape-based structures ($D_{\mathrm{ss}}$)
+    Step 10.14.
+    1: Architect Planning
+    Recall that a target structure consisting of two shape instances is provided as input to the dialogue simulation.
     Next Shape and Its Starting Block
@@
 The algorithm first selects a block within the shape to define spatial relations needed to specify where this shape should be built (e.g.
-wrt the reference block).
-This block also serves as the starting point for B when building the shape (the order of the remaining blocks is decided by B’s planning in Step 5 of the simulation algorithm of Section 5.1.2).
+wrt.
+the reference block).
+This block also serves as the starting point for B when building the shape (the order of the remaining blocks is decided by B’s planning in Step 5 of the simulation algorithm of Section 5).
 Potential starting blocks are always located in the bottom corners of the shape.
@@
     For the second shape being placed, the reference block is the last placed block of the first shape. No reference block is used for the first shape.
-Architect Instruction
+    Step 10.16.
+    3: Generating the Main Dialogue
     Next Shape
     A shape’s size can be described in multiple ways depending on its dimensions.
-For 1D shapes such as rows, the description includes a single number, such as "four blocks long" or "three blocks tall." For 2D shapes like diagonals and planes, various formats can be used, such as "3x3", "two blocks long and three tall", etc.
-For example, in Figure 16, the instructions specify "4 yellow blocks" and "five block long."
+For 1D shapes such as rows, the description includes a single number, such as “four blocks long” or “three blocks tall.” For 2D shapes like diagonals and planes, various formats can be used, such as “3x3”, “two blocks long and three tall”, etc.
     The instruction also includes the shape’s color and name.
 We provide examples here for three shapes – rows, diagonals and planes.
-The primary shape names are "row," "diagonal," and "plane," though synonyms can be used.
-A horizontal row may be called a "line," while a vertical row may be referred to as a "column" or "tower." Similarly, a horizontal plane may be described as a "layer," whereas a vertical plane may be called a "wall." A horizontal diagonal is typically referred to as a "diagonal" or "diagonal line," but a vertical diagonal can also be described as a "staircase" or "stairway." For example, in Figure 16, the terms "pillar" and "line" are used for shape names, while "yellow" and "red" are used for shape colors.
+The primary shape names are “row,” “diagonal,” and “plane,” though synonyms can be used.
+A horizontal row may be called a “line,” while a vertical row may be referred to as a “column”, “pillar”, or “tower.” Similarly, a horizontal plane may be described as a “layer,” whereas a vertical plane may be called a “wall.” A horizontal diagonal is typically referred to as a “diagonal” or “diagonal line,” but a vertical diagonal can also be described as a “staircase” or “stairway.”
     Reference Block
     This applies only to the second shape being placed.
-Since the reference block is the last placed block, we use phrases such as "the last block you placed," similar to the approach in Sections 11.1 and 11.2.
+Since the reference block is the last placed block, we use phrases such as “the last block you placed,” similar to the approach in Sections 10.2 and 10.3.
     Optional Clarification Exchange
-The approach is similar to the one in Sections 11.1 and 11.2, with two additional elements that can be omitted in an instruction.
-Recall that this data also includes spatial relations that use B as a spatial anchor (Section 5.1.3).
-For example, in Figure 16, A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row.
-Such directional spatial relations may be omitted from the instruction.
-Additionally, the dimensions/size of a shape may also be omitted.
-11.4 Shape-based target structures
-The six elementary shapes used to generate the structures are defined as follows:
+    The approach is similar to the one in Sections 10.2 and 10.3, but now, the size/dimension and relative location/orientation of the shape to be placed (using expressions such as “going to the left of you” that use B as a spatial anchor, c.f.
+Section 5.2.3) can also be omitted.
+    10.5 Elementary shapes
+    The six elementary shapes used to generate the structures for $D_{\mathrm{sb}}$ and $D_{\mathrm{ss}}$ are defined as follows:
     Rows
@@
     Diagonals
-consist of at least 3 blocks, and are planar (along any of the $xy$, $yz$, or $xz$ planes). Horizontal diagonals lie along the $xz$ plane, vertical along either $xy$ or $yz$.
+    consist of at least 3 blocks, and are planar (along any of the $xy$, yzyz, or $xz$ planes). Horizontal diagonals lie along the $xz$ plane, vertical along either $xy$ or $yz$.
     T-shapes
@@
     11 Appendix: Experimental Details
-12.1 Curriculum Learning (CL) Algorithm
-We use the Curriculum Arrangement algorithm from the CL framework proposed by Xu et al.
-(2020).
-Specifically, we sort the training dataset $D$ by the difficulty metric(s) and divided into $N$ equal-sized buckets $\{C_{i}:i=1,2,\dotsc,N\}$, corresponding to increasing difficulty levels from $C_{1}$ (easiest) to $C_{N}$ (hardest).
-The curriculum consists of $N$ stages $\{S_{i}:i=1,2,\dotsc,N\}$.
-In each stage $S_{i}$, a $1/N$ fraction of examples is sampled from each preceding bucket $\{C_{j}:j=1,2,\dotsc,i\}$, with the proportion of difficult examples gradually increasing across stages from 0 to eventually match the original dataset distribution.
-Within each stage, examples are still shuffled to maintain local stochastics, and there is no overlap between stages to prevent overfitting.
-At each stage, the model is trained for one epoch.
-After reaching $S_{N}$, the model should be ready for the original distribution in $D$.
-Hence, an additional stage $S_{N+1}$ is introduced, covering the entire dataset $D$, and training continues until convergence.
     11.1 How much of each synthetic dataset to use during train time?
     We provide further details of the procedure in Section 6.1.
-We first train models on the three synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) separately to study how increasing the training data affects performance.
-Focusing on the saturation regions of the learning curves—where performance plateaus and additional data has diminishing returns—we identify various "candidate" data amounts (#items) that yield near-peak performance for each dataset.
+We first train models on the three synthetic datasets ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) separately to study how increasing the training data affects performance.
+Focusing on the saturation regions of the learning curves—where performance plateaus and additional data has diminishing returns—we identify various ”candidate” data amounts (#items) that yield near-peak performance for each dataset.
 Using these data amounts, we construct a grid of combinations across the three datasets, where each grid point corresponds to a different data mix and proportion of each dataset.
-For each grid point, we train a model on the combined $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ and $D_{\mathrm{mc}}$ datasets.
+For each grid point, we train a model on the combined $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ and $D_{\mathrm{mc}}$ datasets.
 The combination that achieves the best performance on $D_{\mathrm{mc}}$ is selected.
 Thus, the approach yields a data mix that optimizes $D_{\mathrm{mc}}$ performance while also being highly performant on the synthetic data as much as possible.
-Also, in the best data mixes, the #items per dataset generally followed the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$, which matches their difficulty ordering (as explored for CL in Section 6.2.1).
+Also, in the best data mixes, the #items per dataset generally followed the order $D_{\mathrm{sb}}$ ¡ $D_{\mathrm{ss}}$ ¡ $D_{\mathrm{r}}$, which matches their difficulty ordering.
     11.2 Experimental Setup for GRU-based Models
@@
 Early stopping is applied when the loss on the held-out development set increases monotonically for ten epochs.
-The development set for $M_{\mathrm{agg}}$ and $M_{\mathrm{agg+cl}}$, used for early stopping, is constructed from the validation splits of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ (Table 3) to match the dataset distribution in the training set.
+The development set for $M_{\mathrm{agg}}$ and $M_{\mathrm{agg+cl}}$, used for early stopping, is constructed from the validation splits of $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ (Table 3) to match the dataset distribution in the training set.
     Model Hyperparameters
     We continue to use the same hyperparameter values as those of the best model in Jayannavar, Narayan-Chen, and Hockenmaier (2020).
-12.4 Experimental Setup for BERT-based Models
-We provide further details for the experimental setup of $M_{\mathrm{mc}}^{\prime}$, $M_{\mathrm{agg}}^{\prime}$, and $M_{\mathrm{agg+cl}}^{\prime}$ (apart from those mentioned in Section 6.3.1).
-Training
-The training setup is similar to Section 12.3, with the following differences: a learning rate of 0.0001 for $M_{\mathrm{mc}}^{\prime}$ and 0.0005 for $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg+cl}}^{\prime}$, a batch size of 16, and a learning rate schedule consisting of linear warmup followed by inverse square root decay with 4000 warmup steps.
-Model Hyperparameters
-The transformer decoder consists of 5 layers with a representation dimensionality of 256, 2 attention heads, and an inner feedforward layer dimensionality of 512.
-The model uses uncased BERT with a maximum input length of 512 tokens, the highest allowed.
-Other hyperparameters are similar to Section 12.3.
-13 Appendix: Additional Results
-This section contains the BAP v2 metrics for the examples in Sections 7.2 and 8.1.
-Example ID Model Type Color Location Shape Overall
-Example 1 Baseline 1.0 1.0 0.0 0.0 0.0
-Best 1.0 1.0 1.0 1.0 1.0
-Example 2 Baseline 1.0 1.0 0.0 0.5 0.0
-Best 1.0 1.0 1.0 1.0 1.0
-Example 3 Baseline 1.0 1.0 0.0 0.0 0.0
-Best 1.0 1.0 1.0 1.0 1.0
-Table 10: BAP v2 metrics by Example ID for Section 7.2 (range $[0,1]$)
-Example ID Model Type Color Location Shape Overall
-Example A Baseline 0.67 0.67 0.0 0.0 0.0
-Best 0.5 0.5 0.5 0.5 0.5
-Example B Baseline 1.0 1.0 0.0 0.33 0.0
-Best 1.0 1.0 0.0 1.0 0.0
-Example C Baseline 0.5 0.5 0.0 0.0 0.0
-Best 0.5 0.5 0.0 0.0 0.0
-Example D Baseline 0.5 0.5 0.5 0.0 0.5
-Best 1.0 1.0 0.5 0.0 0.5
-Table 11: BAP v2 metrics by Example ID for Section 8.1 (range $[0,1]$)
     12 Appendix: Annotation Setup
@@
 The NEB (non-empty board) subset is annotated by the 10 non-expert annotators, with each BAP item annotated by two annotators.
-Disagreements were resolved by the expert annotator, and the workload was evenly distributed among the annotators (in terms of the number of dialogs per annotator).
+Disagreements were resolved by the expert annotator, and the workload was evenly distributed among the annotators (in terms of the number of dialogues per annotator).
 The smaller and easier-to-annotate EB (empty board) subset was exclusively annotated by the expert annotator.
-Annotators were instructed to annotate one dialog at a time, processing items within a dialog in chronological order.
-This ensures they accurately understand the evolving game context and remain unbiased by future "unseen" events in a dialog.
-15 Appendix: Extended Discussion on Concurrent Work
-Apart from the discussion on concurrent work and Nebula’s evaluation in Section 8.3, a slightly more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in this section given some other relevant aspects of their work.
-(The detailed tables for Nebula’s performance on the full v2 benchmark mentioned in Section 8.3 are also provided here (Tables 13 and 14).)
-15.1 Evaluation
-Chaturvedi, Thompson, and Asher (2024) also identify some of the issues with BAP evaluation on $D_{\mathrm{mc}}$ that we do, and hence propose evaluation on synthetic data.
-They design simpler, non-dialog-based data with corresponding evaluation metrics.
-The data is single-turn and evaluates a model’s ability to follow an instruction in two scenarios: when the board is empty and the instruction is to build a single shape, or when a single shape is already present and the instruction involves placing or removing one block.
-While this data is intentionally simple to test elementary Builder capabilities, it overly simplifies the task and therefore deviates significantly from the target MCBT/BAP.
-It lacks a dialog component, game history, and relative spatial references such as "left" and "right" as the Builder’s position and orientation are not considered.
-The world state/built structure is also fairly minimal, either empty or containing a single shape.
-While useful for elementary testing, it is a bit unclear how well insights from this data translate to the full BAP task.
-In contrast, our synthetic data remains simpler than $D_{\mathrm{mc}}$ while also closely emulating the MCBT, addressing these limitations and preserving many of the key BAP features and challenges, such as diverse instructions, rich spatial relations and floating blocks for example (more in Sections 2.1.1 and 5).
-Although designed/motivated for training more than evaluation, its simplicity allows assessment of basic model competencies, while its emulation of the MCBT enhances its applicability to the overall BAP task, thus striking a better balance from an evaluation standpoint.
-Section 8.2 details additional benefits of evaluating on our synthetic data.
-Additionally, our synthetic data evaluation, combined with fairer F1, enables meaningful performance comparisons between synthetic data and $D_{\mathrm{mc}}$ (an example utility of this is the insights it provides to motivate Curriculum Learning).
-Chaturvedi, Thompson, and Asher (2024) are unable to conduct such comparisons (albeit understandably) due to the limitations of the legacy/original F1 metric.
-At a high level, Chaturvedi, Thompson, and Asher (2024) propose evaluation on their synthetic data as a solution to BAP evaluation issues, whereas we take the complementary approach by addressing them directly to propose the BAP v2 evaluation benchmark.
-Additionally, as described above, we also highlight the importance of evaluation on our synthetic data and contrast it with theirs.
-15.2 Nebula+
-While Chaturvedi, Thompson, and Asher (2024) primarily design their synthetic data for evaluation, they also report results from finetuning Nebula (Llama-3-8B) jointly on it and $D_{\mathrm{mc}}$.
-We refer to this model as Nebula+.
-(Recall that the original Nebula model was trained on $D_{\mathrm{mc}}$ alone.) Nebula+ improves performance on their synthetic data but remains similar to Nebula on $D_{\mathrm{mc}}$ (achieving an F1 score of 39.1% under the legacy F1, compared to Nebula’s 39.2%).
-This is fairly understandable given that their synthetic data, designed mainly for evaluation rather than training, does not capture sufficient task complexity, as discussed above.
-In contrast, as shown in Section 6, our approach of training jointly on our richer synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and $D_{\mathrm{mc}}$ improves performance on both (and then further boosted by Curriculum Learning).
-Similar to Nebula’s evaluation in Section 8.3, we evaluate Nebula+ here.
-Detailed results for the full v2 benchmark are provided in Tables 15 and 16.
-Table 12 presents Nebula and Nebula+ performance (micro F1) on both our synthetic data and the original $D_{\mathrm{mc}}$ test sets (extending the previous Table 9 with an additional row for Nebula+).
-The key trends and conclusions remain consistent with those previously discussed for Nebula—demonstrating the impact and robustness of our v2 evaluation, providing a fairer model assessment, and reinforcing the importance of evaluating on our synthetic data.
-Additionally, Table 12 shows that Nebula+ slightly underperforms compared to Nebula on our synthetic data while maintaining similar performance on $D_{\mathrm{mc}}$.
-This suggests that improvements on their synthetic data do not necessarily translate to better performance on ours (fairly understandable as their synthetic data does not capture sufficient task complexity).
-Model Test Data
-$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
-Nebula 51.5 33.1 30.9 50.1
-Nebula+ 48.8 32.4 30.4 50.1
-Table 12: Nebula and Nebula+’s performance across datasets (v2 evaluation on $D_{\mathrm{mc}}$)
-Dataset Type Color Location Shape Overall
-EB 92.8 92.8 78.3 84.8 78.3
-NEB 79.0 77.4 47.2 59.0 46.4
-Overall 80.6 79.1 50.7 61.9 50.1
-Table 13: Micro F1 scores for the Nebula model under the v2 benchmark
-Dataset Type Color Location Shape Overall
-EB 96.8 96.8 84.5 90.5 84.5
-NEB 81.7 80.9 47.1 64.1 46.5
-Overall 82.9 82.2 50.1 66.3 49.6
-Table 14: Macro F1 scores for the Nebula model under the v2 benchmark
-Dataset Type Color Location Shape Overall
-EB 94.8 94.8 80.6 83.7 80.6
-NEB 78.9 77.7 46.9 58.7 46.3
-Overall 80.7 79.6 50.7 61.6 50.1
-Table 15: Micro F1 scores for the Nebula+ model under the v2 benchmark
-Dataset Type Color Location Shape Overall
-EB 97.5 97.5 83.6 88.9 83.6
-NEB 82.8 82.0 48.6 64.1 48.1
-Overall 84.0 83.3 51.5 66.1 51.0
-Table 16: Macro F1 scores for the Nebula+ model under the v2 benchmark
+Annotators were instructed to annotate one dialogue at a time, processing items within a dialogue in chronological order.
+This ensures they accurately understand the evolving game context and remain unbiased by future ”unseen” events in a dialogue.
     Acknowledgements.
-We would like to thank Hetvi Patel, Sana Madhavan, Marc Canby, Risham Sidhu, Siddarth Madala, Rajarshi Haldar, Katya Yegorova, Adam Davies, Ansel Blume, and Sara Aghajanzadeh for their valuable help with the data annotation.
+    We would like to thank Hetvi Patel, Sana Madhavan, Marc Canby, Siddarth Madala, Rajarshi Haldar, Katya Yegorova, Adam Davies, Ansel Blume, and Sara Aghajanzadeh for their valuable help with the annotation, and Dan Roth, Martha Palmer, Janna Doppa and Sriraam Natarajan and all other members of our CwC team for many helpful discussions earlier in this project.
 This work was supported by Contract W911NF-15-1-0461 with the US Defense Advanced Research Projects Agency (DARPA) Communicating with Computers Program and the Army Research Office (ARO).
@@
 The work utilizes resources Kindratenko et al.
-(2020) supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign.
+(2020) supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign, as well as the Delta advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois.
+Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.